### v2

# Online Convex Optimization with Heavy Tails:

Abstract
In Online Convex Optimization (OCO), when the stochastic gradient has a finite variance, many algorithms provably work and guarantee a sublinear regret. However, limited results are known if the gradient estimate has a heavy tail, i.e., the stochastic gradient only admits a finite $\mathsf{p}$-th central moment for some $\mathsf{p}\in\left(1,2\right]$. Motivated by it, this work examines different old algorithms for OCO (e.g., Online Gradient Descent) in the more challenging heavy-tailed setting. Under the standard bounded domain assumption, we establish new regrets for these classical methods without any algorithmic modification. Remarkably, these regret bounds are fully optimal in all parameters (can be achieved even without knowing $\mathsf{p}$), suggesting that OCO with heavy tails can be solved effectively without any extra operation (e.g., gradient clipping). Our new results have several applications. A particularly interesting one is the first provable and optimal convergence result for nonsmooth nonconvex optimization under heavy-tailed noise without gradient clipping. Furthermore, we explore broader settings (e.g., smooth OCO) and extend our ideas to optimistic algorithms to handle different cases simultaneously.
1 Introduction
This paper studies the online learning problem with convex losses, also known as Online Convex Optimization (OCO), a widely applicable framework that learns under streaming data [5, 12, 32, 40]. OCO has tons of implications for both designing and analyzing algorithms in different areas, for example, stochastic optimization [10, 19, 28], PAC learning [4], control theory [1, 13], etc.
In an OCO problem, a learning algorithm $\mathsf{A}$ would interact with the environment in $T$ rounds, where $T\in\mathbb{\mathbb{N}}$ can be either known or unknown. Formally, in each round round $t$, the learner $\mathsf{A}$ first decides an output $\boldsymbol{x}_{t}\in\mathcal{X}$ from a convex feasible set $\mathcal{X}\subseteq\mathbb{R}^{d}$, then the environment reveals a convex loss function $\ell_{t}:\mathcal{X}\to\mathbb{R}$, and $\mathsf{A}$ incurs a loss of $\ell_{t}(\boldsymbol{x}_{t})$. After $T$ many rounds, the quantity measuring the algorithm’s performance is called regret, defined relative to any fixed competitor $\boldsymbol{x}\in\mathcal{X}$ as follows:
$\mathsf{R}_{T}^{\mathsf{A}}(\boldsymbol{x})\triangleq\sum_{t=1}^{T}\ell_{t}(\boldsymbol{x}_{t})-\ell_{t}(\boldsymbol{x}).$
In the classical setting, instead of observing full information about $\ell_{t}$, the learner $\mathsf{A}$ is only guaranteed to receive a subgradient $\nabla\ell_{t}(\boldsymbol{x}_{t})\in\partial\ell_{t}(\boldsymbol{x}_{t})$ at its decision, where $\partial\ell_{t}(\boldsymbol{x}_{t})$ denotes the subdifferential set of $\ell_{t}$ at $\boldsymbol{x}_{t}$ [38]. This turns out to be enough for our purpose of minimizing the regret, since any OCO problem can be reduced to an Online Linear Optimization (OLO) instance via the inequality $\ell_{t}(\boldsymbol{x}_{t})-\ell_{t}(\boldsymbol{x})\leq\left\langle\nabla\ell_{t}(\boldsymbol{x}_{t}),\boldsymbol{x}_{t}-\boldsymbol{x}\right\rangle$, which holds due to convexity. Under the standard bounded domain assumption, i.e., $\mathcal{X}$ has a finite diameter $D$, many classical algorithms, e.g., Online Gradient Descent ($\mathsf{OGD}$) [56], guarantee an optimal sublinear regret $GD\sqrt{T}$ for $G$-Lipschitz $\ell_{t}$. Even better, in the case that computing an exact subgradient is intractable, and one could only query a stochastic estimate $\boldsymbol{g}_{t}$ satisfying $\mathbb{\mathbb{E}}\left[\boldsymbol{g}_{t}\mid\boldsymbol{x}_{t}\right]\in\partial\ell_{t}(\boldsymbol{x}_{t})$, the $\mathsf{OGD}$ algorithm can still solve OCO effectively with a provable $(G+\sigma)D\sqrt{T}$ regret bound in expectation if the stochastic noise $\boldsymbol{g}_{t}-\nabla\ell_{t}(\boldsymbol{x}_{t})$ has a bounded second moment $\sigma^{2}$ for some $\sigma\geq 0$, which is called the finite variance condition.
However, many works have pointed out that even for the easier stochastic optimization (i.e., $\ell_{t}=F$ for a common $F$), the typical finite variance assumption is too optimistic and can be violated in different tasks [14, 42, 51], and their observations suggest that the stochastic gradient only admits a finite $\mathsf{p}$-th central moment upper bounded by $\sigma^{\mathsf{p}}$ for some $\mathsf{p}\in\left(1,2\right]$, which is named heavy-tailed noise. This new assumption generalizes the classical finite variance condition ($\mathsf{p}=2$) and becomes challenging when $\mathsf{p}<2$. A particular evidence is that the famous Stochastic Gradient Descent ($\mathsf{SGD}$) algorithm [37] (which is exactly $\mathsf{OGD}$ for stochastic optimization) provably diverges [51].
Though heavy-tailed stochastic optimization has been extensively studied [24, 31, 39], limited results are known for OCO with heavy tails. The only work under this topic that we are aware of is [53], which established a parameter-free regret bound in high probability (more discussions provided later). However, their algorithm includes many nontrivial modifications like gradient clipping and significantly deviates from the existing simple OCO algorithms used in practice. Especially, consider $\mathsf{OGD}$ as an example. Though the heavy-tailed issue is known, $\mathsf{OGD}$ (or just think of it as $\mathsf{SGD}$) still works (sometimes very well) in practice even without gradient clipping and is arguably one of the most popular optimizers, which seemingly contradicts the theory of unconvergence mentioned before. This indicates that, for classical OCO algorithms under heavy-tailed noise, a huge gap exists between the empirical convergence (or even the effective practical performance) and theoretical guarantees. Therefore, we are naturally led to the following question:
In what context can old OCO algorithms work under heavy tails, in what sense, and to what extent?
1.1 Contributions
Motivated by the above question, we examine three classical algorithms for OCO: Online Gradient Descent ($\mathsf{OGD}$) [56], Dual Averaging ($\mathsf{DA}$) [30, 49], and $\mathsf{AdaGrad}$ [10, 28], and answer it as follows:
Under the standard bounded domain assumption, the in-expectation regret $\mathbb{\mathbb{E}}\left[\mathsf{R}_{T}^{\mathsf{A}}(\boldsymbol{x})\right]$ is finite and optimal for any $\mathsf{A}\in\left\{\mathsf{OGD},\mathsf{DA},\mathsf{AdaGrad}\right\}$, without any algorithmic modification.
In detail, our new results for heavy-tailed OCO are summarized here:
  • We prove the only and the first optimal regret bound $\mathbb{\mathbb{E}}\left[\mathsf{R}_{T}^{\mathsf{A}}(\boldsymbol{x})\right]\lesssim GD\sqrt{T}+\sigma DT^{1/\mathsf{p}},\forall\boldsymbol{x}\in\mathcal{X}$ for any $\mathsf{A}\in\left\{\mathsf{OGD},\mathsf{DA},\mathsf{AdaGrad}\right\}$. Remarkably, $\mathsf{AdaGrad}$ can achieve this result without knowing any of the Lipschitz parameter $G$, noise level $\sigma$, and tail index $\mathsf{p}$.
  • We extend the analysis of $\mathsf{OGD}$ to Online Strongly Convex Optimization with heavy tails and establish the first provable result $\mathbb{\mathbb{E}}\left[\mathsf{R}_{T}^{\mathsf{OGD}}(\boldsymbol{x})\right]\lesssim\frac{G^{2}\log T}{\mu}+\frac{\sigma^{\mathsf{p}}D^{2-\mathsf{p}}}{\mu^{\mathsf{p}-1}}T^{2-\mathsf{p}},\forall\boldsymbol{x}\in\mathcal{X}$, where $\mu>0$ is the modulus of strong convexity and $T^{0}$ should be read as $\log T$.
Based on the new regret bounds for OCO with heavy tails, we provide the following applications:
  • For nonsmooth convex optimization with heavy tails, we show the first optimal in-expectation rate $GD/\sqrt{T}+\sigma D/T^{1-1/\mathsf{p}}$ achieved without gradient clipping, which applies to both the average iterate and last iterate, demonstrating that $\mathsf{SGD}$ does converge once the domain is bounded.
  • For nonsmooth nonconvex optimization with heavy tails, we show the first provable sample complexity of $G^{2}\delta^{-1}\epsilon^{-3}+\sigma^{\frac{\mathsf{p}}{\mathsf{p}-1}}\delta^{-1}\epsilon^{-\frac{2\mathsf{p}-1}{\mathsf{p}-1}}$ for finding a $(\delta,\epsilon)$-stationary point without gradient clipping. In addition, we also establish the first lower bound for nonsmooth nonconvex optimization under heavy tails, matching our sample complexity in the high accuracy and noisy regime (i.e., $\epsilon$ is small enough with $\sigma>0$). These two results together provide a nearly complete characterization of the complexity of finding $(\delta,\epsilon)$-stationary points in the heavy-tailed setting. Moreover, we give the first convergence result when the problem-dependent parameters (like $G$, $\sigma$, and $\mathsf{p}$) are unknown in advance, resolving a question asked by [23].
Furthermore, we explore broader settings. For example, when each $\ell_{t}$ is $H$-smooth, we present new regrets that extend the classical $L^{\star}$ bounds [32, 43] to the heavy-tailed noise case. As an important implication, we show that $\mathsf{SGD}$ converges at a rate of $HD^{2}/T+\sigma D/T^{1-1/\mathsf{p}}$ for smooth convex optimization with heavy tails (even for the last iterate). Finally, we extend our ideas to optimistic algorithms to address various cases simultaneously and employ optimistic algorithms to give the first provable result for Hölder smooth nonconvex optimization under heavy tails, where the problem-dependent parameters can be either known or unknown.
1.2 Discussion on [53]
As noted, [53] is the only work for OCO with heavy tails, as far as we know. There are two major discrepancies between them and us. First, they consider the case where the feasible set $\mathcal{X}$ is unbounded and aim to establish a parameter-free regret bound, i.e., the regret bound has a linear dependency on $\left\|\boldsymbol{x}\right\|$ (up to an extra $\mathrm{polylog}\left\|\boldsymbol{x}\right\|$) for any competitor $\boldsymbol{x}\in\mathcal{X}$. Second, they focus on high-probability rather than in-expectation analysis. As such, their regret is in the form of $\mathsf{R}_{T}^{\mathsf{A}}(\boldsymbol{x})\lesssim(G+\sigma)\left\|\boldsymbol{x}\right\|T^{1/\mathsf{p}},\forall\boldsymbol{x}\in\mathcal{X}$ (up to extra polylogarithmic factors) with high probability. Without a doubt, their setting is harder than ours implying their bound is stronger as it can convert to an in-expectation regret $\mathbb{\mathbb{E}}\left[\mathsf{R}_{T}^{\mathsf{A}}(\boldsymbol{x})\right]\lesssim(G+\sigma)DT^{1/\mathsf{p}}$ for any bounded domain $\mathcal{X}$ with a diameter $D$.
We emphasize that the motivation behind [53] differs heavily from ours. They aim to solve heavy-tailed OCO with a new proposed method that contains many nontrivial technical tricks, including gradient clipping, artificially added regularization, and solving the additional fixed-point equation. However, their result cannot reflect why the existing simple OCO algorithms like $\mathsf{OGD}$ work in practice under heavy-tailed noise. In contrast, our goal is to examine whether, when, and how the classical OCO algorithms work under heavy tails, thereby filling the missing piece in the literature.
Moreover, we would like to mention two limitations of [53]. First, though the $T^{1/\mathsf{p}}$ regret seems tight as it matches the lower bound [29, 35, 47], this may not be the best, since an optimal bound should recover the standard $\sqrt{T}$ regret in the deterministic case (i.e., $\sigma=0$), as one can imagine. This suggests that their bound is not entirely optimal. Second, we remark that they require knowing both problem-dependent parameters $G$, $\sigma$, $\mathsf{p}$ and time horizon $T$ in the algorithm, which may be hard to satisfy in the online setting. In comparison, our regret bound $GD\sqrt{T}+\sigma DT^{1/\mathsf{p}}$ is fully optimal in all parameters. Importantly, $\mathsf{AdaGrad}$ can achieve it while oblivious to the problem information.
2 Preliminary
Notation. $\mathbb{\mathbb{N}}$ denotes the set of natural numbers (excluding 0). $\left[T\right]\triangleq\left\{1,\dots,T\right\},\forall T\in\mathbb{\mathbb{N}}$. $a\land b\triangleq\min\left\{a,b\right\}$ and $a\lor b\triangleq\max\left\{a,b\right\}$. We write $a\lesssim b$ (resp., $a\gtrsim b$) if $a\leq Cb$ (resp., $a\geq Cb$) for a universal constant $C>0$. $\left\lfloor\cdot\right\rfloor$ and $\left\lceil\cdot\right\rceil$ respectively represent the floor and ceiling functions. $\left\langle\cdot,\cdot\right\rangle$ denotes the Euclidean inner product and $\left\|\cdot\right\|\triangleq\sqrt$$\left\langle\cdot$$\cdot\right\rangle$} is the standard $2$-norm. Given $\boldsymbol{x}\in\mathbb{R}^{d}$ and $D>0$, $\mathcal{B}^{d}(\boldsymbol{x},D)$ is the Euclidean ball in $\mathbb{R}^{d}$ centered at $\boldsymbol{x}$ with a radius $D$. In the case $\boldsymbol{x}=\boldsymbol{0}$, we use the shorthand $\mathcal{B}^{d}(D)$. Given $A\subseteq\mathbb{R}^{d}$, $\mathsf{int}A$ stands for the interior points of $A$. For nonempty closed convex $A\subseteq\mathbb{R}^{d}$, $\Pi_{A}$ is the Euclidean projection operator onto $A$. For a convex function $f$, $\partial f(\boldsymbol{x})$ denotes its subgradient set at $\boldsymbol{x}$.
Remark 1.
We choose the Euclidean norm only for simplicity. Extending the results in this work to any general norm via Bregman divergence is straightforward.
This work studies OCO in the context of Assumption 1.
Assumption 1.
We consider the following series of assumptions:
  • $\mathcal{X}\subset\mathbb{R}^{d}$ is a nonempty closed convex set bounded by $D$, i.e., $\sup_{\boldsymbol{x},\boldsymbol{y}\in\mathcal{X}}\left\|\boldsymbol{x}-\boldsymbol{y}\right\|\leq D$.
  • $\ell_{t}:\mathcal{X}\to\mathbb{R}$ is closed convex for all $t\in\left[T\right]$.
  • $\ell_{t}$ is $G$-Lipschitz on $\mathcal{X}$, i.e., $\left\|\nabla\ell_{t}(\boldsymbol{x})\right\|\leq G,\forall\boldsymbol{x}\in\mathcal{X},\nabla\ell_{t}(\boldsymbol{x})\in\partial\ell_{t}(\boldsymbol{x})$, for all $t\in\left[T\right]$.
  • Given a point $\boldsymbol{x}_{t}\in\mathcal{X}$ at the $t$-th iteration, one can query $\boldsymbol{g}_{t}\in\mathbb{R}^{d}$ satisfying $\nabla\ell_{t}(\boldsymbol{x}_{t})\triangleq\mathbb{\mathbb{E}}\left[\boldsymbol{g}_{t}\mid\mathcal{F}_{t-1}\right]\in\partial\ell_{t}(\boldsymbol{x}_{t})$ and $\mathbb{\mathbb{E}}\left[\left\|\boldsymbol{\epsilon}_{t}\right\|^{\mathsf{p}}\right]\leq\sigma^{\mathsf{p}}$ for some $\mathsf{p}\in\left(1,2\right]$ and $\sigma\geq 0$, where $\mathcal{F}_{t}\triangleq\sigma(\boldsymbol{g}_{1},\dots,\boldsymbol{g}_{t})$ denotes the natural filtration and $\boldsymbol{\epsilon}_{t}\triangleq\boldsymbol{g}_{t}-\nabla\ell_{t}(\boldsymbol{x}_{t})$ is the stochastic noise.
Remark 2.
$D$ is recognized as known, like ubiquitously assumed in the OCO literature. Moreover, $\boldsymbol{x}_{t}$ denotes the decision/output of the online learning algorithm by default.
Remark 3.
The idea presented in this work can be extended to the composite setting (i.e., $\ell_{t}+\psi_{t}$, where $\psi_{t}$ is convex and assumed to be known at time t1t-1) via the proximal update, provably.
In Assumption 1, the first three points are standard, and the fourth is the heavy-tailed noise assumption. In particular, $\mathsf{p}=2$ recovers the standard finite variance condition.
3 Old Algorithms under Heavy Tails
In this section, we revisit three classical algorithms for OCO: $\mathsf{OGD}$, $\mathsf{DA}$, and $\mathsf{AdaGrad}$, whose regret bounds are well-studied in the finite variance case but remain unknown under heavy-tailed noise.
The basic idea of proving these algorithms work under heavy tails is to leverage the boundness property of $\mathcal{X}$. We will describe it in more detail using $\mathsf{OGD}$ as an illustrated example. The analysis of $\mathsf{DA}$ follows a similar way at a high level, but differs in some details. However, though $\mathsf{AdaGrad}$ can be viewed as $\mathsf{OGD}$ with an adaptive stepsize, the way to utilize the boundness property is entirely different. All formal proofs are deferred to the appendix due to space limitations.
3.1 New Regret for Online Gradient Descent
Algorithm 1 Online Gradient Descent ($\mathsf{OGD}$) [56]
Input: initial point $\boldsymbol{x}_{1}\in\mathcal{X}$, stepsize $\eta_{t}>0$
for $t=1$ to $T$ do
$\boldsymbol{x}_{t+1}=\Pi_{\mathcal{X}}(\boldsymbol{x}_{t}-\eta_{t}\boldsymbol{g}_{t})$
end for
We begin from arguably the most basic algorithm for OCO, Online Gradient Descent ($\mathsf{OGD}$).
A well known analysis. The regret bound of $\mathsf{OGD}$ has been extensively studied [12, 32, 40]. The most well known analysis is perhaps the following one: for any $\boldsymbol{x}\in\mathcal{X}$, there is
$\left\|\boldsymbol{x}_{t+1}-\boldsymbol{x}\right\|^{2}=\left\|\Pi_{\mathcal{X}}(\boldsymbol{x}_{t}-\eta_{t}\boldsymbol{g}_{t})-\Pi_{\mathcal{X}}(\boldsymbol{x})\right\|^{2}\leq\left\|\boldsymbol{x}_{t}-\eta_{t}\boldsymbol{g}_{t}-\boldsymbol{x}\right\|^{2},$
where the inequality holds by the nonexpansive property of $\Pi_{\mathcal{X}}$. Expanding both sides and rearranging terms yield that
$\left\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t}-\boldsymbol{x}\right\rangle\leq\frac{\left\|\boldsymbol{x}_{t}-\boldsymbol{x}\right\|^{2}-\left\|\boldsymbol{x}_{t+1}-\boldsymbol{x}\right\|^{2}}{2\eta_{t}}+\frac{\eta_{t}\left\|\boldsymbol{g}_{t}\right\|^{2}}{2}.$ (1)
If $\boldsymbol{g}_{t}$ admits a finite variance, i.e., $\mathsf{p}=2$ in Assumption 1, taking expectations on both sides, then following a standard analysis for $\eta_{t}=\frac{D}{(G+\sigma)\sqrt{t}}$ (or $\eta_{t}=\frac{D}{(G+\sigma)\sqrt{T}}$ if $T$ is known) gives the regret
$\mathbb{\mathbb{E}}\left[\mathsf{R}_{T}^{\mathsf{OGD}}(\boldsymbol{x})\right]\lesssim(G+\sigma)D\sqrt{T},\forall\boldsymbol{x}\in\mathcal{X}.$
However, the step of taking expectations on the R.H.S. of (1) crucially relies on the finite variance condition of $\boldsymbol{g}_{t}$. Therefore, one may naturally think $\mathsf{OGD}$ would not guarantee a finite regret if $\mathsf{p}<2$.
A less well known analysis [^1] To clarify, the phrase “less well known” is compared to the first one. This analysis itself is also well known [22, 32].. As discussed, the failure of the above proof under heavy-tailed noise is due to (1). Therefore, if a tighter inequality than (1) exists, then it might be possible to show that $\mathsf{OGD}$ still works for $\mathsf{p}<2$. However, does it exist?
Actually, there is another less well known analysis to produce a better inequality than (1). That is, first showing for any $\boldsymbol{x}\in\mathcal{X}$, by the optimality condition of the update rule,
$\left\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t+1}-\boldsymbol{x}\right\rangle\leq\frac{\left\langle\boldsymbol{x}_{t}-\boldsymbol{x}_{t+1},\boldsymbol{x}_{t+1}-\boldsymbol{x}\right\rangle}{\eta_{t}}=\frac{\left\|\boldsymbol{x}_{t}-\boldsymbol{x}\right\|^{2}-\left\|\boldsymbol{x}_{t+1}-\boldsymbol{x}\right\|^{2}-\left\|\boldsymbol{x}_{t}-\boldsymbol{x}_{t+1}\right\|^{2}}{2\eta_{t}},$
and then obtaining
$\left\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t}-\boldsymbol{x}\right\rangle\leq\frac{\left\|\boldsymbol{x}_{t}-\boldsymbol{x}\right\|^{2}-\left\|\boldsymbol{x}_{t+1}-\boldsymbol{x}\right\|^{2}}{2\eta_{t}}+\left\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t}-\boldsymbol{x}_{t+1}\right\rangle-\frac{\left\|\boldsymbol{x}_{t}-\boldsymbol{x}_{t+1}\right\|^{2}}{2\eta_{t}}.$ (2)
Note that (2) is tighter than (1) as $\left\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t}-\boldsymbol{x}_{t+1}\right\rangle\leq\left\|\boldsymbol{g}_{t}\right\|\left\|\boldsymbol{x}_{t}-\boldsymbol{x}_{t+1}\right\|\leq\frac{\eta_{t}\left\|\boldsymbol{g}_{t}\right\|^{2}}{2}+\frac{\left\|\boldsymbol{x}_{t}-\boldsymbol{x}_{t+1}\right\|^{2}}{2\eta_{t}}$, where the first step is due to Cauchy-Schwarz inequality and the second one is by AM-GM inequality.
Handle $\mathsf{p}<2$ in a simple way. Though we have tightened (1) into (2), can inequality (2) help to overcome heavy tails? The answer is surprisingly positive, and our solution is fairly simple. Instead of directly applying AM-GM inequality in the second step, we recall $\boldsymbol{g}_{t}=\nabla\ell_{t}(\boldsymbol{x}_{t})+\boldsymbol{\epsilon}_{t}$ and use triangle inequality to obtain
$\left\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t}-\boldsymbol{x}_{t+1}\right\rangle\leq\left\|\boldsymbol{g}_{t}\right\|\left\|\boldsymbol{x}_{t}-\boldsymbol{x}_{t+1}\right\|\leq\left(\left\|\nabla\ell_{t}(\boldsymbol{x}_{t})\right\|+\left\|\boldsymbol{\epsilon}_{t}\right\|\right)\left\|\boldsymbol{x}_{t}-\boldsymbol{x}_{t+1}\right\|.$ (3)
On the one hand, by $\left\|\nabla\ell_{t}(\boldsymbol{x}_{t})\right\|\leq G$ and AM-GM inequality, there is
$\left\|\nabla\ell_{t}(\boldsymbol{x}_{t})\right\|\left\|\boldsymbol{x}_{t}-\boldsymbol{x}_{t+1}\right\|\leq G\left\|\boldsymbol{x}_{t}-\boldsymbol{x}_{t+1}\right\|\leq\eta_{t}G^{2}+\frac{\left\|\boldsymbol{x}_{t}-\boldsymbol{x}_{t+1}\right\|^{2}}{4\eta_{t}}.$ (4)
On the other hand, let $\mathsf{p}_{\star}\triangleq\frac{\mathsf{p}}{\mathsf{p}-1}$ and $\mathsf{C}(\mathsf{p})\triangleq\frac{(4\mathsf{p}-4)^{\mathsf{p}-1}}{\mathsf{p}^{\mathsf{p}}}$, we have
$\left\|\boldsymbol{\epsilon}_{t}\right\|\left\|\boldsymbol{x}_{t}-\boldsymbol{x}_{t+1}\right\|$ $=\left(\frac{4\eta_{t}}{\mathsf{p}_{\star}}\right)^{\frac{1}{\mathsf{p}_{\star}}}\left\|\boldsymbol{\epsilon}_{t}\right\|\left\|\boldsymbol{x}_{t}-\boldsymbol{x}_{t+1}\right\|^{1-\frac{2}{\mathsf{p}_{\star}}}\cdot\left(\frac{\mathsf{p}_{\star}\left\|\boldsymbol{x}_{t}-\boldsymbol{x}_{t+1}\right\|^{2}}{4\eta_{t}}\right)^{\frac{1}{\mathsf{p}_{\star}}}$
$\overset{(a)}{\leq}\frac{\left(\frac{4\eta_{t}}{\mathsf{p}_{\star}}\right)^{\frac{\mathsf{p}}{\mathsf{p}_{\star}}}\left\|\boldsymbol{\epsilon}_{t}\right\|^{\mathsf{p}}\left\|\boldsymbol{x}_{t}-\boldsymbol{x}_{t+1}\right\|^{\mathsf{p}-\frac{2\mathsf{p}}{\mathsf{p}_{\star}}}}{\mathsf{p}}+\frac{\left\|\boldsymbol{x}_{t}-\boldsymbol{x}_{t+1}\right\|^{2}}{4\eta_{t}}$
$\overset{(b)}{\leq}\mathsf{C}(\mathsf{p})\eta_{t}^{\mathsf{p}-1}\left\|\boldsymbol{\epsilon}_{t}\right\|^{\mathsf{p}}D^{2-\mathsf{p}}+\frac{\left\|\boldsymbol{x}_{t}-\boldsymbol{x}_{t+1}\right\|^{2}}{4\eta_{t}},$ (5)
where $(a)$ is by Young’s inequality and $(b)$ is due to $\left\|\boldsymbol{x}_{t}-\boldsymbol{x}_{t+1}\right\|\leq D$, $\mathsf{p}_{\star}=\frac{\mathsf{p}}{\mathsf{p}-1}$, and $\mathsf{C}(\mathsf{p})=\frac{(4\mathsf{p}-4)^{\mathsf{p}-1}}{\mathsf{p}^{\mathsf{p}}}$. Next, we plug (4) and (5) back into (3), then combine with (2) to know
$\left\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t}-\boldsymbol{x}\right\rangle\leq\frac{\left\|\boldsymbol{x}_{t}-\boldsymbol{x}\right\|^{2}-\left\|\boldsymbol{x}_{t+1}-\boldsymbol{x}\right\|^{2}}{2\eta_{t}}+\eta_{t}G^{2}+\mathsf{C}(\mathsf{p})\eta_{t}^{\mathsf{p}-1}\left\|\boldsymbol{\epsilon}_{t}\right\|^{\mathsf{p}}D^{2-\mathsf{p}}.$ (6)
Notably, the term $\left\|\boldsymbol{\epsilon}_{t}\right\|^{\mathsf{p}}$ has a correct exponent $\mathsf{p}$. Thus, we can safely take expectations on both sides. Finally, a standard analysis yields the following Theorem 1 (see Appendix A for a formal proof).
Theorem 1.
Under Assumption 1, taking $\eta_{t}=\frac{D}{G\sqrt{t}}\land\frac{D}{\sigma t^{1/\mathsf{p}}}$ in $\mathsf{OGD}$ (Algorithm 1), we have
$\mathbb{\mathbb{E}}\left[\mathsf{R}_{T}^{\mathsf{OGD}}(\boldsymbol{x})\right]\lesssim GD\sqrt{T}+\sigma DT^{1/\mathsf{p}},\forall\boldsymbol{x}\in\mathcal{X}.$
As far as we know, Theorem 1 is the first and the only provable result for $\mathsf{OGD}$ under heavy tails. Remarkably, it is not only tight in $T$ [29, 35, 47] but also fully optimal in all parameters, in contrast to the bound $(G+\sigma)DT^{1/\mathsf{p}}$ of [53]. This reveals that OCO with heavy tails can be optimally solved as effectively as the finite variance case once the domain is bounded, a classical condition adapted in many existing works.
Strongly convex functions. We highlight that the above idea can also be applied to Online Strongly Convex Optimization and leads to a sublinear regret $T^{2-\mathsf{p}}$ better than $T^{1/\mathsf{p}}$. This extension can be found in Appendix A.
3.2 New Regret for Dual Averaging
Algorithm 2 Dual Averaging ($\mathsf{DA}$) [30, 49]
Input: initial point $\boldsymbol{x}_{1}\in\mathcal{X}$, stepsize $\eta_{t}>0$
for $t=1$ to $T$ do
$\boldsymbol{x}_{t+1}=\Pi_{\mathcal{X}}(\boldsymbol{x}_{1}-\eta_{t}\sum_{s=1}^{t}\boldsymbol{g}_{s})$
end for
Remark 4.
It is known that $\mathsf{DA}$ is a special realization of the more general Follow-the-Regularized-Leader $(\mathsf{FTRL})$ framework [27]. To keep the work concise, we focus only on $\mathsf{DA}$. The key idea for proving Theorem 2 can be directly extended to show new regret for $\mathsf{FTRL}$ under heavy-tailed noise.
We turn our attention to the second candidate, the Dual Averaging ($\mathsf{DA}$) algorithm, which is given in Algorithm 2. Though $\mathsf{DA}$ coincides with $\mathsf{OGD}$ when $\mathcal{X}=\mathbb{R}^{d}$ and $\eta_{t}=\eta$, these two methods in general are not equivalent and can have significant performance differences in practice. Therefore, it is also important to understand $\mathsf{DA}$ under heavy tails.
Despite the proof strategies for $\mathsf{OGD}$ and $\mathsf{DA}$ are in different flavors (even for $\mathsf{p}=2$), the basic idea presented before for $\mathsf{OGD}$ still works here, i.e., apply the boundness property of $\mathcal{X}$ to make the term $\left\|\boldsymbol{\epsilon}_{t}\right\|$ have a correct exponent. Armed with this thought, we can prove the following new regret bound for $\mathsf{DA}$ under heavy-tailed noise. We refer the reader to Appendix B for its proof.
Theorem 2.
Under Assumption 1, taking $\eta_{t}=\frac{D}{G\sqrt{t}}\land\frac{D}{\sigma t^{1/\mathsf{p}}}$ in $\mathsf{DA}$ (Algorithm 2), we have
$\mathbb{\mathbb{E}}\left[\mathsf{R}_{T}^{\mathsf{DA}}(\boldsymbol{x})\right]\lesssim GD\sqrt{T}+\sigma DT^{1/\mathsf{p}},\forall\boldsymbol{x}\in\mathcal{X}.$
As far as we know, Theorem 2 is the first provable and optimal regret for $\mathsf{DA}$ under heavy tails. It guarantees the same tight bound as in Theorem 1 up to different constants.
3.3 New Regret for $\mathsf{AdaGrad}$
Algorithm 3 $\mathsf{AdaGrad}$ [10, 28]
Input: initial point $\boldsymbol{x}_{1}\in\mathcal{X}$, stepsize $\eta>0$
for $t=1$ to $T$ do
$\eta_{t}=\eta V_{t}^{-1/2}$ where $V_{t}=\sum_{s=1}^{t}\left\|\boldsymbol{g}_{s}\right\|^{2}$
$\boldsymbol{x}_{t+1}=\Pi_{\mathcal{X}}(\boldsymbol{x}_{t}-\eta_{t}\boldsymbol{g}_{t})$
end for
Remark 5.
Algorithm 3 is also named $\mathsf{AdaGrad\text{-}Norm}$ (e.g., [48]). We simply call it $\mathsf{AdaGrad}$. It is straightforward to generalize Theorem 3 below to the per-coordinate update version.
Although Theorems 1 and 2 are optimal, they both suffer from an undesired point. That is, the stepsize $\eta_{t}=\frac{D}{G\sqrt{t}}\land\frac{D}{\sigma t^{1/\mathsf{p}}}$ requires knowing all problem-dependent parameters. However, it may not be easy to obtain them in an online setting. Especially, it heavily depends on the prior information about the tail index $\mathsf{p}$, which is hard to know (even approximately) in advance. In other words, they both lack the adaptive property to an unknown environment.
To handle this issue, we consider $\mathsf{AdaGrad}$, a classical adaptive algorithm for OCO. As can be seen, $\mathsf{AdaGrad}$ is just $\mathsf{OGD}$ with an adaptive stepsize. However, it is this adaptive stepsize that can help us to overcome the above undesired point.
Theorem 3.
Under Assumption 1, taking $\eta=D/\sqrt{2}$ in $\mathsf{AdaGrad}$ (Algorithm 3), we have
$\mathbb{\mathbb{E}}\left[\mathsf{R}_{T}^{\mathsf{AdaGrad}}(\boldsymbol{x})\right]\lesssim GD\sqrt{T}+\sigma DT^{1/\mathsf{p}},\forall\boldsymbol{x}\in\mathcal{X}.$
Remark 6.
We also establish a similar result for $\mathsf{DA}$ with an adaptive stepsize. See Theorem 9 in Appendix B for details.
Theorem 3 provides the first regret bound for $\mathsf{AdaGrad}$ under heavy tails. Impressively, it is optimal even without knowing any of $G$, $\sigma$, and $\mathsf{p}$. This surprising result once again demonstrates the power of the adaptive method, indicating it is robust to an unknown environment and even heavy-tailed noise, which may partially explain the favorable performance of many adaptive optimizers designed based on $\mathsf{AdaGrad}$ like $\mathsf{RMSProp}$ [46] and $\mathsf{Adam}$ [19].
We point out that the key to establishing Theorem 3 differs from the idea used before for $\mathsf{OGD}$ and $\mathsf{DA}$. Actually, Theorem 3 can be obtained in an embarrassingly simple way. It is known that $\mathsf{AdaGrad}$ with $\eta=D/\sqrt{2}$ on a bounded domain guarantees the following path-wise regret
$\sum_{t=1}^{T}\left\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t}-\boldsymbol{x}\right\rangle\lesssim D\sqrt{\sum_{t=1}^{T}\left\|\boldsymbol{g}_{t}\right\|^{2}}.$ (7)
Observe that $\sqrt{\sum_{t=1}^{T}\left\|\boldsymbol{g}_{t}\right\|^{2}}\lesssim\sqrt{\sum_{t=1}^{T}\left\|\nabla\ell_{t}(\boldsymbol{x}_{t})\right\|^{2}}+\sqrt{\sum_{t=1}^{T}\left\|\boldsymbol{\epsilon}_{t}\right\|^{2}}\leq G\sqrt{T}+\left(\sum_{t=1}^{T}\left\|\boldsymbol{\epsilon}_{t}\right\|^{\mathsf{p}}\right)^{\frac{1}{\mathsf{p}}}$, where the last step is due to $\left\|\cdot\right\|_{2}\leq\left\|\cdot\right\|_{\mathsf{p}}$ for any $\mathsf{p}\in\left[1,2\right]$. After taking expectations on both sides of (7) and applying Hölder’s inequality to obtain $\mathbb{\mathbb{E}}\left[\left(\sum_{t=1}^{T}\left\|\boldsymbol{\epsilon}_{t}\right\|^{\mathsf{p}}\right)^{\frac{1}{\mathsf{p}}}\right]\leq\left(\sum_{t=1}^{T}\mathbb{\mathbb{E}}\left[\left\|\boldsymbol{\epsilon}_{t}\right\|^{\mathsf{p}}\right]\right)^{\frac{1}{\mathsf{p}}}\leq\sigma T^{\frac{1}{\mathsf{p}}}$, we conclude Theorem 3. To make the work self-consistent, we produce the formal proof of Theorem 3 in Appendix C.
4 Applications
We provide some applications based on the new regret bounds established in Section 3. The basic problem we study is optimizing a single objective $F$, which could be either convex or nonconvex.
4.1 Nonsmooth Convex Optimization
In this section, we consider nonsmooth convex optimization with heavy tails.
Convergence of the average iterate. First, we focus on the average-iterate convergence. By the classical online-to-batch conversion [4], the following corollary holds.
Corollary 1.
Under Assumption 1 for $\ell_{t}(\boldsymbol{x})=\left\langle\nabla F(\boldsymbol{x}_{t}),\boldsymbol{x}\right\rangle$ and let $\bar{\boldsymbol{x}}_{T}\triangleq\frac{1}{T}\sum_{t=1}^{T}\boldsymbol{x}_{t}$, for any $\mathsf{A}\in\left\{\mathsf{OGD},\mathsf{DA},\mathsf{AdaGrad}\right\}$, we have
$\mathbb{\mathbb{E}}\left[F(\bar{\boldsymbol{x}}_{T})-F(\boldsymbol{x})\right]\leq\frac{\mathbb{\mathbb{E}}\left[\mathsf{R}_{T}^{\mathsf{A}}(\boldsymbol{x})\right]}{T}\lesssim\frac{GD}{\sqrt{T}}+\frac{\sigma D}{T^{1-\frac{1}{\mathsf{p}}}},\forall\boldsymbol{x}\in\mathcal{X}.$
Proof.
By convexity, $F(\bar{\boldsymbol{x}}_{T})-F(\boldsymbol{x})\leq\frac{\sum_{t=1}^{T}F(\boldsymbol{x}_{t})-F(\boldsymbol{x})}{T}\leq\frac{\mathsf{R}_{T}^{\mathsf{A}}(\boldsymbol{x})}{T}$ is valid for any OCO algorithm $\mathsf{A}$. We conclude from invoking Theorems 1, 2 and 3. ∎
To the best of our knowledge, Corollary 1 gives the first and optimal convergence rate for these three algorithms in stochastic optimization with heavy tails. Especially, it implies that once the domain is bounded, the widely implemented $\mathsf{SGD}$ algorithm provably converges under heavy-tailed noise without any algorithmic change considered in many prior works, e.g., gradient clipping [24, 31].
We are only aware of two works [25, 47] based on Stochastic Mirror Descent ($\mathsf{SMD}$) [29] that gave convergence results without clipping. However, they share a common shortcoming, i.e., their bounds are both in the form of $(G+\sigma)D/T^{1-1/\mathsf{p}}$, which cannot recover the optimal rate $GD/\sqrt{T}$ when $\sigma=0$.
Lastly, we highlight that for $\mathsf{A}=\mathsf{AdaGrad}$, Corollary 1 is not only optimal but also adaptive to the tail index $\mathsf{p}$. As far as we know, no result has achieved this property before. This once again evidences the benefit of adaptive gradient methods.
Convergence of the last iterate. Next, we consider the more challenging last-iterate convergence, which has a long history in stochastic optimization and fruitful results in the case of $\mathsf{p}=2$ (see, e.g., [11, 16, 33, 41, 55]). However, less is known about heavy-tailed problems. So far, only two works [25, 34] have established the last-iterate convergence. The former is based on $\mathsf{SMD}$, and the latter employs gradient clipping in $\mathsf{SGD}$. Unfortunately, their rates are both in the suboptimal order $(G+\sigma)D/T^{1-1/\mathsf{p}}$.
We will provide an optimal last-iterate rate based on the following lemma, which reduces the last-iterate convergence to an online learning problem.
Lemma 1 (Theorem 1 of [9]).
Suppose $\boldsymbol{x}_{1},\dots,\boldsymbol{x}_{T}$ and $\boldsymbol{y}_{1},\dots,\boldsymbol{y}_{T}$ are two sequences of vectors satisfying $\boldsymbol{x}_{t}\in\mathcal{X},$ $\boldsymbol{x}_{1}=\boldsymbol{y}_{1}$ and
$\boldsymbol{y}_{t+1}=\boldsymbol{y}_{t}+\frac{T-t}{T}\left(\boldsymbol{x}_{t+1}-\boldsymbol{x}_{t}\right).$ (8)
Given a convex function $F(\boldsymbol{x})$, let $\ell_{t}(\boldsymbol{x})=\left\langle\nabla F(\boldsymbol{y}_{t}),\boldsymbol{x}\right\rangle$. Then for any online learner $\mathsf{A}$, we have
$F(\boldsymbol{y}_{T})-F(\boldsymbol{x})\leq\frac{\mathsf{R}_{T}^{\mathsf{A}}(\boldsymbol{x})}{T},\forall\boldsymbol{x}\in\mathcal{X}.$
We emphasize that the stochastic gradient $\boldsymbol{g}_{t}$ received by $\mathsf{A}$ is an estimate of $\nabla F(\boldsymbol{y}_{t})$ instead of $\nabla F(\boldsymbol{x}_{t})$. This flexibility is due to the generality of the OCO framework. Moreover, for $\mathsf{OGD}$, suppose there is no projection step, then (8) is equivalent to $\boldsymbol{y}_{t+1}=\boldsymbol{y}_{t}-\frac{T-t}{T}\eta_{t}\boldsymbol{g}_{t}$, which can be viewed as $\mathsf{SGD}$ with a stepsize $\frac{T-t}{T}\eta_{t}$. For proof of Lemma 1, we refer the interested reader to [9].
Corollary 2.
Under Assumption 1 for $\ell_{t}(\boldsymbol{x})=\left\langle\nabla F(\boldsymbol{y}_{t}),\boldsymbol{x}\right\rangle$, where $\boldsymbol{y}_{t}$ satisfies (8), for any $\mathsf{A}\in\left\{\mathsf{OGD},\mathsf{DA},\mathsf{AdaGrad}\right\}$, we have
$\mathbb{\mathbb{E}}\left[F(\boldsymbol{y}_{T})-F(\boldsymbol{x})\right]\leq\frac{\mathbb{\mathbb{E}}\left[\mathsf{R}_{T}^{\mathsf{A}}(\boldsymbol{x})\right]}{T}\lesssim\frac{GD}{\sqrt{T}}+\frac{\sigma D}{T^{1-\frac{1}{\mathsf{p}}}},\forall\boldsymbol{x}\in\mathcal{X}.$
Proof.
Combine Lemma 1 and Theorems 1, 2 and 3 to conclude. ∎
As far as we know, Corollary 2 is the first optimal last-iterate convergence rate for stochastic convex optimization with heavy tails, closing the gap in existing works.
One may notice that $\boldsymbol{y}_{t}$ itself is not the decision made by the online learner and naturally may ask whether $\boldsymbol{x}_{t}$ ensures the last-iterate convergence if we simply pick $\ell_{t}=F$. The answer turns out to be positive at least for $\mathsf{OGD}$ (which is equivalent to $\mathsf{SGD}$ now). However, to prove this result, we rely on a technique specialized to stochastic optimization recently developed by [25, 50]. To not diverge from the topic of OCO, we defer the last-iterate convergence of $\mathsf{OGD}$ (i.e., $\mathbb{\mathbb{E}}\left[F(\boldsymbol{x}_{T})\right]$) to Appendix D, in which Theorem 10 gives a general result for any stepsize $\eta_{t}$ (even for a broader function class) and Corollary 5 shows a last-iterate rate similar to Corollary 2 (up to an extra logarithmic factor) under the same stepsize $\eta_{t}=\frac{D}{G\sqrt{t}}\land\frac{D}{\sigma t^{1/\mathsf{p}}}$ as in Theorem 1. If $T$ is assumed to be known, Corollary 5 also shows how to set the stepsize in $\mathsf{OGD}$ to converge in the optimal rate $\frac{GD}{\sqrt{T}}+\frac{\sigma D}{T^{1-\frac{1}{\mathsf{p}}}}$. If strong convexity is additionally assumed, Corollary 6 gives a last-iterate rate in the order of $\frac{\log T}{T^{\mathsf{p}-1}}$.
4.2 Nonsmooth Nonconvex Optimization
This section contains another application, nonsmooth nonconvex optimization with heavy tails. Due to limited space, we will provide only the necessary background. For more details, we refer the reader to [8, 17, 20, 21, 44, 45] for recent progress. We start with a new set of conditions.
Assumption 2.
We consider the following series of assumptions:
  • The objective $F$ is lower bounded by $F_{\star}\triangleq\inf_{\boldsymbol{x}\in\mathbb{R}^{d}}F(\boldsymbol{x})\in\mathbb{R}$.
  • $F$ is differentiable and well-behaved, i.e., $F(\boldsymbol{x})-F(\boldsymbol{y})=\int_{0}^{1}\left\langle\nabla F(\boldsymbol{y}+t(\boldsymbol{x}-\boldsymbol{y})),\boldsymbol{x}-\boldsymbol{y}\right\rangle\mathrm{d}t$.
  • $F$ is $G$-Lipschitz on $\mathbb{R}^{d}$, i.e., $\left\|\nabla F(\boldsymbol{x})\right\|\leq G,\forall\boldsymbol{x}\in\mathbb{R}^{d}$.
  • Given $\boldsymbol{z}_{t}\in\mathbb{R}^{d}$ at the $t$-th iteration, one can query $\boldsymbol{g}_{t}\in\mathbb{R}^{d}$ satisfying $\mathbb{\mathbb{E}}\left[\boldsymbol{g}_{t}\mid\mathcal{F}_{t-1}\right]=\nabla F(\boldsymbol{z}_{t})$ and $\mathbb{\mathbb{E}}\left[\left\|\boldsymbol{\epsilon}_{t}\right\|^{\mathsf{p}}\right]\leq\sigma^{\mathsf{p}}$ for some $\mathsf{p}\in\left(1,2\right]$ and $\sigma\geq 0$, where $\mathcal{F}_{t}$ denotes the natural filtration and $\boldsymbol{\epsilon}_{t}\triangleq\boldsymbol{g}_{t}-\nabla F(\boldsymbol{z}_{t})$ is the stochastic noise.
Remark 7.
The second point is a mild regularity condition introduced by [7] and becomes standard in the literature [2, 23, 54]. See Definition 1 and Proposition 2 of [7] for more details. In the fourth point, we use the same notation $\boldsymbol{z}_{t}$ as in the algorithm being studied later. In fact, it can be arbitrary.
In nonsmooth nonconvex optimization, we aim to find a $(\delta,\epsilon)$-stationary point [52] (see the formal Definition 2 in Appendix E). This goal can be reduced to finding a point $\boldsymbol{x}\in\mathbb{R}^{d}$ such that $\left\|\nabla F(\boldsymbol{x})\right\|_{\delta}\leq\epsilon$, where $\left\|\nabla F(\boldsymbol{x})\right\|_{\delta}$ is a quantity introduced by [7] as follows.
Definition 1 (Definition 5 of [7]).
Given a point $\boldsymbol{x}\in\mathbb{R}^{d}$, a number $\delta>0$ and an almost-everywhere differentiable function $F$, define $\left\|\nabla F(\boldsymbol{x})\right\|_{\delta}\triangleq\inf_{S\subset\mathcal{B}(\boldsymbol{x},\delta),\frac{1}{\left|S\right|}\sum_{\boldsymbol{y}\in S}\boldsymbol{y}=\boldsymbol{x}}\left\|\frac{1}{\left|S\right|}\sum_{\boldsymbol{y}\in S}\nabla F(\boldsymbol{y})\right\|.$
The only existing sample complexity under Assumption 2 is $(G+\sigma)^{\frac{\mathsf{p}}{\mathsf{p}-1}}\delta^{-1}\epsilon^{-\frac{2\mathsf{p}-1}{\mathsf{p}-1}}$ in high probability [23], where we only report the dominant term and hide the dependency on the failure probability.
However, on the theoretical side, their result cannot recover the optimal bound $G^{2}\delta^{-1}\epsilon^{-3}$ [7] in the deterministic case. On the practical side, their method also employs the gradient clipping step, which introduces a new clipping parameter to tune. In fact, as stated in their Section 5, they observed in experiments that their algorithm without the clipping operation (exactly the algorithm we study next) still works under heavy tails. In addition, in their Section 6, they also explicitly ask whether the requirement to know $G$ and $\mathsf{A}$ can be removed.
As will be seen later, we can address these points with the new regret bounds presented before.
4.2.1 Online-to-Nonconvex Conversion under Heavy Tails
Algorithm 4 Online-to-Nonconvex Conversion ($\mathsf{O2NC}$) [7]
Input: initial point $\boldsymbol{y}_{0}\in\mathbb{R}^{d}$, $K\in\mathbb{\mathbb{N}}$, $T\in\mathbb{\mathbb{N}}$, online learning algorithm $\mathsf{A}$.
for $n=1$ to $KT$ do
 Receive $\boldsymbol{x}_{n}$ from $\mathsf{A}$
$\boldsymbol{y}_{n}=\boldsymbol{y}_{n-1}+\boldsymbol{x}_{n}$
$\boldsymbol{z}_{n}=\boldsymbol{y}_{n-1}+s_{n}\boldsymbol{x}_{n}$ where $s_{n}\sim\mathsf{Uniform}\left[0,1\right]$ i.i.d.
 Query a stochastic gradient $\boldsymbol{g}_{n}$ at $\boldsymbol{z}_{n}$
 Send $\boldsymbol{g}_{n}$ to $\mathsf{A}$
end for
Remark 8.
Note that $\mathsf{O2NC}$ is a randomized algorithm. Therefore, the definition of the natural filtration is adjusted to $\mathcal{F}_{n}\triangleq\sigma(s_{1},\boldsymbol{g}_{1},\dots,s_{n},\boldsymbol{g}_{n},s_{n+1})$ accordingly.
We provide the Online-to-Nonconvex Conversion ($\mathsf{O2NC}$) framework in Algorithm 4, which serves as a meta algorithm. Roughly speaking, Algorithm 4 reduces a nonconvex optimization problem to an OCO (in fact, OLO) problem, for which the $K$-shifting regret (see (9)) of the online learner $\mathsf{A}$ crucially affects the final convergence rate. However, the existing Theorem 8 of [7], a general convergence result for the above reduction, cannot directly apply to heavy-tailed noise, since its proof relies on the finite variance condition on $\boldsymbol{g}_{n}$ (see Appendix E for more details).
Theorem 4.
Under Assumption 2 and let $\boldsymbol{v}_{k}\triangleq-D\frac{\sum_{n=(k-1)T+1}^{kT}\nabla F(\boldsymbol{z}_{n})}{\left\|\sum_{n=(k-1)T+1}^{kT}\nabla F(\boldsymbol{z}_{n})\right\|},\forall k\in\left[K\right]$ for arbitrary $D>0$, then for any online learning algorithm $\mathsf{A}$ in $\mathsf{O2NC}$ (Algorithm 4), we have
$\mathbb{\mathbb{E}}\left[\sum_{k=1}^{K}\frac{1}{K}\left\|\frac{1}{T}\sum_{n=(k-1)T+1}^{kT}\nabla F(\boldsymbol{z}_{n})\right\|\right]\lesssim\frac{F(\boldsymbol{y}_{0})-F_{\star}}{DKT}+\frac{\mathbb{\mathbb{E}}\left[\mathsf{R}_{T}^{\mathsf{A}}(\boldsymbol{v}_{1},\dots,\boldsymbol{v}_{K})\right]}{DKT}+\frac{\sigma}{T^{1-\frac{1}{\mathsf{p}}}}.$
$\mathsf{R}_{T}^{\mathsf{A}}(\boldsymbol{v}_{1},\dots,\boldsymbol{v}_{K})$ in Theorem 4 is called $K$-shifting regret [7], defined as follows:
$\mathsf{R}_{T}^{\mathsf{A}}\left(\boldsymbol{v}_{1},\dots,\boldsymbol{v}_{K}\right)\triangleq\sum_{k=1}^{K}\sum_{n=(k-1)T+1}^{kT}\ell_{n}(\boldsymbol{x}_{n})-\ell_{n}(\boldsymbol{v}_{k})$ where $\ell_{n}(\boldsymbol{x})\triangleq\left\langle\boldsymbol{g}_{n},\boldsymbol{x}\right\rangle.$ (9)
Theorem 4 here provides a new and the first theoretical guarantee for $\mathsf{O2NC}$ under heavy tails. Especially, it recovers Theorem 8 of [7] when $\mathsf{p}=2$. A remarkable point is that the $\mathsf{O2NC}$ algorithm itself does not need any information about $\mathsf{p}$. The proof of Theorem 4 can be found in Appendix E.
4.2.2 Convergence Rates
Theorem 4 enables us to apply the results presented in Section 3. Concretely, for $\mathcal{X}=\mathcal{B}^{d}(D)$ and any $\mathsf{A}\in\left\{\mathsf{OGD},\mathsf{DA},\mathsf{AdaGrad}\right\}$, if we reset the stepsize in $\mathsf{A}$ after every $T$ iterations, there will be $\mathbb{\mathbb{E}}\left[\mathsf{R}_{T}^{\mathsf{A}}(\boldsymbol{v}_{1},\dots,\boldsymbol{v}_{K})\right]\lesssim GDK\sqrt{T}+\sigma DKT^{1/\mathsf{p}}$ by our new regret bounds, since $\boldsymbol{v}_{k}\in\mathcal{X}$. With a carefully picked $D$, we obtain the following Theorem 5. Its proof is deferred to Appendix E.
Theorem 5.
Under Assumption 2 and let $\Delta\triangleq F(\boldsymbol{y}_{0})-F_{\star}$ and $\bar{\boldsymbol{z}}_{k}\triangleq\frac{1}{T}\sum_{n=(k-1)T+1}^{kT}\boldsymbol{z}_{n},\forall k\in\left[K\right]$, setting any $\mathsf{A}\in\left\{\mathsf{OGD},\mathsf{DA},\mathsf{AdaGrad}\right\}$ in $\mathsf{O2NC}$ (Algorithm 4) with a domain $\mathcal{X}=\mathcal{B}^{d}(D)$ for $D=\delta/T$ and resetting the stepsize in $\mathsf{A}$ after every $T$ iterations, we have
$\mathbb{\mathbb{E}}\left[\frac{1}{K}\sum_{k=1}^{K}\left\|\nabla F(\bar{\boldsymbol{z}}_{k})\right\|_{\delta}\right]\lesssim\frac{\Delta}{\delta K}+\frac{G}{\sqrt{T}}+\frac{\sigma}{T^{1-\frac{1}{\mathsf{p}}}}.$
Notably, this is the first time confirming that gradient clipping is indeed unnecessary for the $\mathsf{O2NC}$ framework, matching the experimental observation of [23].
Corollary 3.
Under the same setting of Theorem 5, suppose we have $N\geq 2$ stochastic gradient budgets, taking $K=\left\lfloor N/T\right\rfloor$ and $T=\left\lceil N/2\right\rceil\land\left(\left\lceil(\delta GN/\Delta)^{\frac{2}{3}}\right\rceil\lor\left\lceil(\delta\sigma N/\Delta)^{\frac{\mathsf{p}}{2\mathsf{p}-1}}\right\rceil\right)$, we have
$\mathbb{\mathbb{E}}\left[\frac{1}{K}\sum_{k=1}^{K}\left\|\nabla F(\bar{\boldsymbol{z}}_{k})\right\|_{\delta}\right]\lesssim\frac{G}{\sqrt{N}}+\frac{\sigma}{N^{1-\frac{1}{\mathsf{p}}}}+\frac{\Delta}{\delta N}+\frac{G^{\frac{2}{3}}\Delta^{\frac{1}{3}}}{(\delta N)^{\frac{1}{3}}}+\frac{\sigma^{\frac{\mathsf{p}}{2\mathsf{p}-1}}\Delta^{\frac{\mathsf{p}-1}{2\mathsf{p}-1}}}{(\delta N)^{\frac{\mathsf{p}-1}{2\mathsf{p}-1}}}.$
Corollary 3 is obtained by optimizing $K$ and $T$ in Theorem 5. It implies a sample complexity of $G^{2}\delta^{-1}\epsilon^{-3}+\sigma^{\frac{\mathsf{p}}{\mathsf{p}-1}}\delta^{-1}\epsilon^{-\frac{2\mathsf{p}-1}{\mathsf{p}-1}}$ for finding a $(\delta,\epsilon)$-stationary point, improved over the previous bound $(G+\sigma)^{\frac{\mathsf{p}}{\mathsf{p}-1}}\delta^{-1}\epsilon^{-\frac{2\mathsf{p}-1}{\mathsf{p}-1}}$ [23]. Furthermore, leveraging the adaptive feature of $\mathsf{AdaGrad}$, Corollary 7 in Appendix E shows how to set $K$ and $T$ without $G$, $\sigma$, and $\mathsf{p}$, resulting in the first provably rate for $\mathsf{O2NC}$ when no problem information is known in advance, which solves the problem asked by [23].
4.2.3 Lower Bounds
In this part, we provide the first lower bound for nonsmooth nonconvex optimization under heavy tails in the following Theorem 6, the proof of which follows the framework first established in [3] and later developed by [7] but with some necessary (though minor) variation to make it compatible with heavy-tailed noise.
Theorem 6.
For any given $\Delta>0$, $G>0$, $\mathsf{p}\in\left(1,2\right]$, $\sigma\geq 0$, $\delta>0$ and $0<\epsilon\lesssim\frac{\Delta}{\delta}\land\frac{G^{2}\delta}{\Delta}$, there exists a dimension $d>0$ depending on the previous parameters such that, for any randomized first-order algorithm (see Definition 4 for a formal description), there exists a $G$-Lipschitz differentiable function $F:\mathbb{R}^{d}\to\mathbb{R}$ satisfying $F(\boldsymbol{0})-F_{\star}\leq\Delta$ and a function $\boldsymbol{g}:\mathbb{R}^{d}\times\left\{0,1\right\}\to\mathbb{R}$ satisfying $\mathbb{\mathbb{E}}_{r}\left[\boldsymbol{g}(\boldsymbol{x},r)\right]=\nabla F(\boldsymbol{x})$ and $\mathbb{\mathbb{E}}_{r}\left[\left\|\boldsymbol{g}(\boldsymbol{x},r)-\nabla F(\boldsymbol{x})\right\|^{\mathsf{p}}\right]\leq\sigma^{\mathsf{p}}$ where $r$ follows a certain probability distribution over $\left\{0,1\right\}$ such that the algorithm requires $\gtrsim\Delta\delta^{-1}\epsilon^{-1}+\Delta\sigma^{\frac{\mathsf{p}}{\mathsf{p}-1}}\delta^{-1}\epsilon^{-\frac{2\mathsf{p}-1}{\mathsf{p}-1}}$ queries of $\boldsymbol{g}$ to find a point $\boldsymbol{z}$ such that $\mathbb{\mathbb{E}}\left[\left\|\nabla F(\boldsymbol{z})\right\|_{\delta}\right]\leq\epsilon$.
For small enough $\epsilon$ and $\sigma>0$, Theorem 6 can be further simplified into a lower bound of $\Delta\sigma^{\frac{\mathsf{p}}{\mathsf{p}-1}}\delta^{-1}\epsilon^{-\frac{2\mathsf{p}-1}{\mathsf{p}-1}}$, matching the leading term in the sample complexity derived from the previous Corollary 3. Therefore, Theorem 6 suggests that our Corollary 3 is tight in the high accuracy and noisy regime. As such, Corollary 3 and Theorem 6 together provide a nearly complete characterization of the complexity of finding $(\delta,\epsilon)$-stationary points in the heavy-tailed setting.
However, for any general $\epsilon>0$ or the case $\sigma=0$, there is still a gap between the upper and lower bounds. Closing this gap could be an interesting direction for the future.
5 Further Extensions
So far, we revisit three classical OCO algorithms, $\mathsf{OGD}$, $\mathsf{DA}$, and $\mathsf{AdaGrad}$, show that they all guarantee optimal regrets, and provide some applications based on their new bounds. However, all the results are restricted to the Lipschitz case, which is standard in the literature, but sometimes inadequate to derive a better bound. For example, if $\ell_{t}$ is smooth (i.e., the gradient of $\ell_{t}$ is Lipschitz), then one could establish a bound depending on the cumulative competitor loss, i.e., $\sum_{t=1}^{T}\ell_{t}(\boldsymbol{x})$.
In this section, we first show that our idea presented in Section 3 can be directly extended to the smooth case, then discuss what more we can do. Formally, we need the following condition.
Condition 1.
$\ell_{t}$ is $H$-smooth on $\mathbb{R}^{d}$, i.e., $\left\|\nabla\ell_{t}(\boldsymbol{x})-\nabla\ell_{t}(\boldsymbol{y})\right\|\leq H\left\|\boldsymbol{x}-\boldsymbol{y}\right\|,\forall\boldsymbol{x},\boldsymbol{y}\in\mathbb{R}^{d}$, for all $t\in\left[T\right]$.
Remark 9.
Strictly speaking, Condition 1 may not be well-defined, since Assumption 1 only requires $\ell_{t}$ to be defined on $\mathcal{X}$. To avoid any further complications, this section recognizes $\ell_{t}$ as a real-valued convex function defined on $\mathbb{R}^{d}$, and $\mathcal{X}$ is the constraint set, i.e., the domain of the problem.
5.1 New Regrets for Old Algorithms under Smooth $\ell_{t}$ and Applications
Theorem 7.
Under Assumption 1 (with replacing the third point by Condition 1) and additionally assuming $\ell_{t}\geq 0$ on $\mathbb{R}^{d}$:
  • taking $\eta_{t}=\frac{1}{4H}\land\frac{\gamma D}{\sqrt{H}}\land\frac{D}{\sigma t^{1/\mathsf{p}}}$ for any $\gamma>0$ in $\mathsf{A}\in\left\{\mathsf{OGD},\mathsf{DA}\right\}$, we have
    $\mathbb{\mathbb{E}}\left[\mathsf{R}_{T}^{\mathsf{A}}(\boldsymbol{x})\right]\lesssim HD^{2}+\sqrt{H}D\left(\frac{1}{\gamma}+\gamma\sum_{t=1}^{T}\ell_{t}(\boldsymbol{x})\right)+\sigma DT^{1/\mathsf{p}},\forall\boldsymbol{x}\in\mathcal{X}.$
  • taking $\eta=D/\sqrt{2}$ in $\mathsf{A}=\mathsf{AdaGrad}$, we have
    $\mathbb{\mathbb{E}}\left[\mathsf{R}_{T}^{\mathsf{A}}(\boldsymbol{x})\right]\lesssim HD^{2}+D\sqrt{H\sum_{t=1}^{T}\ell_{t}(\boldsymbol{x})}+\sigma DT^{1/\mathsf{p}},\forall\boldsymbol{x}\in\mathcal{X}.$
Theorem 7 provides new regrets that extend the classical $L^{\star}$ bounds [32] to the heavy-tailed noise case, under an additional nonnegative condition $\ell_{t}\geq 0$ that is widely used in the literature [43]. The full version without such a requirement is provided in Theorem 13 in Appendix F.
As one can see, the optimal value of $\gamma$ for $\mathsf{OGD}$ and $\mathsf{DA}$ should be in the order of $1/\sqrt{\sum_{t=1}^{T}\ell_{t}(\boldsymbol{x})}$, which is however not possible to take as the competitor $\boldsymbol{x}$ is not fixed here. In contrast, $\mathsf{AdaGrad}$ contains the improved term $D\sqrt{H\sum_{t=1}^{T}\ell_{t}(\boldsymbol{x})}$, which once again suggests the benefit of adaptive methods. More importantly, the regret for $\mathsf{AdaGrad}$ also indicates that it can be oblivious to the problem class and adapt to the best-possible bound automatically, even under heavy tails.
Theorem 7 could be particularly useful for smooth convex optimization when a global minimizer is contained in $\mathcal{X}$, as given in the following corollary.
Corollary 4.
Under Assumption 1 (with replacing the third point by Condition 1) for $\ell_{t}(\boldsymbol{x})=F(\boldsymbol{x})$ and additionally assuming that $\mathcal{X}$ contains a global minimzer $\boldsymbol{x}^{\star}$ of $F$ (i.e., $\boldsymbol{x}^{\star}\in\mathrm{arginf}_{\boldsymbol{x}\in\mathbb{R}^{d}}F(\boldsymbol{x})$), let $\bar{\boldsymbol{x}}_{T}\triangleq\frac{1}{T}\sum_{t=1}^{T}\boldsymbol{x}_{t}$, taking $\eta_{t}=\frac{1}{4H}\land\frac{D}{\sigma t^{1/\mathsf{p}}}$ in $\mathsf{A}\in\left\{\mathsf{OGD},\mathsf{DA}\right\}$ or $\eta=D/\sqrt{2}$ in $\mathsf{A}=\mathsf{AdaGrad}$, we have
$\mathbb{\mathbb{E}}\left[F(\bar{\boldsymbol{x}}_{T})-F(\boldsymbol{x}^{\star})\right]\leq\frac{\mathbb{\mathbb{E}}\left[\mathsf{R}_{T}^{\mathsf{A}}(\boldsymbol{x}^{\star})\right]}{T}\lesssim\frac{HD^{2}}{T}+\frac{\sigma D}{T^{1-\frac{1}{\mathsf{p}}}}.$
Proof.
Equivalently, we can redefine $\ell_{t}(\boldsymbol{x})=F(\boldsymbol{x})-F(\boldsymbol{x}^{\star})$, then $\ell_{t}$ is nonnegative on $\mathbb{R}^{d}$. By convexity, $F(\bar{\boldsymbol{x}}_{T})-F(\boldsymbol{x}^{\star})\leq\frac{\sum_{t=1}^{T}F(\boldsymbol{x}_{t})-F(\boldsymbol{x}^{\star})}{T}=\frac{\mathsf{R}_{T}^{\mathsf{A}}(\boldsymbol{x}^{\star})}{T}$ is valid for any OCO algorithm $\mathsf{A}$. We conclude from invoking Theorem 7 with $\gamma=\frac{1}{4D\sqrt{H}}$ for $\mathsf{A}\in\left\{\mathsf{OGD},\mathsf{DA}\right\}$ and $\eta=D/\sqrt{2}$ for $\mathsf{A}=\mathsf{AdaGrad}$ and combining the fact $\ell_{t}(\boldsymbol{x}^{\star})=0$. ∎
An important implication of Corollary 4 is the first to show that $\mathsf{SGD}$ can converge at the optimal rate of $HD^{2}/T+\sigma D/T^{1-1/\mathsf{p}}$ for heavy-tailed smooth convex optimization (though under an extra condition). In fact, a valid but worse bound without assuming the existence of $\boldsymbol{x}^{\star}\in\mathcal{X}$ is possible, which we defer to Corollary 8 in Appendix F.
5.2 Optimistic Algorithms for Broader Cases
As presented, we have successfully handled smooth OCO with heavy tails for classical $\mathsf{OGD}$, $\mathsf{DA}$, and $\mathsf{AdaGrad}$. In fact, our ideas can also be applied to another famous family of methods known as optimistic algorithms [6, 18, 36] to deal with broader cases. Specifically, we will study an optimistic version of $\mathsf{AdaGrad}$ called $\mathsf{OAdaGrad}$ in Appendix F to address various cases simultaneously (see Theorems 14 and 15) and provide more applications for stochastic optimization in Appendix G.
For example, in the special case of heavy-tailed Hölder smooth optimization (i.e., $\left\|\nabla F(\boldsymbol{x})-\nabla F(\boldsymbol{y})\right\|\leq H\left\|\boldsymbol{x}-\boldsymbol{y}\right\|^{\nu}$ for some $H>0$ and $\nu\in\left(0,1\right]$), Corollary 9 provides a rate $\frac{HD^{1+\nu}}{T^{\frac{1+\nu}{2}}}+\frac{\sigma D}{T^{1-\frac{1}{\mathsf{p}}}}$ for convex problems, Corollaries 12 and 13 establish provable results for nonconvex problems, where the problem-dependent parameters may be known or unknown. All of these results are new to the best of our knowledge.
6 Conclusion and Future Work
This paper shows that three classical OCO algorithms, $\mathsf{OGD}$, $\mathsf{DA}$, and $\mathsf{AdaGrad}$, can achieve the optimal in-expectation regret under heavy tails without any algorithmic modification if the feasible set is bounded, and provides some applications in stochastic optimization. The main limitation of our work is that all the proof crucially relies on the bounded domain assumption, which may not always be suitable in practice. Finding a weaker sufficient condition, under which the classical OCO algorithms work with heavy tails provably, is a direction worth studying in the future.
Appendix A Missing Proofs for Online Gradient Descent
This section provides missing proofs for regret bounds of $\mathsf{OGD}$. Before showing the formal proof, we recall the following core inequality that holds for any $\boldsymbol{x}\in\mathcal{X}$ given in (6):
$\left\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t}-\boldsymbol{x}\right\rangle\leq\frac{\left\|\boldsymbol{x}_{t}-\boldsymbol{x}\right\|^{2}-\left\|\boldsymbol{x}_{t+1}-\boldsymbol{x}\right\|^{2}}{2\eta_{t}}+\eta_{t}G^{2}+\mathsf{C}(\mathsf{p})\eta_{t}^{\mathsf{p}-1}\left\|\boldsymbol{\epsilon}_{t}\right\|^{\mathsf{p}}D^{2-\mathsf{p}}.$ (10)
The key to establishing the above result is showing
$\left\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t}-\boldsymbol{x}_{t+1}\right\rangle-\frac{\left\|\boldsymbol{x}_{t}-\boldsymbol{x}_{t+1}\right\|^{2}}{2\eta_{t}}\leq\eta_{t}G^{2}+\mathsf{C}(\mathsf{p})\eta_{t}^{\mathsf{p}-1}\left\|\boldsymbol{\epsilon}_{t}\right\|^{\mathsf{p}}D^{2-\mathsf{p}},$ (11)
the proof of which is by combining (3), (4), and (5) established in the main text.
A.1 Proof of Theorem 1
Proof.
For any $\boldsymbol{x}\in\mathcal{X}$, sum up (10) from $t=1$ to $T$ and drop the term $\frac{\left\|\boldsymbol{x}_{T+1}-\boldsymbol{x}\right\|^{2}}{2\eta_{T}}$ to obtain
$\sum_{t=1}^{T}\left\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t}-\boldsymbol{x}\right\rangle$
$\leq$ $\frac{\left\|\boldsymbol{x}_{1}-\boldsymbol{x}\right\|^{2}}{2\eta_{1}}+\sum_{t=1}^{T-1}\left(\frac{1}{\eta_{t+1}}-\frac{1}{\eta_{t}}\right)\frac{\left\|\boldsymbol{x}_{t+1}-\boldsymbol{x}\right\|^{2}}{2}+\sum_{t=1}^{T}\eta_{t}G^{2}+\mathsf{C}(\mathsf{p})\eta_{t}^{\mathsf{p}-1}\left\|\boldsymbol{\epsilon}_{t}\right\|^{\mathsf{p}}D^{2-\mathsf{p}}$ (12)
$\leq$ $\frac{D^{2}}{\eta_{T}}+\sum_{t=1}^{T}\eta_{t}G^{2}+\mathsf{C}(\mathsf{p})\eta_{t}^{\mathsf{p}-1}\left\|\boldsymbol{\epsilon}_{t}\right\|^{\mathsf{p}}D^{2-\mathsf{p}},$ (13)
where the last step is due to $\left\|\boldsymbol{x}_{t}-\boldsymbol{x}\right\|\leq D,\forall t\in\left[T\right]$ and $\eta_{t+1}\leq\eta_{t},\forall t\in\left[T-1\right]$.
Taking expectations on both sides of (13) yields that
$\mathbb{\mathbb{E}}\left[\mathsf{R}_{T}^{\mathsf{OGD}}(\boldsymbol{x})\right]\leq\frac{D^{2}}{\eta_{T}}+\sum_{t=1}^{T}\eta_{t}G^{2}+\mathsf{C}(\mathsf{p})\eta_{t}^{\mathsf{p}-1}\sigma^{\mathsf{p}}D^{2-\mathsf{p}},$ (14)
where for the L.H.S., we use $\mathbb{\mathbb{E}}\left[\left\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t}-\boldsymbol{x}\right\rangle\right]=\mathbb{\mathbb{E}}\left[\mathbb{\mathbb{E}}\left[\left\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t}-\boldsymbol{x}\right\rangle\mid\mathcal{F}_{t-1}\right]\right]$ and
$\mathbb{\mathbb{E}}\left[\left\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t}-\boldsymbol{x}\right\rangle\mid\mathcal{F}_{t-1}\right]=\left\langle\mathbb{\mathbb{E}}\left[\boldsymbol{g}_{t}\mid\mathcal{F}_{t-1}\right],\boldsymbol{x}_{t}-\boldsymbol{x}\right\rangle=\left\langle\nabla\ell_{t}(\boldsymbol{x}_{t}),\boldsymbol{x}_{t}-\boldsymbol{x}\right\rangle\geq\ell_{t}(\boldsymbol{x}_{t})-\ell_{t}(\boldsymbol{x}),$ (15)
for the R.H.S., we use $\mathbb{\mathbb{E}}\left[\left\|\boldsymbol{\epsilon}_{t}\right\|^{\mathsf{p}}\right]\leq\sigma^{\mathsf{p}}$.
Finally, we plug $\eta_{t}=\frac{D}{G\sqrt{t}}\land\frac{D}{\sigma t^{1/\mathsf{p}}},\forall t\in\left[T\right]$ into (14), then use $\sum_{t=1}^{T}\frac{1}{\sqrt{t}}\lesssim\sqrt{T}$ and $\sum_{t=1}^{T}\frac{1}{t^{1-1/\mathsf{p}}}\lesssim T^{1/\mathsf{p}}$ to conclude
$\mathbb{\mathbb{E}}\left[\mathsf{R}_{T}^{\mathsf{OGD}}(\boldsymbol{x})\right]\lesssim GD\sqrt{T}+\sigma DT^{1/\mathsf{p}}.$
A.2 Extension to Online Strongly Convex Optimization
Next, we extend Theorem 1 to the strongly convex case, i.e., $\exists\mu>0$ such that for all $t\in\left[T\right]$,
$\frac{\mu}{2}\left\|\boldsymbol{x}-\boldsymbol{y}\right\|^{2}+\left\langle\nabla\ell_{t}(\boldsymbol{y}),\boldsymbol{x}-\boldsymbol{y}\right\rangle+\ell_{t}(\boldsymbol{y})\leq\ell_{t}(\boldsymbol{x}),\forall\boldsymbol{x},\boldsymbol{y}\in\mathcal{X},\nabla\ell_{t}(\boldsymbol{y})\in\partial\ell_{t}(\boldsymbol{y}).$ (16)
In this setting, it is well known that $\mathsf{OGD}$ achieves a logarithmic regret bound when $\mathsf{p}=2$ [12, 32]. Theorem 8 below provides the first provable result for $\mathsf{p}<2$.
Theorem 8.
Under Assumption 1 and additionally assuming (16), taking $\eta_{t}=\frac{1}{\mu t}$ in $\mathsf{OGD}$ (Algorithm 1), we have
$\mathbb{\mathbb{E}}\left[\mathsf{R}_{T}^{\mathsf{OGD}}(\boldsymbol{x})\right]\lesssim\frac{G^{2}\left(1+\log T\right)}{\mu}+\frac{\sigma^{\mathsf{p}}D^{2-\mathsf{p}}}{\mu^{\mathsf{p}-1}}\times\begin{cases}T^{2-\mathsf{p}}&\mathsf{p}\in\left(1,2\right)\\ 1+\log T&\mathsf{p}=2\end{cases},\forall\boldsymbol{x}\in\mathcal{X}.$
Theorem 8 shows that under strongly convexity, $\mathsf{OGD}$ for $\mathsf{p}\in\left(1,2\right)$ achieves a better sublinear regret $T^{2-\mathsf{p}}$ than $T^{1/\mathsf{p}}$ in Theorem 1 as $2-\mathsf{p}\leq 1/\mathsf{p},\forall\mathsf{p}>0$. One point we highlight here is that the stepsize $\eta_{t}=\frac{1}{\mu t}$ is commonly used in the OCO literature and is independent of the tail index $\mathsf{p}$.
However, in contrast to Theorem 1, we suspect Theorem 8 is not tight in $T$ for $\mathsf{p}\in\left(1,2\right)$. The reason is that for nonsmooth strongly convex optimization with heavy tails (i.e., $\ell_{t}=F,\forall t\in\left[T\right]$ where $F$ is strongly convex), Theorem 8 can convert to a convergence rate only in the order of $1/T^{\mathsf{p}-1}$, which is worse than the lower bound $1/T^{2-2/\mathsf{p}}$ [51]. Therefore, we conjecture that a way to obtain a better regret bound than $T^{2-\mathsf{p}}$ exists, which we leave as future work.
Proof of Theorem 8.
For any $\boldsymbol{x}\in\mathcal{X}$, we take expectations on both sides of (12) to have
$\mathbb{\mathbb{E}}\left[\mathsf{R}_{T}^{\mathsf{OGD}}(\boldsymbol{x})\right]\leq$ $\left(\frac{1}{\eta_{1}}-\mu\right)\frac{\left\|\boldsymbol{x}_{1}-\boldsymbol{x}\right\|^{2}}{2}+\sum_{t=1}^{T-1}\left(\frac{1}{\eta_{t+1}}-\frac{1}{\eta_{t}}-\mu\right)\frac{\mathbb{\mathbb{E}}\left[\left\|\boldsymbol{x}_{t+1}-\boldsymbol{x}\right\|^{2}\right]}{2}$
$+\sum_{t=1}^{T}\eta_{t}G^{2}+\mathsf{C}(\mathsf{p})\eta_{t}^{\mathsf{p}-1}\sigma^{\mathsf{p}}D^{2-\mathsf{p}},$ (17)
where for the L.H.S., we follow a similar step of reasoning out (15) but instead using
$\left\langle\nabla\ell_{t}(\boldsymbol{x}_{t}),\boldsymbol{x}_{t}-\boldsymbol{x}\right\rangle\geq\ell_{t}(\boldsymbol{x}_{t})-\ell_{t}(\boldsymbol{x})+\frac{\mu}{2}\left\|\boldsymbol{x}_{t}-\boldsymbol{x}\right\|^{2},$
for the R.H.S., we use $\mathbb{\mathbb{E}}\left[\left\|\boldsymbol{\epsilon}_{t}\right\|^{\mathsf{p}}\right]\leq\sigma^{\mathsf{p}}$.
Next, we plug $\eta_{t}=\frac{1}{\mu t},\forall t\in\left[T\right]$ into (17) to obtain
$\mathbb{\mathbb{E}}\left[\mathsf{R}_{T}^{\mathsf{OGD}}(\boldsymbol{x})\right]\lesssim\sum_{t=1}^{T}\frac{G^{2}}{\mu t}+\frac{\sigma^{\mathsf{p}}D^{2-\mathsf{p}}}{\mu^{\mathsf{p}-1}t^{\mathsf{p}-1}}\lesssim\frac{G^{2}\left(1+\log T\right)}{\mu}+\frac{\sigma^{\mathsf{p}}D^{2-\mathsf{p}}}{\mu^{\mathsf{p}-1}}\times\begin{cases}T^{2-\mathsf{p}}&\mathsf{p}\in\left(1,2\right)\\ 1+\log T&\mathsf{p}=2\end{cases}.$
Appendix B Missing Proofs for Dual Averaging
This section provides missing proofs for regret bounds of $\mathsf{DA}$.
B.1 Proof of Theorem 2
Proof.
Let $L_{t}(\boldsymbol{x})\triangleq\frac{\left\|\boldsymbol{x}-\boldsymbol{x}_{1}\right\|^{2}}{2\eta_{t-1}}+\sum_{s=1}^{t-1}\left\langle\boldsymbol{g}_{s},\boldsymbol{x}\right\rangle,\forall t\in\left[T+1\right]$, where $\eta_{0}\triangleq\eta_{1}$. Then $\mathsf{DA}$ can be equivalently written as
$\boldsymbol{x}_{t}=\mathbb{\mathrm{argmin}}_{\boldsymbol{x}\in\mathcal{X}}L_{t}(\boldsymbol{x}),\forall t\in\left[T+1\right].$
By Lemma 7.1 of [32], for any $\boldsymbol{x}\in\mathcal{X}$,
$\sum_{t=1}^{T}\left\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t}-\boldsymbol{x}\right\rangle$ $=\frac{\left\|\boldsymbol{x}-\boldsymbol{x}_{1}\right\|^{2}}{2\eta_{T}}+L_{T+1}(\boldsymbol{x}_{T+1})-L_{T+1}(\boldsymbol{x})+\sum_{t=1}^{T}L_{t}(\boldsymbol{x}_{t})+\left\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t}\right\rangle-L_{t+1}(\boldsymbol{x}_{t+1})$
$\leq\frac{\left\|\boldsymbol{x}-\boldsymbol{x}_{1}\right\|^{2}}{2\eta_{T}}+\sum_{t=1}^{T}L_{t}(\boldsymbol{x}_{t})-L_{t+1}(\boldsymbol{x}_{t+1})+\left\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t}\right\rangle,$
where the inequality holds by $L_{T+1}(\boldsymbol{x}_{T+1})\leq L_{T+1}(\boldsymbol{x}),\forall\boldsymbol{x}\in\mathcal{X}$ due to $\boldsymbol{x}_{T+1}=\mathbb{\mathrm{argmin}}_{\boldsymbol{x}\in\mathcal{X}}L_{T+1}(\boldsymbol{x})$. Note that for any $t\in\left[T\right]$,
$L_{t}(\boldsymbol{x}_{t})-L_{t+1}(\boldsymbol{x}_{t+1})+\left\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t}\right\rangle$
$=$ $L_{t}(\boldsymbol{x}_{t})-L_{t}(\boldsymbol{x}_{t+1})+\left\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t}-\boldsymbol{x}_{t+1}\right\rangle+\frac{\left\|\boldsymbol{x}_{t+1}-\boldsymbol{x}_{1}\right\|^{2}}{2\eta_{t-1}}-\frac{\left\|\boldsymbol{x}_{t+1}-\boldsymbol{x}_{1}\right\|^{2}}{2\eta_{t}}$
$\overset{(a)}{\leq}$ $L_{t}(\boldsymbol{x}_{t})-L_{t}(\boldsymbol{x}_{t+1})+\left\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t}-\boldsymbol{x}_{t+1}\right\rangle$
$\overset{(b)}{\leq}$ $\left\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t}-\boldsymbol{x}_{t+1}\right\rangle-\frac{\left\|\boldsymbol{x}_{t}-\boldsymbol{x}_{t+1}\right\|^{2}}{2\eta_{t-1}},$
where $(a)$ is by $\eta_{t}\leq\eta_{t-1},\forall t\in\left[T\right]$ and $(b)$ holds because $L_{t}$ is $\frac{1}{\eta_{t-1}}$-strongly convex and $\boldsymbol{x}_{t}=\mathbb{\mathrm{argmin}}_{\boldsymbol{x}\in\mathcal{X}}L_{t}(\boldsymbol{x})$, which together imply
$L_{t}(\boldsymbol{x}_{t})-L_{t}(\boldsymbol{x}_{t+1})\leq\left\langle\nabla L_{t}(\boldsymbol{x}_{t}),\boldsymbol{x}_{t}-\boldsymbol{x}_{t+1}\right\rangle-\frac{\left\|\boldsymbol{x}_{t}-\boldsymbol{x}_{t+1}\right\|^{2}}{2\eta_{t-1}}\leq-\frac{\left\|\boldsymbol{x}_{t}-\boldsymbol{x}_{t+1}\right\|^{2}}{2\eta_{t-1}}.$
Therefore, we have
$\sum_{t=1}^{T}\left\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t}-\boldsymbol{x}\right\rangle\leq\frac{\left\|\boldsymbol{x}-\boldsymbol{x}_{1}\right\|^{2}}{2\eta_{T}}+\sum_{t=1}^{T}\left\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t}-\boldsymbol{x}_{t+1}\right\rangle-\frac{\left\|\boldsymbol{x}_{t}-\boldsymbol{x}_{t+1}\right\|^{2}}{2\eta_{t-1}}.$ (18)
By the same argument as proving (11) but replacing $\eta_{t}$ with $\eta_{t-1}$, there is
$\left\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t}-\boldsymbol{x}_{t+1}\right\rangle-\frac{\left\|\boldsymbol{x}_{t}-\boldsymbol{x}_{t+1}\right\|^{2}}{2\eta_{t-1}}\leq\eta_{t-1}G^{2}+\mathsf{C}(\mathsf{p})\eta_{t-1}^{\mathsf{p}-1}\left\|\boldsymbol{\epsilon}_{t}\right\|^{\mathsf{p}}D^{2-\mathsf{p}}.$
As such, we know
$\sum_{t=1}^{T}\left\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t}-\boldsymbol{x}\right\rangle\leq\frac{\left\|\boldsymbol{x}-\boldsymbol{x}_{1}\right\|^{2}}{2\eta_{T}}+\sum_{t=1}^{T}\eta_{t-1}G^{2}+\mathsf{C}(\mathsf{p})\eta_{t-1}^{\mathsf{p}-1}\left\|\boldsymbol{\epsilon}_{t}\right\|^{\mathsf{p}}D^{2-\mathsf{p}}.$ (19)
Finally, following similar steps in proving Theorem 1 in Appendix A, we conclude
$\mathbb{\mathbb{E}}\left[\mathsf{R}_{T}^{\mathsf{DA}}(\boldsymbol{x})\right]\lesssim GD\sqrt{T}+\sigma DT^{1/\mathsf{p}}.$
B.2 Dual Averaging with an Adaptive Stepsize
We show that $\mathsf{DA}$ with an adaptive stepsize can also achieve the optimal regret $GD\sqrt{T}+\sigma DT^{1/\mathsf{p}}$.
Theorem 9.
Under Assumption 1, taking $\eta_{t}=2DV_{t}^{-1/2}$ and $V_{t}=\sum_{s=1}^{t}\left\|\boldsymbol{g}_{s}\right\|^{2}$ in $\mathsf{DA}$ (Algorithm 2), we have
$\mathbb{\mathbb{E}}\left[\mathsf{R}_{T}^{\mathsf{DA}}(\boldsymbol{x})\right]\lesssim GD\sqrt{T}+\sigma DT^{1/\mathsf{p}},\forall\boldsymbol{x}\in\mathcal{X}.$
Proof.
For any $\boldsymbol{x}\in\mathcal{X}$, we have
[…]
where $\eta_{0}\triangleq\eta_{1}$. On the one hand, we can use AM-GM inequality to bound
$\left\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t}-\boldsymbol{x}_{t+1}\right\rangle-\frac{\left\|\boldsymbol{x}_{t}-\boldsymbol{x}_{t+1}\right\|^{2}}{2\eta_{t-1}}\leq\frac{\eta_{t-1}\left\|\boldsymbol{g}_{t}\right\|^{2}}{2}.$
On the other hand, we know
$\left\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t}-\boldsymbol{x}_{t+1}\right\rangle-\frac{\left\|\boldsymbol{x}_{t}-\boldsymbol{x}_{t+1}\right\|^{2}}{2\eta_{t-1}}\leq\left\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t}-\boldsymbol{x}_{t+1}\right\rangle\leq\left\|\boldsymbol{g}_{t}\right\|\left\|\boldsymbol{x}_{t}-\boldsymbol{x}_{t+1}\right\|\leq\left\|\boldsymbol{g}_{t}\right\|D,$ (21)
where the second step is by Cauchy-Schwarz inequality. Therefore, for any $t\geq 2$,
$\left\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t}-\boldsymbol{x}_{t+1}\right\rangle-\frac{\left\|\boldsymbol{x}_{t}-\boldsymbol{x}_{t+1}\right\|^{2}}{2\eta_{t-1}}$ $\leq\frac{\eta_{t-1}\left\|\boldsymbol{g}_{t}\right\|^{2}}{2}\land\left\|\boldsymbol{g}_{t}\right\|D\overset{(a)}{\leq}\frac{2}{\frac{2}{\eta_{t-1}\left\|\boldsymbol{g}_{t}\right\|^{2}}+\frac{1}{\left\|\boldsymbol{g}_{t}\right\|D}}$
$\overset{(b)}{=}\frac{2D\left\|\boldsymbol{g}_{t}\right\|^{2}}{\sqrt{\sum_{s=1}^{t-1}\left\|\boldsymbol{g}_{s}\right\|^{2}}+\left\|\boldsymbol{g}_{t}\right\|}\overset{(c)}{\leq}\frac{2D\left\|\boldsymbol{g}_{t}\right\|^{2}}{\sqrt{\sum_{s=1}^{t}\left\|\boldsymbol{g}_{s}\right\|^{2}}},$ (22)
where $(a)$ is due to $x\land y\leq\frac{2}{x^{-1}+y^{-1}},\forall x,y>0$, $(b)$ is by $\eta_{t-1}=\frac{2D}{\sqrt{\sum_{s=1}^{t-1}\left\|\boldsymbol{g}_{s}\right\|^{2}}}$, and $(c)$ holds because of $\sqrt{\sum_{s=1}^{t}\left\|\boldsymbol{g}_{s}\right\|^{2}}\leq\sqrt{\sum_{s=1}^{t-1}\left\|\boldsymbol{g}_{s}\right\|^{2}}+\left\|\boldsymbol{g}_{t}\right\|$. Note that (22) is also true for $t=1$ by (21).
Combine (20) and (22) and use $\left\|\boldsymbol{x}-\boldsymbol{x}_{1}\right\|\leq D$ to obtain
$\sum_{t=1}^{T}\left\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t}-\boldsymbol{x}\right\rangle\leq\frac{D^{2}}{2\eta_{T}}+\sum_{t=1}^{T}\frac{2D\left\|\boldsymbol{g}_{t}\right\|^{2}}{\sqrt{\sum_{s=1}^{t}\left\|\boldsymbol{g}_{s}\right\|^{2}}}=\frac{D^{2}}{2\eta_{T}}+\sum_{t=1}^{T}\eta_{t}\left\|\boldsymbol{g}_{t}\right\|^{2},$
which only differs from (23) by a constant. Hence, by a similar proof for (25), there is
$\sum_{t=1}^{T}\left\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t}-\boldsymbol{x}\right\rangle\lesssim D\left[\sqrt{\sum_{t=1}^{T}\left\|\nabla\ell_{t}(\boldsymbol{x}_{t})\right\|^{2}}+\left(\sum_{t=1}^{T}\left\|\boldsymbol{\epsilon}_{t}\right\|^{\mathsf{p}}\right)^{\frac{1}{\mathsf{p}}}\right],$
implying
$\mathbb{\mathbb{E}}\left[\mathsf{R}_{T}^{\mathsf{DA}}(\boldsymbol{x})\right]\lesssim GD\sqrt{T}+\sigma DT^{1/\mathsf{p}}.$
Appendix C Missing Proofs for $\mathsf{AdaGrad}$
This section provides missing proofs for regret bounds of $\mathsf{AdaGrad}$.
C.1 Proof of Theorem 3
Proof.
As mentioned, $\mathsf{AdaGrad}$ can be viewed as $\mathsf{OGD}$ with a stepsize $\eta_{t}=\frac{\eta}{\sqrt{V_{t}}}=\frac{\eta}{\sqrt{\sum_{s=1}^{t}\left\|\boldsymbol{g}_{s}\right\|^{2}}}$. Therefore, we can use (1) for $\mathsf{AdaGrad}$ to know for any $\boldsymbol{x}\in\mathcal{X}$,
$\left\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t}-\boldsymbol{x}\right\rangle\leq\frac{\left\|\boldsymbol{x}_{t}-\boldsymbol{x}\right\|^{2}-\left\|\boldsymbol{x}_{t+1}-\boldsymbol{x}\right\|^{2}}{2\eta_{t}}+\frac{\eta_{t}\left\|\boldsymbol{g}_{t}\right\|^{2}}{2}.$
Sum up the above inequality from $t=1$ to $T$ and drop the term $\frac{\left\|\boldsymbol{x}_{T+1}-\boldsymbol{x}\right\|^{2}}{2\eta_{T}}$ to have
$\sum_{t=1}^{T}\left\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t}-\boldsymbol{x}\right\rangle$ $\leq\frac{\left\|\boldsymbol{x}_{1}-\boldsymbol{x}\right\|^{2}}{2\eta_{1}}+\sum_{t=1}^{T-1}\left(\frac{1}{\eta_{t+1}}-\frac{1}{\eta_{t}}\right)\frac{\left\|\boldsymbol{x}_{t+1}-\boldsymbol{x}\right\|^{2}}{2}+\sum_{t=1}^{T}\frac{\eta_{t}\left\|\boldsymbol{g}_{t}\right\|^{2}}{2}$
$\leq\frac{D^{2}}{2\eta_{T}}+\sum_{t=1}^{T}\frac{\eta_{t}\left\|\boldsymbol{g}_{t}\right\|^{2}}{2},$ (23)
where the last step is by $\left\|\boldsymbol{x}_{t}-\boldsymbol{x}\right\|\leq D,\forall t\in\left[T\right]$ and $\eta_{t+1}\leq\eta_{t},\forall t\in\left[T-1\right]$.
Next, observe that for any $t\in\left[T\right]$,
$\left\|\boldsymbol{g}_{t}\right\|^{2}=\frac{\eta^{2}}{\eta_{t}^{2}}-\frac{\eta^{2}}{\eta_{t-1}^{2}}=\eta^{2}\left(\frac{1}{\eta_{t}}-\frac{1}{\eta_{t-1}}\right)\left(\frac{1}{\eta_{t}}+\frac{1}{\eta_{t-1}}\right)\leq\frac{2\eta^{2}}{\eta_{t}}\left(\frac{1}{\eta_{t}}-\frac{1}{\eta_{t-1}}\right),$
where $1/\eta_{0}$ should be read as 0. The above inequality implies
$\sum_{t=1}^{T}\frac{\eta_{t}\left\|\boldsymbol{g}_{t}\right\|^{2}}{2}\leq\eta^{2}\sum_{t=1}^{T}\frac{1}{\eta_{t}}-\frac{1}{\eta_{t-1}}=\frac{\eta^{2}}{\eta_{T}}.$ (24)
Combine (23) and (24) to have
$\sum_{t=1}^{T}\left\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t}-\boldsymbol{x}\right\rangle\leq\frac{D^{2}}{2\eta_{T}}+\frac{\eta^{2}}{\eta_{T}}=\left(\frac{D^{2}}{2\eta}+\eta\right)\sqrt{\sum_{t=1}^{T}\left\|\boldsymbol{g}_{t}\right\|^{2}}.$
Note that there is
$\sqrt{\sum_{t=1}^{T}\left\|\boldsymbol{g}_{t}\right\|^{2}}$ $\leq\sqrt{\sum_{t=1}^{T}2\left\|\nabla\ell_{t}(\boldsymbol{x}_{t})\right\|^{2}+2\left\|\boldsymbol{\epsilon}_{t}\right\|^{2}}\leq\sqrt{2\sum_{t=1}^{T}\left\|\nabla\ell_{t}(\boldsymbol{x}_{t})\right\|^{2}}+\sqrt{2\sum_{t=1}^{T}\left\|\boldsymbol{\epsilon}_{t}\right\|^{2}}$
$\leq\sqrt{2\sum_{t=1}^{T}\left\|\nabla\ell_{t}(\boldsymbol{x}_{t})\right\|^{2}}+\sqrt{2}\left(\sum_{t=1}^{T}\left\|\boldsymbol{\epsilon}_{t}\right\|^{\mathsf{p}}\right)^{\frac{1}{\mathsf{p}}},$
where the last step is due to $\left\|\cdot\right\|_{2}\leq\left\|\cdot\right\|_{\mathsf{p}}$ for any $\mathsf{p}\in\left[1,2\right]$. Hence, we obtain
$\sum_{t=1}^{T}\left\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t}-\boldsymbol{x}\right\rangle\leq\sqrt{2}\left(\frac{D^{2}}{2\eta}+\eta\right)\left[\sqrt{\sum_{t=1}^{T}\left\|\nabla\ell_{t}(\boldsymbol{x}_{t})\right\|^{2}}+\left(\sum_{t=1}^{T}\left\|\boldsymbol{\epsilon}_{t}\right\|^{\mathsf{p}}\right)^{\frac{1}{\mathsf{p}}}\right].$ (25)
We take expectations on both sides of (25), then apply Hölder’s inequality to have
$\mathbb{\mathbb{E}}\left[\left(\sum_{t=1}^{T}\left\|\boldsymbol{\epsilon}_{t}\right\|^{\mathsf{p}}\right)^{\frac{1}{\mathsf{p}}}\right]\leq\left(\sum_{t=1}^{T}\mathbb{\mathbb{E}}\left[\left\|\boldsymbol{\epsilon}_{t}\right\|^{\mathsf{p}}\right]\right)^{\frac{1}{\mathsf{p}}}\leq\sigma T^{\frac{1}{\mathsf{p}}},$
and finally plug in $\eta=D/\sqrt{2}$ to conclude
$\mathbb{\mathbb{E}}\left[\mathsf{R}_{T}^{\mathsf{AdaGrad}}(\boldsymbol{x})\right]\lesssim GD\sqrt{T}+\sigma DT^{1/\mathsf{p}}.$
Appendix D Missing Proofs for Applications: Nonsmooth Convex Optimization
In this section, we prove the last-iterate convergence for $\mathsf{SGD}$ (i.e., $\mathsf{OGD}$ for stochastic optimization) under heavy-tailed noise. Following the existing literature [22, 25], we will move beyond the standard $G$-Lipschitz convex function considered before and instead focus on a broader class of nonsmooth functions, i.e., $\exists\mu\geq 0$, $G\geq 0$, and $H\geq 0$, such that $G+H>0$ and
$\frac{\mu}{2}\left\|\boldsymbol{x}-\boldsymbol{y}\right\|^{2}\leq\ell_{t}(\boldsymbol{x})-\ell_{t}(\boldsymbol{y})-\left\langle\nabla\ell_{t}(\boldsymbol{y}),\boldsymbol{x}-\boldsymbol{y}\right\rangle\leq 2G\left\|\boldsymbol{x}-\boldsymbol{y}\right\|+\frac{H}{2}\left\|\boldsymbol{x}-\boldsymbol{y}\right\|^{2},\forall\boldsymbol{x},\boldsymbol{y}\in\mathcal{X},\nabla\ell_{t}(\boldsymbol{y})\in\partial\ell_{t}(\boldsymbol{y}).$ (26)
Notably, such a condition encompasses various commonly studied objectives simultaneously (i.e., strong convexity and smoothness).
Theorem 10 first shows the general last-iterate convergence result, the proof of which is inspired by [25, 50].
Theorem 10.
Under Assumption 1 (without the third point) and additionally assuming (26) for $\ell_{t}(\boldsymbol{x})=F(\boldsymbol{x})$, for any stepsize $0<\eta_{t}\leq\frac{1}{2H\lor\mu}$ in $\mathsf{OGD}$ (Algorithm 1), let $\gamma_{t}\triangleq\frac{\eta_{t}}{\prod_{s=2}^{t}(1-\mu\eta_{s})}$, we have
$\mathbb{\mathbb{E}}\left[F(\boldsymbol{x}_{T})-F(\boldsymbol{x})\right]\lesssim\frac{(1-\mu\eta_{1})D^{2}}{\sum_{t=1}^{T-1}\gamma_{t}}+G^{2}\sum_{t=1}^{T-1}\frac{\gamma_{t}\eta_{t}}{\sum_{s=t}^{T-1}\gamma_{s}}+\sigma^{\mathsf{p}}D^{2-\mathsf{p}}\sum_{t=1}^{T-1}\frac{\gamma_{t}\eta_{t}^{\mathsf{p}-1}}{\sum_{s=t}^{T-1}\gamma_{s}},\forall\boldsymbol{x}\in\mathcal{X}.$
Proof.
Given $\boldsymbol{x}\in\mathcal{X}$, we recursively define
$\boldsymbol{y}_{0}\triangleq\boldsymbol{x}$ and $\boldsymbol{y}_{t}\triangleq\left(1-\frac{w_{t-1}}{w_{t}}\right)\boldsymbol{x}_{t}+\frac{w_{t-1}}{w_{t}}\boldsymbol{y}_{t-1},\forall t\in\left[T\right],$ (27)
in which
$w_{t}\triangleq\frac{\gamma_{T}}{\sum_{s=t}^{T}\gamma_{s}},\forall t\in\left[T\right]$ and $w_{0}\triangleq w_{1}.$ (28)
Equivalently, $\boldsymbol{y}_{t}$ can be written into a convex combination of $\boldsymbol{x},\boldsymbol{x}_{1},\dots,\boldsymbol{x}_{t}$ as
$\boldsymbol{y}_{t}=\frac{w_{0}}{w_{t}}\boldsymbol{x}+\sum_{s=1}^{t}\frac{w_{s}-w_{s-1}}{w_{t}}\boldsymbol{x}_{s},\forall t\in\left\{0\right\}\cup\left[T\right].$ (29)
Therefore, $\boldsymbol{y}_{t}$ also falls into $\mathcal{X}$ and satisfies $\boldsymbol{y}_{t}\in\mathcal{F}_{t-1}$.
We invoke (2) for $\boldsymbol{y}_{t}$ to obtain
$\left\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t}-\boldsymbol{y}_{t}\right\rangle\leq\frac{\left\|\boldsymbol{x}_{t}-\boldsymbol{y}_{t}\right\|^{2}-\left\|\boldsymbol{x}_{t+1}-\boldsymbol{y}_{t}\right\|^{2}}{2\eta_{t}}+\left\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t}-\boldsymbol{x}_{t+1}\right\rangle-\frac{\left\|\boldsymbol{x}_{t}-\boldsymbol{x}_{t+1}\right\|^{2}}{2\eta_{t}},$
and then bound
$\left\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t}-\boldsymbol{x}_{t+1}\right\rangle=\left\langle\nabla F(\boldsymbol{x}_{t}),\boldsymbol{x}_{t}-\boldsymbol{x}_{t+1}\right\rangle+\left\langle\boldsymbol{\epsilon}_{t},\boldsymbol{x}_{t}-\boldsymbol{x}_{t+1}\right\rangle$
$\leq$ $F(\boldsymbol{x}_{t})-F(\boldsymbol{x}_{t+1})+8\eta_{t}G^{2}+2^{\mathsf{p}-1}\mathsf{C}(\mathsf{p})\eta_{t}^{\mathsf{p}-1}\left\|\boldsymbol{\epsilon}_{t}\right\|^{\mathsf{p}}D^{2-\mathsf{p}}+\frac{\left\|\boldsymbol{x}_{t}-\boldsymbol{x}_{t+1}\right\|^{2}}{2\eta_{t}},$
where, in the second step, we use (5) to bound $\left\langle\boldsymbol{\epsilon}_{t},\boldsymbol{x}_{t}-\boldsymbol{x}_{t+1}\right\rangle$ (up to change $\eta_{t}$ to $2\eta_{t}$) and the following step to bound $\left\langle\nabla F(\boldsymbol{x}_{t}),\boldsymbol{x}_{t}-\boldsymbol{x}_{t+1}\right\rangle$,
[…]
$\leq F(\boldsymbol{x}_{t})-F(\boldsymbol{x}_{t+1})+8\eta_{t}G^{2}+\frac{\left\|\boldsymbol{x}_{t}-\boldsymbol{x}_{t+1}\right\|^{2}}{8\eta_{t}}+\frac{H}{2}\left\|\boldsymbol{x}_{t}-\boldsymbol{x}_{t+1}\right\|^{2}$
$\overset{\eta_{t}\leq\frac{1}{2H}}{\leq}F(\boldsymbol{x}_{t})-F(\boldsymbol{x}_{t+1})+8\eta_{t}G^{2}+\frac{3\left\|\boldsymbol{x}_{t}-\boldsymbol{x}_{t+1}\right\|^{2}}{8\eta_{t}}.$
Hence, we know
$\left\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t}-\boldsymbol{y}_{t}\right\rangle\leq F(\boldsymbol{x}_{t})-F(\boldsymbol{x}_{t+1})+\frac{\left\|\boldsymbol{x}_{t}-\boldsymbol{y}_{t}\right\|^{2}-\left\|\boldsymbol{x}_{t+1}-\boldsymbol{y}_{t}\right\|^{2}}{2\eta_{t}}+8\eta_{t}G^{2}+2^{\mathsf{p}-1}\mathsf{C}(\mathsf{p})\eta_{t}^{\mathsf{p}-1}\left\|\boldsymbol{\epsilon}_{t}\right\|^{\mathsf{p}}D^{2-\mathsf{p}}.$ (30)
Since $\boldsymbol{x}_{t},\boldsymbol{y}_{t}\in\mathcal{F}_{t-1}$, there is
[…]
As such, we can take expectations on both sides of (30) and rearrange terms to have
$\mathbb{\mathbb{E}}\left[F(\boldsymbol{x}_{t+1})-F(\boldsymbol{y}_{t})\right]\leq\frac{(1-\mu\eta_{t})\mathbb{\mathbb{E}}\left[\left\|\boldsymbol{x}_{t}-\boldsymbol{y}_{t}\right\|^{2}\right]-\mathbb{\mathbb{E}}\left[\left\|\boldsymbol{x}_{t+1}-\boldsymbol{y}_{t}\right\|^{2}\right]}{2\eta_{t}}+8\eta_{t}G^{2}+2^{\mathsf{p}-1}\mathsf{C}(\mathsf{p})\eta_{t}^{\mathsf{p}-1}\left\|\boldsymbol{\epsilon}_{t}\right\|^{\mathsf{p}}D^{2-\mathsf{p}}$
$\leq$ $\frac{(1-\mu\eta_{t})\mathbb{\mathbb{E}}\left[\frac{w_{t-1}}{w_{t}}\left\|\boldsymbol{x}_{t}-\boldsymbol{y}_{t-1}\right\|^{2}\right]-\mathbb{\mathbb{E}}\left[\left\|\boldsymbol{x}_{t+1}-\boldsymbol{y}_{t}\right\|^{2}\right]}{2\eta_{t}}+8\eta_{t}G^{2}+2^{\mathsf{p}-1}\mathsf{C}(\mathsf{p})\eta_{t}^{\mathsf{p}-1}\left\|\boldsymbol{\epsilon}_{t}\right\|^{\mathsf{p}}D^{2-\mathsf{p}},$ (31)
where the second step is due to $\left\|\boldsymbol{x}_{t}-\boldsymbol{y}_{t}\right\|^{2}\leq\left(1-\frac{w_{t-1}}{w_{t}}\right)\left\|\boldsymbol{x}_{t}-\boldsymbol{x}_{t}\right\|^{2}+\frac{w_{t-1}}{w_{t}}\left\|\boldsymbol{x}_{t}-\boldsymbol{y}_{t-1}\right\|^{2}=\frac{w_{t-1}}{w_{t}}\left\|\boldsymbol{x}_{t}-\boldsymbol{y}_{t-1}\right\|^{2}$ by (27) and the convexity of $\left\|\boldsymbol{x}_{t}-\cdot\right\|^{2}$. Multiply both sides of (31) by $w_{t}\gamma_{t}$ and sum up from $t=1$ to $T$ to obtain (note that $\frac{\gamma_{t}}{\eta_{t}}=\frac{1}{\prod_{s=2}^{t}(1-\mu\eta_{s})}=\frac{\gamma_{t+1}(1-\mu\eta_{t+1})}{\eta_{t+1}},\forall t\in\left[T-1\right]$)
$\mathbb{\mathbb{E}}\left[\sum_{t=1}^{T}w_{t}\gamma_{t}\left(F(\boldsymbol{x}_{t+1})-F(\boldsymbol{y}_{t})\right)\right]$
$\leq$ $\frac{\gamma_{1}(1-\mu\eta_{1})w_{0}\left\|\boldsymbol{x}_{1}-\boldsymbol{y}_{0}\right\|^{2}}{2\eta_{1}}-\frac{\gamma_{T}\mathbb{\mathbb{E}}\left[w_{T}\left\|\boldsymbol{x}_{T+1}-\boldsymbol{y}_{T}\right\|^{2}\right]}{2\eta_{T}}+\sum_{t=1}^{T}8w_{t}\gamma_{t}\eta_{t}G^{2}+2^{\mathsf{p}-1}\mathsf{C}(\mathsf{p})w_{t}\gamma_{t}\eta_{t}^{\mathsf{p}-1}\sigma^{\mathsf{p}}D^{2-\mathsf{p}}$
$\leq$ $\frac{(1-\mu\eta_{1})w_{0}D^{2}}{2}+\sum_{t=1}^{T}8w_{t}\gamma_{t}\eta_{t}G^{2}+2^{\mathsf{p}-1}\mathsf{C}(\mathsf{p})w_{t}\gamma_{t}\eta_{t}^{\mathsf{p}-1}\sigma^{\mathsf{p}}D^{2-\mathsf{p}}.$ (32)
Now observe that
[…]
which implies that
$\sum_{t=1}^{T}w_{t}\gamma_{t}\left(F(\boldsymbol{y}_{t})-F(\boldsymbol{x})\right)$ $\leq\sum_{t=1}^{T}\sum_{s=1}^{t}\left(w_{s}-w_{s-1}\right)\gamma_{t}\left(F(\boldsymbol{x}_{s})-F(\boldsymbol{x})\right)$
$=\sum_{t=1}^{T}\left(w_{t}-w_{t-1}\right)\left(\sum_{s=t}^{T}\gamma_{s}\right)\left(F(\boldsymbol{x}_{t})-F(\boldsymbol{x})\right).$
Thus, we can lower bound the L.H.S. of (32) by
$\sum_{t=1}^{T}w_{t}\gamma_{t}\left(F(\boldsymbol{x}_{t+1})-F(\boldsymbol{y}_{t})\right)$ $=\sum_{t=1}^{T}w_{t}\gamma_{t}\left(F(\boldsymbol{x}_{t+1})-F(\boldsymbol{x})\right)-w_{t}\gamma_{t}\left(F(\boldsymbol{y}_{t})-F(\boldsymbol{x})\right)$
$\geq w_{T}\gamma_{T}\left(F(\boldsymbol{x}_{T+1})-F(\boldsymbol{x})\right)-\left(w_{1}-w_{0}\right)\left(\sum_{s=1}^{T}\gamma_{1}\right)\left(F(\boldsymbol{x}_{1})-F(\boldsymbol{x})\right)$
$+\sum_{t=2}^{T}\left[w_{t-1}\gamma_{t-1}-\left(w_{t}-w_{t-1}\right)\left(\sum_{s=t}^{T}\gamma_{s}\right)\right]\left(F(\boldsymbol{x}_{t})-F(\boldsymbol{x})\right)$
$=w_{T}\gamma_{T}\left(F(\boldsymbol{x}_{T+1})-F(\boldsymbol{x})\right),$ (33)
where the last step is due to, for $t\geq 2$,
[…]
$=\frac{\gamma_{T}}{\sum_{s=t-1}^{T}\gamma_{s}}\cdot\gamma_{t-1}-\frac{\gamma_{T}}{\sum_{s=t-1}^{T}\gamma_{s}}\cdot\gamma_{t-1}=0,$
[…]
We plug (33) back into (32) and divide both sides by $w_{T}\gamma_{T}$ to obtain
$\mathbb{\mathbb{E}}\left[F(\boldsymbol{x}_{T+1})-F(\boldsymbol{x})\right]$ $\leq\frac{(1-\mu\eta_{1})w_{0}D^{2}}{2w_{T}\gamma_{T}}+\sum_{t=1}^{T}\frac{8w_{t}\gamma_{t}\eta_{t}}{w_{T}\gamma_{T}}G^{2}+\frac{2^{\mathsf{p}-1}\mathsf{C}(\mathsf{p})w_{t}\gamma_{t}\eta_{t}^{\mathsf{p}-1}}{w_{T}\gamma_{T}}\sigma^{\mathsf{p}}D^{2-\mathsf{p}}$
[…]
Finally, relabel $T$ by T1T-1 to complete the proof. ∎
Equipped with Theorem 10, we show the following last-iterate convergence rate for $\mathsf{SGD}$$\mathsf{OGD}$. As far as we know, this is the first and the only provable result demonstrating that the last iterate of $\mathsf{SGD}$ can converge in heavy-tailed stochastic optimization without gradient clipping. When $T$ is unknown, we only incur an extra logarithmic factor compared to the best possible rate. In particular, the first stepsize degenerates to $\eta_{t}=\frac{D}{G\sqrt{t}}\land\frac{D}{\sigma t^{1/\mathsf{p}}}$ employed in Theorem 1 if $H=0$. When $T$ is known, we employ the linear decay rate proposed by [50] to give an optimal rate.
Corollary 5.
Under Assumption 1 (without the third point) and additionally assuming (26) (with $\mu=0$) for $\ell_{t}(\boldsymbol{x})=F(\boldsymbol{x})$:
  • taking $\eta_{t}=\frac{1}{2H}\land\frac{D}{G\sqrt{t}}\land\frac{D}{\sigma t^{1/\mathsf{p}}}$ in $\mathsf{OGD}$ (Algorithm 1), we have
    $\mathbb{\mathbb{E}}\left[F(\boldsymbol{x}_{T})-F(\boldsymbol{x})\right]\lesssim\frac{HD^{2}\left(1+\log T\right)}{T}+\frac{GD\left(1+\log T\right)}{\sqrt{T}}+\frac{\sigma D\left(1+\log T\right)}{T^{1-\frac{1}{\mathsf{p}}}},\forall\boldsymbol{x}\in\mathcal{X}.$
  • taking $\eta_{t}=\frac{T-t}{2HT}\land\frac{D(T-t)}{GT^{3/2}}\land\frac{D(T-t)}{\sigma T^{1+1/\mathsf{p}}}$ in $\mathsf{OGD}$ (Algorithm 1), we have
    $\mathbb{\mathbb{E}}\left[F(\boldsymbol{x}_{T})-F(\boldsymbol{x})\right]\lesssim\frac{HD^{2}}{T}+\frac{GD}{\sqrt{T}}+\frac{\sigma D}{T^{1-\frac{1}{\mathsf{p}}}},\forall\boldsymbol{x}\in\mathcal{X}.$
Proof.
Since both kinds of stepsize satisfy $\eta_{t}\leq\frac{1}{2H\lor\mu}\overset{\mu=0}{=}\frac{1}{2H}$, by Theorem 10, we have
$\mathbb{\mathbb{E}}\left[F(\boldsymbol{x}_{T})-F(\boldsymbol{x})\right]$ $\lesssim\frac{(1-\mu\eta_{1})D^{2}}{\sum_{t=1}^{T-1}\gamma_{t}}+G^{2}\sum_{t=1}^{T-1}\frac{\gamma_{t}\eta_{t}}{\sum_{s=t}^{T-1}\gamma_{s}}+\sigma^{\mathsf{p}}D^{2-\mathsf{p}}\sum_{t=1}^{T-1}\frac{\gamma_{t}\eta_{t}^{\mathsf{p}-1}}{\sum_{s=t}^{T-1}\gamma_{s}}$
$=\frac{D^{2}}{\sum_{t=1}^{T-1}\eta_{t}}+G^{2}\sum_{t=1}^{T-1}\frac{\eta_{t}^{2}}{\sum_{s=t}^{T-1}\eta_{s}}+\sigma^{\mathsf{p}}D^{2-\mathsf{p}}\sum_{t=1}^{T-1}\frac{\eta_{t}^{\mathsf{p}}}{\sum_{s=t}^{T-1}\eta_{s}},$ (34)
where the second step is due $\gamma_{t}=\frac{\eta_{t}}{\prod_{s=2}^{t}(1-\mu\eta_{s})}=\eta_{t}$ when $\mu=0$.
For $\eta_{t}=\frac{1}{2H}\land\frac{D}{G\sqrt{t}}\land\frac{D}{\sigma t^{1/\mathsf{p}}}$, we observe that by Cauchy-Schwarz inequality
$(T-t)^{2}\leq\left(\sum_{s=t}^{T-1}\frac{1}{\eta_{s}}\right)\left(\sum_{s=t}^{T-1}\eta_{s}\right)\Rightarrow\frac{1}{\sum_{s=t}^{T-1}\eta_{s}}\leq\frac{\sum_{s=t}^{T-1}\frac{1}{\eta_{s}}}{(T-t)^{2}}.$
Thus, there is
[…]
We first bound
$\sum_{t=1}^{T-1}\frac{1}{\eta_{t}}=\sum_{t=1}^{T-1}2H\lor\frac{G\sqrt{t}}{D}\lor\frac{\sigma t^{1/\mathsf{p}}}{D}\leq\sum_{t=1}^{T-1}2H+\frac{G\sqrt{t}}{D}+\frac{\sigma t^{1/\mathsf{p}}}{D}\lesssim HT+\frac{G}{D}T^{3/2}+\frac{\sigma}{D}T^{1+1/\mathsf{p}},$
which implies
$\frac{D^{2}}{(T-1)^{2}}\sum_{t=1}^{T-1}\frac{1}{\eta_{t}}\lesssim\frac{HD^{2}}{T}+\frac{GD}{\sqrt{T}}+\frac{\sigma D}{T^{1-\frac{1}{\mathsf{p}}}}.$ (36)
Next, we know
$\sum_{t=1}^{T-1}\frac{\eta_{t}^{2}\sum_{s=t}^{T-1}\frac{1}{\eta_{s}}}{(T-t)^{2}}$ $\overset{(a)}{\leq}\sum_{t=1}^{T-1}\left[\frac{2HD^{2}}{G^{2}}\cdot\frac{1}{t(T-t)}+\frac{D}{G}\cdot\frac{\sum_{s=t}^{T-1}\sqrt{s}}{t(T-t)^{2}}+\frac{\sigma D}{G^{2}}\cdot\frac{\sum_{s=t}^{T-1}s^{1/\mathsf{p}}}{t(T-t)^{2}}\right]$
[…]
where $(a)$ is by $\eta_{t}\leq\frac{D}{G\sqrt{t}}$ and $\frac{1}{\eta_{s}}\leq 2H+\frac{G\sqrt{s}}{D}+\frac{\sigma s^{1/\mathsf{p}}}{D}$. Hence, there is
$G^{2}\sum_{t=1}^{T-1}\frac{\eta_{t}^{2}\sum_{s=t}^{T-1}\frac{1}{\eta_{s}}}{(T-t)^{2}}\lesssim\frac{HD^{2}\left(1+\log T\right)}{T}+\frac{GD\left(1+\log T\right)}{\sqrt{T}}+\frac{\sigma D\left(1+\log T\right)}{T^{1-\frac{1}{\mathsf{p}}}}.$ (37)
Similarly, we can bound
$\sigma^{\mathsf{p}}D^{2-\mathsf{p}}\sum_{t=1}^{T-1}\frac{\eta_{t}^{\mathsf{p}}\sum_{s=t}^{T-1}\frac{1}{\eta_{s}}}{(T-t)^{2}}\lesssim\frac{HD^{2}\left(1+\log T\right)}{T}+\frac{GD\left(1+\log T\right)}{\sqrt{T}}+\frac{\sigma D\left(1+\log T\right)}{T^{1-\frac{1}{\mathsf{p}}}}.$ (38)
Finally, we plug (36), (37) and (38) back into (35) to conclude.
For $\eta_{t}=\frac{T-t}{2HT}\land\frac{D(T-t)}{GT^{3/2}}\land\frac{D(T-t)}{\sigma T^{1+1/\mathsf{p}}}$, we denote by $\eta_{t}=\eta(T-t)$ for $\eta\triangleq\frac{1}{2HT}\land\frac{D}{GT^{3/2}}\land\frac{D}{\sigma T^{1+1/\mathsf{p}}}$ to have
[…]
$=\frac{2D^{2}}{\eta T(T-1)}+2\eta G^{2}\sum_{t=1}^{T-1}\frac{T-t}{T-t+1}+2\eta^{\mathsf{p}-1}\sigma^{\mathsf{p}}D^{2-\mathsf{p}}\sum_{t=1}^{T-1}\frac{(T-t)^{\mathsf{p}-1}}{T-t+1}$
$\lesssim\frac{D^{2}}{\eta T^{2}}+\eta G^{2}T+\eta^{\mathsf{p}-1}\sigma^{\mathsf{p}}D^{2-\mathsf{p}}T^{\mathsf{p}-1}\lesssim\frac{HD^{2}}{T}+\frac{GD}{\sqrt{T}}+\frac{\sigma D}{T^{1-\frac{1}{\mathsf{p}}}}.$
Next, Corollary 6 provides the first last-iterate convergence results in the strongly convex case. In particular, when $H=0\Rightarrow\kappa=0$, the first stepsize degenerates to $\eta_{t}=\frac{1}{\mu t}$, as used in Theorem 8, and yields a last-iterate rate $\frac{G^{2}\left(1+\log T\right)}{\mu T}+\frac{\sigma^{\mathsf{p}}D^{2-\mathsf{p}}\left(1+\log T\right)}{\mu^{\mathsf{p}-1}T^{\mathsf{p}-1}}$, matching the average-iterate rate (up to an extra logarithmic factor) implied by the online-to-batch conversion for Theorem 8.
Corollary 6.
Under Assumption 1 (without the third point) and additionally assuming (26) (with $\mu>0$) for $\ell_{t}(\boldsymbol{x})=F(\boldsymbol{x})$, let $\kappa\triangleq\frac{H}{\mu}$:
  • taking $\eta_{t}=\frac{1}{\mu(t+2\kappa)}$ in $\mathsf{OGD}$ (Algorithm 1), we have
    $\mathbb{\mathbb{E}}\left[F(\boldsymbol{x}_{T})-F(\boldsymbol{x})\right]\lesssim\frac{HD^{2}}{T}+\frac{G^{2}\left(1+\log T\right)}{\mu\left(T+\kappa\right)}+\frac{\sigma^{\mathsf{p}}D^{2-\mathsf{p}}\left(1+\log T\right)}{\mu^{\mathsf{p}-1}\left(T+\kappa\right)^{\mathsf{p}-1}},\forall\boldsymbol{x}\in\mathcal{X}.$
  • taking $\eta_{t}=\begin{cases}\frac{1}{\mu(1+2\kappa)}&t\leq\tau\\ \frac{2}{\mu(t-\tau+2+4\kappa)}&t\geq\tau+1\end{cases}$ where $\tau\triangleq\left\lceil\frac{T}{2}\right\rceil$ in $\mathsf{OGD}$ (Algorithm 1), we have
    $\mathbb{\mathbb{E}}\left[F(\boldsymbol{x}_{T})-F(\boldsymbol{x})\right]\lesssim\frac{HD^{2}}{\exp\left(\frac{T}{2+4\kappa}\right)}+\frac{G^{2}\left(1+\log T\right)}{\mu\left(T+\kappa\right)}+\frac{\sigma^{\mathsf{p}}D^{2-\mathsf{p}}\left(1+\log T\right)}{\mu^{\mathsf{p}-1}\left(T+\kappa\right)^{\mathsf{p}-1}},\forall\boldsymbol{x}\in\mathcal{X}.$
Proof.
Since both kinds of stepsize satisfy $\eta_{t}\leq\frac{1}{2H\lor\mu}$, by Theorem 10, we have
$\mathbb{\mathbb{E}}\left[F(\boldsymbol{x}_{T})-F(\boldsymbol{x})\right]\lesssim\frac{(1-\mu\eta_{1})D^{2}}{\sum_{t=1}^{T-1}\gamma_{t}}+G^{2}\sum_{t=1}^{T-1}\frac{\gamma_{t}\eta_{t}}{\sum_{s=t}^{T-1}\gamma_{s}}+\sigma^{\mathsf{p}}D^{2-\mathsf{p}}\sum_{t=1}^{T-1}\frac{\gamma_{t}\eta_{t}^{\mathsf{p}-1}}{\sum_{s=t}^{T-1}\gamma_{s}}.$ (39)
For $\eta_{t}=\frac{1}{\mu(t+2\kappa)}$, we can find
$\gamma_{t}=\frac{\eta_{t}}{\prod_{s=2}^{t}(1-\mu\eta_{s})}=\frac{1}{\mu(t+2\kappa)\prod_{s=2}^{t}\frac{s-1+2\kappa}{s+2\kappa}}=\frac{1}{\mu(1+2\kappa)}=\eta_{1}.$ (40)
Hence, there is
[…]
$=\frac{2HD^{2}}{T-1}+\frac{G^{2}}{\mu}\sum_{t=1}^{T-1}\frac{1}{(T-t)(t+2\kappa)}+\frac{\sigma^{\mathsf{p}}D^{2-\mathsf{p}}}{\mu^{\mathsf{p}-1}}\sum_{t=1}^{T-1}\frac{1}{(T-t)(t+2\kappa)^{\mathsf{p}-1}}.$ (41)
Given $a\in\left(0,1\right]$, we can bound
$\sum_{t=1}^{T-1}\frac{1}{(T-t)(t+2\kappa)^{a}}=\frac{1}{T+2\kappa}\sum_{t=1}^{T-1}\frac{(t+2\kappa)^{1-a}}{T-t}+\frac{1}{(t+2\kappa)^{a}}\lesssim\frac{1+\log T}{(T+2\kappa)^{a}},$
where the last step is due to
$\sum_{t=1}^{T-1}\frac{(t+2\kappa)^{1-a}}{T-t}\lesssim(T+2\kappa)^{1-a}(1+\log T)$ and $\sum_{t=1}^{T-1}\frac{1}{(t+2\kappa)^{a}}\lesssim\begin{cases}T^{1-a}&a\in\left(0,1\right)\\ 1+\log T&a=1\end{cases}.$
Finally, we apply the above inequality to (41) with $a=1$ and $a=\mathsf{p}-1$ to obtain
$\mathbb{\mathbb{E}}\left[F(\boldsymbol{x}_{T})-F(\boldsymbol{x})\right]\lesssim\frac{HD^{2}}{T}+\frac{G^{2}\left(1+\log T\right)}{\mu\left(T+\kappa\right)}+\frac{\sigma^{\mathsf{p}}D^{2-\mathsf{p}}\left(1+\log T\right)}{\mu^{\mathsf{p}-1}\left(T+\kappa\right)^{\mathsf{p}-1}}.$
For $\eta_{t}=\begin{cases}\frac{1}{\mu(1+2\kappa)}&t\leq\tau\\ \frac{2}{\mu(t-\tau+2+4\kappa)}&t\geq\tau+1\end{cases}$, let $\eta\triangleq\frac{1}{\mu(1+2\kappa)}$, we can find
$\gamma_{t}=\frac{\eta_{t}}{\prod_{s=2}^{t}(1-\mu\eta_{s})}=\begin{cases}\frac{\eta}{(1-\mu\eta)^{t-1}}&t\leq\tau\\ \frac{\eta(t-\tau+1+4\kappa)}{(1-\mu\eta)^{\tau-1}(1+4\kappa)}&t\geq\tau+1\end{cases}.$ (42)
Hence, there is
$\frac{1-\mu\eta_{1}}{\sum_{t=1}^{T-1}\gamma_{t}}$ $\leq\frac{1-\mu\eta_{1}}{\sum_{t=1}^{\tau}\gamma_{t}}=\frac{1-\mu\eta}{\sum_{t=1}^{\tau}\frac{\eta}{(1-\mu\eta)^{t-1}}}=\frac{(1-\mu\eta)^{\tau}}{\eta\sum_{t=1}^{\tau}(1-\mu\eta)^{\tau-t}}=\frac{\mu(1-\mu\eta)^{\tau}}{1-(1-\mu\eta)^{\tau}}$
$\overset{(1-\mu\eta)^{\tau}\leq 1-\mu\eta}{\leq}\frac{(1-\mu\eta)^{\tau}}{\eta}=2H(1-\mu\eta)^{\tau-1}\leq 2H\exp\left(-\mu\eta(\tau-1)\right)$
$=2H\exp\left(\frac{-\tau+1}{1+2\kappa}\right)\overset{\kappa\geq 0}{\leq}2eH\exp\left(\frac{-\tau}{1+2\kappa}\right)\overset{\tau\geq\frac{T}{2}}{\lesssim}\frac{H}{\exp\left(\frac{T}{2+4\kappa}\right)},$
implying that
[…]
Given $a\in\left(0,1\right]$, we can write
$\sum_{t=1}^{T-1}\frac{\gamma_{t}\eta_{t}^{a}}{\sum_{s=t}^{T-1}\gamma_{s}}=\sum_{t=1}^{\tau}\frac{\gamma_{t}\eta_{t}^{a}}{\sum_{s=t}^{T-1}\gamma_{s}}+\sum_{t=\tau+1}^{T-1}\frac{\gamma_{t}\eta_{t}^{a}}{\sum_{s=t}^{T-1}\gamma_{s}}=\sum_{t=1}^{\tau}\frac{\gamma_{t}\eta^{a}}{\sum_{s=t}^{T-1}\gamma_{s}}+\sum_{t=\tau+1}^{T-1}\frac{\gamma_{t}\eta_{t}^{a}}{\sum_{s=t}^{T-1}\gamma_{s}}.$ (44)
When $t\leq\tau$, we know
[…]
which implies
$\sum_{t=1}^{\tau}\frac{\gamma_{t}\eta^{a}}{\sum_{s=t}^{T-1}\gamma_{s}}\leq\frac{2(1+4\kappa)\tau}{\mu^{a}(1+2\kappa)^{a}(T-\tau)(T-\tau+1+8\kappa)}\lesssim\frac{(1+\kappa)^{1-a}}{\mu^{a}(T+\kappa)}\leq\frac{1}{\mu^{a}(T+\kappa)^{a}}.$ (45)
When $t\geq\tau+1$, we know
[…]
$=\frac{2}{2T-2\tau+1+8\kappa}\left(\frac{t-\tau+1+4\kappa}{T-t}+\frac{t-\tau+1+4\kappa}{T+t-2\tau+1+8\kappa}\right)$
$\overset{\tau\leq\frac{T+1}{2},t\geq\tau+1}{\lesssim}\frac{1}{T+\kappa}\left(\frac{t-\tau+1+4\kappa}{T-t}+\frac{t-\tau+1+4\kappa}{T-\tau+2+8\kappa}\right),$
which implies
$\sum_{t=\tau+1}^{T-1}\frac{\gamma_{t}\eta_{t}^{a}}{\sum_{s=t}^{T-1}\gamma_{s}}$ $\lesssim\frac{1}{\mu^{a}(T+\kappa)}\sum_{t=\tau+1}^{T-1}\frac{(t-\tau+1+4\kappa)^{1-a}}{T-t}+\frac{(t-\tau+1+4\kappa)^{1-a}}{T-\tau+2+8\kappa}$
$\lesssim\frac{1}{\mu^{a}(T+\kappa)}\left((T+\kappa)^{1-a}(1+\log T)+(T+\kappa)^{1-a}\right)\lesssim\frac{1+\log T}{\mu^{a}(T+\kappa)^{a}}.$ (46)
Plug (45) and (46) back into (44) to have
$\sum_{t=1}^{T-1}\frac{\gamma_{t}\eta_{t}^{a}}{\sum_{s=t}^{T-1}\gamma_{s}}\lesssim\frac{1+\log T}{\mu^{a}(T+\kappa)^{a}}.$
Applying the above inequality (for $a=1$ and $a=\mathsf{p}-1$) to (43), we conclude
$\mathbb{\mathbb{E}}\left[F(\boldsymbol{x}_{T})-F(\boldsymbol{x})\right]\lesssim\frac{HD^{2}}{\exp\left(\frac{T}{2+4\kappa}\right)}+\frac{G^{2}\left(1+\log T\right)}{\mu\left(T+\kappa\right)}+\frac{\sigma^{\mathsf{p}}D^{2-\mathsf{p}}\left(1+\log T\right)}{\mu^{\mathsf{p}-1}\left(T+\kappa\right)^{\mathsf{p}-1}}.$
Appendix E Missing Proofs for Applications: Nonsmooth Nonconvex Optimization
E.1 $(\delta,\epsilon)$-Stationary Points
Definition 2 (Definition 4 of [7]).
A point $\boldsymbol{x}\in\mathbb{R}^{d}$ is a $(\delta,\epsilon)$-stationary point of an almost-everywhere differentiable function $F$ if there is a finite subset $S\subset\mathcal{B}^{d}(\boldsymbol{x},\delta)$ such that for $\boldsymbol{y}$ selected uniformly at random from $S$, $\mathbb{\mathbb{E}}\left[\boldsymbol{y}\right]=\boldsymbol{x}$ and $\left\|\mathbb{\mathbb{E}}\left[\nabla F(\boldsymbol{y})\right]\right\|\leq\epsilon$.
The concept of the $(\delta,\epsilon)$-stationary point presented here is due to [7], which is mildly more stringent than the notion of [52], since the latter does not require $\mathbb{\mathbb{E}}\left[\boldsymbol{y}\right]=\boldsymbol{x}$. For more discussions, see Section 2.1 of [7].
E.2 Proof of Theorem 4
In this section, our ultimate goal is to prove Theorem 4 for the $\mathsf{O2NC}$ algorithm, extending Theorem 8 of [7] from $\mathsf{p}=2$ to any $\mathsf{p}\in\left(1,2\right]$. Notably, our new result does not require any modification to the $\mathsf{O2NC}$ method, but is obtained only from a more careful analysis, indicating that $\mathsf{O2NC}$ is a robust and powerful algorithmic framework.
We begin with Lemma 2, which lies as the cornerstone for establishing the convergence of $\mathsf{O2NC}$.
Lemma 2 (Theorem 7 of [7]).
Under Assumption 2 (only need the second point and the unbiased part in the fourth point), for any sequence of vectors $\boldsymbol{u}_{1},\dots,\boldsymbol{u}_{KT}\in\mathbb{R}^{d}$, $\mathsf{O2NC}$ (Algorithm 4) guarantees
$\mathbb{\mathbb{E}}\left[F(\boldsymbol{y}_{KT})\right]=F(\boldsymbol{y}_{0})+\mathbb{\mathbb{E}}\left[\sum_{n=1}^{KT}\left\langle\boldsymbol{g}_{n},\boldsymbol{x}_{n}-\boldsymbol{u}_{n}\right\rangle\right]+\mathbb{\mathbb{E}}\left[\sum_{n=1}^{KT}\left\langle\boldsymbol{g}_{n},\boldsymbol{u}_{n}\right\rangle\right].$ (47)
To relate Lemma 2 to the concept of $K$-shifting regret introduced before (see (9)), suppose now a sequence of vectors $\boldsymbol{v}_{1},\dots,\boldsymbol{v}_{K}$ is given, if we set $\boldsymbol{u}_{n}=\boldsymbol{v}_{k}$ for all $n\in\left\{(k-1)T+1,\dots,kT\right\}$ and $k\in\left[K\right]$, then the second term on the R.H.S. of (47) can be written as $\mathbb{\mathbb{E}}\left[\mathsf{R}_{T}^{\mathsf{A}}(\boldsymbol{v}_{1},\dots,\boldsymbol{v}_{K})\right]$, and the third term can be simplified into $\sum_{k=1}^{K}\mathbb{\mathbb{E}}\left[\left\langle\sum_{n=(k-1)T+1}^{kT}\boldsymbol{g}_{n},\boldsymbol{v}_{k}\right\rangle\right]$.
Same as [7], we pick $\boldsymbol{v}_{k}\triangleq-D\frac{\sum_{n=(k-1)T+1}^{kT}\nabla F(\boldsymbol{z}_{n})}{\left\|\sum_{n=(k-1)T+1}^{kT}\nabla F(\boldsymbol{z}_{n})\right\|}$ for some constant $D>0$, which gives us
$\mathbb{\mathbb{E}}\left[\left\langle\sum_{n=(k-1)T+1}^{kT}\boldsymbol{g}_{n},\boldsymbol{v}_{k}\right\rangle\right]$ $=\mathbb{\mathbb{E}}\left[\left\langle\sum_{n=(k-1)T+1}^{kT}\boldsymbol{\epsilon}_{n},\boldsymbol{v}_{k}\right\rangle\right]-D\mathbb{\mathbb{E}}\left[\left\|\sum_{n=(k-1)T+1}^{kT}\nabla F(\boldsymbol{z}_{n})\right\|\right]$
$\leq D\mathbb{\mathbb{E}}\left[\left\|\sum_{n=(k-1)T+1}^{kT}\boldsymbol{\epsilon}_{n}\right\|\right]-D\mathbb{\mathbb{E}}\left[\left\|\sum_{n=(k-1)T+1}^{kT}\nabla F(\boldsymbol{z}_{n})\right\|\right].$
If $\boldsymbol{\epsilon}_{n}$ has a finite variance (i.e., $\mathsf{p}=2$), then like [7], one can invoke Hölder’s inequality and use the fact $\mathbb{\mathbb{E}}\left[\left\langle\boldsymbol{\epsilon}_{m},\boldsymbol{\epsilon}_{n}\right\rangle\right]=0,\forall m\neq n\in\left[KT\right]$ to obtain for any $k\in\left[K\right]$,
$\mathbb{\mathbb{E}}\left[\left\|\sum_{n=(k-1)T+1}^{kT}\boldsymbol{\epsilon}_{n}\right\|\right]\leq\sqrt{\mathbb{\mathbb{E}}\left[\left\|\sum_{n=(k-1)T+1}^{kT}\boldsymbol{\epsilon}_{n}\right\|^{2}\right]}=\sqrt{\sum_{n=(k-1)T+1}^{kT}\mathbb{\mathbb{E}}\left[\left\|\boldsymbol{\epsilon}_{n}\right\|^{2}\right]}\leq\sigma\sqrt{T}.$
However, this argument immediately fails when $\mathsf{p}<2$ as $\mathbb{\mathbb{E}}\left[\left\|\boldsymbol{\epsilon}_{n}\right\|^{2}\right]$ can be $\infty$. To handle this potential issue, we require the following Lemma 3.
Lemma 3 (Lemma 4.3 of [26]).
Given a vector-valued martingale difference sequence $\boldsymbol{w}_{1},\dots,\boldsymbol{w}_{T}$, there is
$\mathbb{\mathbb{E}}\left[\left\|\sum_{t=1}^{T}\boldsymbol{w}_{t}\right\|\right]\leq 2\sqrt{2}\mathbb{\mathbb{E}}\left[\left(\sum_{t=1}^{T}\left\|\boldsymbol{w}_{t}\right\|^{\mathsf{p}}\right)^{\frac{1}{\mathsf{p}}}\right],\forall\mathsf{p}\in\left[1,2\right].$
Equipped with Lemmas 2 and 3, we are ready to formally prove Theorem 4, demonstrating that the $\mathsf{O2NC}$ framework provably works under heavy-tailed noise.
Proof of Theorem 4.
We invoke Lemma 2 with $\boldsymbol{u}_{n}=\boldsymbol{v}_{\left\lceil n/T\right\rceil},\forall n\in\left[KT\right]$ (equivalently, $\boldsymbol{u}_{n}=\boldsymbol{v}_{k}$ if $n\in\left\{(k-1)T+1,\dots,kT\right\}$) and use the definition of $K$-shifting regret (see (9)) to obtain
$\mathbb{\mathbb{E}}\left[F(\boldsymbol{y}_{KT})\right]=F(\boldsymbol{y}_{0})+\mathbb{\mathbb{E}}\left[\mathsf{R}_{T}^{\mathsf{A}}\left(\boldsymbol{v}_{1},\dots,\boldsymbol{v}_{K}\right)\right]+\sum_{k=1}^{K}\mathbb{\mathbb{E}}\left[\left\langle\sum_{n=(k-1)T+1}^{kT}\boldsymbol{g}_{n},\boldsymbol{v}_{k}\right\rangle\right].$ (48)
Recall that $\boldsymbol{g}_{n}=\nabla F(\boldsymbol{z}_{n})+\boldsymbol{\epsilon}_{n}$, which implies for any $k\in\left[K\right]$,
$\mathbb{\mathbb{E}}\left[\left\langle\sum_{n=(k-1)T+1}^{kT}\boldsymbol{g}_{n},\boldsymbol{v}_{k}\right\rangle\right]$ $=\mathbb{\mathbb{E}}\left[\left\langle\sum_{n=(k-1)T+1}^{kT}\boldsymbol{\epsilon}_{n},\boldsymbol{v}_{k}\right\rangle\right]+\mathbb{\mathbb{E}}\left[\left\langle\sum_{n=(k-1)T+1}^{kT}\nabla F(\boldsymbol{z}_{n}),\boldsymbol{v}_{k}\right\rangle\right]$
$\leq\mathbb{\mathbb{E}}\left[\left\|\sum_{n=(k-1)T+1}^{kT}\boldsymbol{\epsilon}_{n}\right\|\left\|\boldsymbol{v}_{k}\right\|\right]+\mathbb{\mathbb{E}}\left[\left\langle\sum_{n=(k-1)T+1}^{kT}\nabla F(\boldsymbol{z}_{n}),\boldsymbol{v}_{k}\right\rangle\right]$
$=D\mathbb{\mathbb{E}}\left[\left\|\sum_{n=(k-1)T+1}^{kT}\boldsymbol{\epsilon}_{n}\right\|\right]-D\mathbb{\mathbb{E}}\left[\left\|\sum_{n=(k-1)T+1}^{kT}\nabla F(\boldsymbol{z}_{n})\right\|\right],$ (49)
where the second step is by Cauchy-Schwarz inequality and the last equation holds due to
$\boldsymbol{v}_{k}=-D\frac{\sum_{n=(k-1)T+1}^{kT}\nabla F(\boldsymbol{z}_{n})}{\left\|\sum_{n=(k-1)T+1}^{kT}\nabla F(\boldsymbol{z}_{n})\right\|},\forall k\in\left[K\right].$ (50)
Combine (48) and (49), apply $F(\boldsymbol{y}_{KT})\geq F_{\star}$, and rearrange terms to have
$\mathbb{\mathbb{E}}\left[\sum_{k=1}^{K}\frac{1}{K}\left\|\frac{1}{T}\sum_{n=(k-1)T+1}^{kT}\nabla F(\boldsymbol{z}_{n})\right\|\right]$
$\leq$ $\frac{F(\boldsymbol{y}_{0})-F_{\star}}{DKT}+\frac{\mathbb{\mathbb{E}}\left[R_{T}^{\mathsf{A}}\left(\boldsymbol{v}_{1},\dots,\boldsymbol{v}_{K}\right)\right]}{DKT}+\frac{\sum_{k=1}^{K}\mathbb{\mathbb{E}}\left[\left\|\sum_{n=(k-1)T+1}^{kT}\boldsymbol{\epsilon}_{n}\right\|\right]}{KT}.$ (51)
For any fixed $k\in\left[K\right]$, we apply Lemma 3 with $\boldsymbol{w}_{t}=\boldsymbol{\epsilon}_{(k-1)T+t},\forall t\in\left[T\right]$ to know
$\mathbb{\mathbb{E}}\left[\left\|\sum_{n=(k-1)T+1}^{kT}\boldsymbol{\epsilon}_{n}\right\|\right]$ $\leq 2\sqrt{2}\mathbb{\mathbb{E}}\left[\left(\sum_{n=(k-1)T+1}^{kT}\left\|\boldsymbol{\epsilon}_{n}\right\|^{\mathsf{p}}\right)^{\frac{1}{\mathsf{p}}}\right]$
$\leq 2\sqrt{2}\left(\sum_{n=(k-1)T+1}^{kT}\mathbb{\mathbb{E}}\left[\left\|\boldsymbol{\epsilon}_{n}\right\|^{\mathsf{p}}\right]\right)^{\frac{1}{\mathsf{p}}}\leq 2\sqrt{2}\sigma T^{\frac{1}{\mathsf{p}}},$ (52)
where the second step is by Hölder’s inequality (note that $\mathsf{p}>1$). Finally, we conclude the proof after plugging (52) back into (51). ∎
E.3 Proof of Theorem 5
Proof.
By Theorem 4, there is
$\mathbb{\mathbb{E}}\left[\sum_{k=1}^{K}\frac{1}{K}\left\|\frac{1}{T}\sum_{n=(k-1)T+1}^{kT}\nabla F(\boldsymbol{z}_{n})\right\|\right]\lesssim\frac{F(\boldsymbol{y}_{0})-F_{\star}}{DKT}+\frac{\mathbb{\mathbb{E}}\left[\mathsf{R}_{T}^{\mathsf{A}}(\boldsymbol{v}_{1},\dots,\boldsymbol{v}_{K})\right]}{DKT}+\frac{\sigma}{T^{1-\frac{1}{\mathsf{p}}}}.$ (53)
Note that $\mathsf{A}$ has the domain $\mathcal{X}=\mathcal{B}^{d}(D)$ and $s_{n}\sim\mathsf{Uniform}\left[0,1\right]$. Thus, for any $n\in\left[KT\right]$,
$\left\|\boldsymbol{x}_{n}\right\|\leq D$ and $s_{n}\in\left[0,1\right].$ (54)
We first lower bound the L.H.S. of (53). Given $k\in\left[K\right]$, for any $m<n\in\left\{(k-1)T+1,\dots,kT\right\}$, observe that
$\left\|\boldsymbol{z}_{n}-\boldsymbol{z}_{m}\right\|$ $=\left\|\boldsymbol{y}_{n-1}+s_{n}\boldsymbol{x}_{n}-\boldsymbol{y}_{m-1}-s_{m}\boldsymbol{x}_{m}\right\|=\left\|s_{n}\boldsymbol{x}_{n}-s_{m}\boldsymbol{x}_{m}+\sum_{i=m}^{n-1}\boldsymbol{x}_{i}\right\|$
[…]
Recall that $\bar{\boldsymbol{z}}_{k}=\frac{1}{T}\sum_{n=(k-1)T+1}^{kT}\boldsymbol{z}_{n}$ and $D=\delta/T$ now, then the above inequality implies
$\left\|\boldsymbol{z}_{n}-\bar{\boldsymbol{z}}_{k}\right\|\leq DT=\delta,\forall n\in\left\{(k-1)T+1,\dots,kT\right\},$ (55)
which means
$\boldsymbol{z}_{n}\in\mathcal{B}^{d}(\bar{\boldsymbol{z}}_{k},\delta),\forall n\in\left\{(k-1)T+1,\dots,kT\right\}.$
By the definition of $\left\|\nabla F(\bar{\boldsymbol{z}}_{k})\right\|_{\delta}$ (see Definition 1), there is
$\left\|\nabla F(\bar{\boldsymbol{z}}_{k})\right\|_{\delta}\leq\left\|\frac{1}{T}\sum_{n=(k-1)T+1}^{kT}\nabla F(\boldsymbol{z}_{n})\right\|.$ (56)
Next, we upper bound the R.H.S. of (53). By the definition of $K$-shifting regret (see (9)), there is
$\mathbb{\mathbb{E}}\left[\mathsf{R}_{T}^{\mathsf{A}}(\boldsymbol{v}_{1},\dots,\boldsymbol{v}_{K})\right]=\sum_{k=1}^{K}\mathbb{\mathbb{E}}\left[\sum_{n=(k-1)T+1}^{kT}\left\langle\boldsymbol{g}_{n},\boldsymbol{x}_{n}-\boldsymbol{v}_{k}\right\rangle\right].$
Note that we reset the stepsize in $\mathsf{A}$ after every $T$ iterations and $\boldsymbol{v}_{k}\in\mathcal{B}^{d}(D)$ by its definition (see (50)). Then for any $\mathsf{A}\in\left\{\mathsf{OGD},\mathsf{DA},\mathsf{AdaGrad}\right\}$, we can invoke its regret bound [^2] A minor point here is that the current function $\ell_{n}(\boldsymbol{x})=\left\langle\boldsymbol{g}_{n},\boldsymbol{x}\right\rangle$ does not entirely fit Assumption 1. We clarify that one does not need to worry about it, since all results proved in Section 3 hold under this change. For example, in the proof of Theorem 1, we can safely replace the L.H.S. of (14) with $\mathbb{\mathbb{E}}\left[\sum_{t=1}^{T}\left\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t}-\boldsymbol{x}\right\rangle\right]$. (i.e., Theorems 1, 2 and 3) to obtain
$\mathbb{\mathbb{E}}\left[\sum_{n=(k-1)T+1}^{kT}\left\langle\boldsymbol{g}_{n},\boldsymbol{x}_{n}-\boldsymbol{v}_{k}\right\rangle\right]\lesssim GD\sqrt{T}+\sigma DT^{1/\mathsf{p}},\forall k\in\left[K\right],$
which implies
$\mathbb{\mathbb{E}}\left[\mathsf{R}_{T}^{\mathsf{A}}(\boldsymbol{v}_{1},\dots,\boldsymbol{v}_{K})\right]\lesssim GDK\sqrt{T}+\sigma DKT^{1/\mathsf{p}}.$ (57)
Finally, we plug (56) and (57) back into (53), then use $D=\delta/T$ and $\Delta=F(\boldsymbol{y}_{0})-F_{\star}$ to have
$\mathbb{\mathbb{E}}\left[\frac{1}{K}\sum_{k=1}^{K}\left\|\nabla F(\bar{\boldsymbol{z}}_{k})\right\|_{\delta}\right]\lesssim\frac{\Delta}{\delta K}+\frac{G}{\sqrt{T}}+\frac{\sigma}{T^{1-\frac{1}{\mathsf{p}}}}.$
E.4 Proof of Corollary 3
Proof.
Recall that we pick
$K=\left\lfloor\frac{N}{T}\right\rfloor$ and $T=\left\lceil\frac{N}{2}\right\rceil\land\left(\left\lceil\left(\frac{\delta GN}{\Delta}\right)^{\frac{2}{3}}\right\rceil\lor\left\lceil\left(\frac{\delta\sigma N}{\Delta}\right)^{\frac{\mathsf{p}}{2\mathsf{p}-1}}\right\rceil\right),$
where $\Delta=F(\boldsymbol{y}_{0})-F_{\star}$. We invoke Theorem 5 and use $KT\geq N/4$ (see Fact 2) to obtain
$\mathbb{\mathbb{E}}\left[\frac{1}{K}\sum_{k=1}^{K}\left\|\nabla F(\bar{\boldsymbol{z}}_{k})\right\|_{\delta}\right]\lesssim\frac{\Delta T}{\delta N}+\frac{G}{\sqrt{T}}+\frac{\sigma}{T^{1-\frac{1}{\mathsf{p}}}}.$
By the definition of $T$, we know
$\frac{\Delta T}{\delta N}\lesssim\frac{\Delta}{\delta N}\left[1+\left(\frac{\delta GN}{\Delta}\right)^{\frac{2}{3}}+\left(\frac{\delta\sigma N}{\Delta}\right)^{\frac{\mathsf{p}}{2\mathsf{p}-1}}\right]=\frac{\Delta}{\delta N}+\frac{G^{\frac{2}{3}}\Delta^{\frac{1}{3}}}{(\delta N)^{\frac{1}{3}}}+\frac{\sigma^{\frac{\mathsf{p}}{2\mathsf{p}-1}}\Delta^{\frac{\mathsf{p}-1}{2\mathsf{p}-1}}}{(\delta N)^{\frac{\mathsf{p}-1}{2\mathsf{p}-1}}},$
and
$\frac{G}{\sqrt{T}}\lesssim\frac{G}{\sqrt{N}}+\frac{G^{\frac{2}{3}}\Delta^{\frac{1}{3}}}{(\delta N)^{\frac{1}{3}}},$ $\frac{\sigma}{T^{1-\frac{1}{\mathsf{p}}}}\lesssim\frac{\sigma}{N^{1-\frac{1}{\mathsf{p}}}}+\frac{\sigma^{\frac{\mathsf{p}}{2\mathsf{p}-1}}\Delta^{\frac{\mathsf{p}-1}{2\mathsf{p}-1}}}{(\delta N)^{\frac{\mathsf{p}-1}{2\mathsf{p}-1}}}.$
Therefore, there is
$\mathbb{\mathbb{E}}\left[\frac{1}{K}\sum_{k=1}^{K}\left\|\nabla F(\bar{\boldsymbol{z}}_{k})\right\|_{\delta}\right]\lesssim\frac{G}{\sqrt{N}}+\frac{\sigma}{N^{1-\frac{1}{\mathsf{p}}}}+\frac{\Delta}{\delta N}+\frac{G^{\frac{2}{3}}\Delta^{\frac{1}{3}}}{(\delta N)^{\frac{1}{3}}}+\frac{\sigma^{\frac{\mathsf{p}}{2\mathsf{p}-1}}\Delta^{\frac{\mathsf{p}-1}{2\mathsf{p}-1}}}{(\delta N)^{\frac{\mathsf{p}-1}{2\mathsf{p}-1}}}.$
E.5 Extension to the Case of Unknown Problem-Dependent Parameters
In Corollary 7, we show how to set $K$ and $T$ when all problem-dependent parameters are unknown. It is particularly meaningful for $\mathsf{AdaGrad}$. As in that case, the rate is achieved without knowing any problem-dependent parameter. This kind of result is the first to appear for nonsmooth nonconvex optimization with heavy tails. However, the rate is not as good as Corollary 3. It is currently unclear whether the same bound $1/(\delta N)^{\frac{\mathsf{p}-1}{2\mathsf{p}-1}}$ as in Corollary 3 can be obtained when no information about the problem is known.
Corollary 7.
Under the same setting of Theorem 5, suppose we have $N\geq 2$ stochastic gradient budgets, taking $K=\left\lfloor N/T\right\rfloor$ and $T=\left\lceil N/2\right\rceil\land\left\lceil(\delta N)^{\frac{2}{3}}\right\rceil$, we have
$\mathbb{\mathbb{E}}\left[\frac{1}{K}\sum_{k=1}^{K}\left\|\nabla F(\bar{\boldsymbol{z}}_{k})\right\|_{\delta}\right]\lesssim\frac{\Delta}{(\delta N)\land(\delta N)^{\frac{1}{3}}}+\frac{G}{\sqrt{N}\land(\delta N)^{\frac{1}{3}}}+\frac{\sigma}{N^{1-\frac{1}{\mathsf{p}}}\land(\delta N)^{\frac{2(\mathsf{p}-1)}{3\mathsf{p}}}}.$
Proof.
We invoke Theorem 5 and use $KT\geq N/4$ (see Fact 2) to obtain
$\mathbb{\mathbb{E}}\left[\frac{1}{K}\sum_{k=1}^{K}\left\|\nabla F(\bar{\boldsymbol{z}}_{k})\right\|_{\delta}\right]\lesssim\frac{\Delta T}{\delta N}+\frac{G}{\sqrt{T}}+\frac{\sigma}{T^{1-\frac{1}{\mathsf{p}}}}.$
By the definition of $T$, we know
$\frac{\Delta T}{\delta N}\lesssim\frac{\Delta}{\delta N}\left[1+(\delta N)^{\frac{2}{3}}\right]\lesssim\frac{\Delta}{(\delta N)\land(\delta N)^{\frac{1}{3}}}.$
and
$\frac{G}{\sqrt{T}}$ $\lesssim\frac{G}{\sqrt{N}}+\frac{G}{(\delta N)^{\frac{1}{3}}}\lesssim\frac{G}{\sqrt{N}\land(\delta N)^{\frac{1}{3}}},$
$\frac{\sigma}{T^{1-\frac{1}{\mathsf{p}}}}$ $\lesssim\frac{\sigma}{N^{1-\frac{1}{\mathsf{p}}}}+\frac{\sigma}{(\delta N)^{\frac{2(\mathsf{p}-1)}{3\mathsf{p}}}}\lesssim\frac{\sigma}{N^{1-\frac{1}{\mathsf{p}}}\land(\delta N)^{\frac{2(\mathsf{p}-1)}{3\mathsf{p}}}}.$
Therefore, there is
$\mathbb{\mathbb{E}}\left[\frac{1}{K}\sum_{k=1}^{K}\left\|\nabla F(\bar{\boldsymbol{z}}_{k})\right\|_{\delta}\right]\lesssim\frac{\Delta}{(\delta N)\land(\delta N)^{\frac{1}{3}}}+\frac{G}{\sqrt{N}\land(\delta N)^{\frac{1}{3}}}+\frac{\sigma}{N^{1-\frac{1}{\mathsf{p}}}\land(\delta N)^{\frac{2(\mathsf{p}-1)}{3\mathsf{p}}}}.$
E.6 Proof of Theorem 6
In this section, our ultimate goal is to prove Theorem 6. The analysis follows the framework first established in [3] and later developed by [7] but with some necessary (though minor) variation to make it compatible with heavy-tailed noise. In the following, we will slightly abuse the notation $\mathsf{A}$ to denote any possible randomized first-order algorithm instead of an online learning algorithm.
E.6.1 Basic Definitions
To begin with, we introduce some basic definitions given in [3].
Definition 3 (stochastic first-order oracle).
Given a differentiable function $F:\mathbb{R}^{d}\to\mathbb{R}$, a tuple $(\boldsymbol{g},\mathcal{R},\mathbb{P}_{r})$ is called a stochastic first-order oracle of $F$ if $\mathbb{P}_{r}$ is a probability distribution on the measurable space $\mathcal{R}$ and $\boldsymbol{g}:\mathbb{R}^{d}\times\mathcal{R}\to\mathbb{R}^{d}$ satisfies $\mathbb{\mathbb{E}}_{r\sim\mathbb{P}_{r}}\left[\boldsymbol{g}(\boldsymbol{x},r)\right]=\nabla F(\boldsymbol{x}),\forall\boldsymbol{x}\in\mathbb{R}^{d}$.
Remark 10.
When the context is clear, we will omit $F$, $\mathcal{R}$ and $\mathbb{P}_{r}$, and simply call $\boldsymbol{g}$ as a stochastic first-order oracle.
Definition 4 (randomized algorithm).
A randomized algorithm $\mathsf{A}$ consists of a probability distribution $\mathbb{P}_{s}$ over a measurable space $\mathcal{S}$ and a sequence of measurable mappings $\mathsf{A}_{t},\forall t\in\mathbb{\mathbb{N}}$ such that every $\mathsf{A}_{t}$ takes a common random seed $s\in\mathcal{S}$ and the first t1t-1 oracle responses to produce the $t$-th query. Concretely, given a differentiable function $F$ equipped with a stochastic first-order oracle $(\boldsymbol{g},\mathcal{R},\mathbb{P}_{r})$, the sequence $\boldsymbol{x}_{t},\forall t\in\mathbb{\mathbb{N}}$ produced by $\mathsf{A}$ to optimize $F$ is recursively defined as
$\boldsymbol{x}_{t}=\mathsf{A}_{t}(s,\boldsymbol{g}(\boldsymbol{x}_{t-1},r_{t-1}),\dots,\boldsymbol{g}(\boldsymbol{x}_{1},r_{1})),\forall t\in\mathbb{\mathbb{N}},$
where $s\sim\mathbb{P}_{s}$ is drawn a single time at the beginning of the algorithm and $r_{t}\sim\mathbb{P}_{r},\forall t\in\mathbb{\mathbb{N}}$ is a sequence of i.i.d. random variables. Moreover, $\mathsf{A}_{\mathsf{rand}}$ denotes the set containing all randomized algorithms.
Next, we require a useful concept named probability-$q$ zero-chain. Before formally stating what it is, we need some notations. Given $\boldsymbol{x}\in\mathbb{R}^{d}$ and $\alpha\in\left[0,1\right]$, $\mathsf{prog}_{\alpha}(\boldsymbol{x})$ denotes the largest index whose entry is $\alpha$-far from 0, i.e.,
$\mathsf{prog}_{\alpha}(\boldsymbol{x})\triangleq\max\left\{i\in\left[d\right]:\left|\boldsymbol{x}\left[i\right]\right|>\alpha\right\}$ where $\max\varnothing\triangleq 0.$
In addition, for any $j\in\left\{0\right\}\cup\mathbb{\mathbb{N}}$, let $\boldsymbol{x}_{\leq j}\left[i\right]\triangleq\boldsymbol{x}\left[i\right]\mathds{1}\left[i\leq j\right],\forall i\in\left[d\right]$ be the truncated version of $\boldsymbol{x}$. Now we are ready to provide the definition of the probability-$q$ zero-chain.
Definition 5 (probability-$q$ zero-chain).
A stochastic first-order oracle $(\boldsymbol{g},\mathcal{R},\mathbb{P}_{r})$ is called a probability-$q$ zero-chain if and only if
$\mathbb{P}_{r}\left[\forall\boldsymbol{x}\in\mathbb{R}^{d}:\mathsf{prog}_{0}(\boldsymbol{g}(\boldsymbol{x},r))\leq\mathsf{prog}_{1/4}(\boldsymbol{x})\text{ and }\boldsymbol{g}(\boldsymbol{x},r)=\boldsymbol{g}(\boldsymbol{x}_{\leq\mathsf{prog}_{1/4}(\boldsymbol{x})},r)\right]\geq 1-q,$
and
[…]
E.6.2 Useful Existing Results
In this part, we list some useful existing results from [3] and [7].
Given $d\geq T\in\mathbb{\mathbb{N}}$ and a differentiable function $F_{T}:\mathbb{R}^{T}\to\mathbb{R}$ with a stochastic first-order oracle $\boldsymbol{g}_{T}$ that is a probability-$q$ zero-chain, their rotated variants parametrized by a matrix $U\in\mathsf{Ortho}(d,T)\triangleq\left\{U\in\mathbb{R}^{d\times T}:U^{\top}U=I_{T}\right\}$ (where $I_{T}$ is the $T$-dimensional identity matrix) are defined as
$\bar{F}_{T,U}(\boldsymbol{x})\triangleq F_{T}(U^{\top}\boldsymbol{x})$ and $\bar{\boldsymbol{g}}_{T,U}(\boldsymbol{x},r)\triangleq U\boldsymbol{g}_{T}(U^{\top}\boldsymbol{x},r).$ (58)
Clearly, $\bar{\boldsymbol{g}}_{T,U}$ is a stochastic first-order oracle of $\bar{F}_{T,U}$. In addition, we emphasize that the input variable $\boldsymbol{x}$ is from $\mathbb{R}^{d}$ instead of $\mathbb{R}^{T}$ now.
With $\bar{F}_{T,U}$ and $\bar{\boldsymbol{g}}_{T,U}$, we state the following lemma due to [3].
Lemma 4 (Lemma 5 of [3]).
Given $q,\iota\in\left(0,1\right)$, $R>0$, $T\in\mathbb{\mathbb{N}}$, $d\in\mathbb{\mathbb{N}}$ satisfying $d\geq T+32R^{2}\log\frac{2T^{2}}{q\iota}$ and a randomized algorithm $\mathsf{A}\in\mathsf{A}_{\mathsf{rand}}$, suppose the output of $\mathsf{A}$ always has a norm bounded by $R$, let $\boldsymbol{x}_{t},\forall t\in\mathbb{\mathbb{N}}$ be the trajectory produced by applying $\mathsf{A}$ to optimize $\bar{F}_{T,U}$ interacting with the stochastic first-order oracle $\bar{\boldsymbol{g}}_{T,U}$, where $U$ is drawn from $\mathsf{Ortho}(d,T)$ uniformly, then there is
$\Pr\left[\mathsf{prog}_{1/4}\left(U^{\top}\boldsymbol{x}_{t}\right)<T,\forall t\leq\frac{T-\log(2/\iota)}{2q}\right]\geq 1-\iota.$
Here $\Pr$ takes into account all randomness over $\mathsf{A}$, $\boldsymbol{g}_{T}$, and $U$.
Note that Lemma 4 can be only applied to a randomized algorithm with bounded outputs. To overcome this issue, we need the following variants of $\bar{F}_{T,U}$ and $\bar{\boldsymbol{g}}_{T,U}$ introduced by [7]:
$\widehat{F}_{T,U}(\boldsymbol{x})$ $\triangleq\bar{F}_{T,U}(\boldsymbol{\rho}_{R,d}(\boldsymbol{x}))+\eta\boldsymbol{x}^{\top}\boldsymbol{\rho}_{R,d}(\boldsymbol{x}),$ (59)
$\widehat{\boldsymbol{g}}_{T,U}(\boldsymbol{x},r)$ $\triangleq\mathcal{J}_{R,d}(\boldsymbol{x})^{\top}\bar{\boldsymbol{g}}_{T,U}(\boldsymbol{\rho}_{R,d}(\boldsymbol{x}),r)+\eta\nabla(\boldsymbol{x}^{\top}\boldsymbol{\rho}_{R,d}(\boldsymbol{x})),$ (60)
where $\boldsymbol{\rho}_{R,d}(\boldsymbol{x})\triangleq\frac{\boldsymbol{x}}{\sqrt{1+\left\|\boldsymbol{x}\right\|^{2}/R^{2}}}$ is a bijection from $\mathbb{R}^{d}$ to $\mathsf{int}\mathcal{B}^{d}(R)$ [^3] As one can check, $\boldsymbol{\rho}_{R,d}^{-1}(\boldsymbol{x}):\mathsf{int}\mathcal{B}^{d}(R)\to\mathbb{R}^{d}$ exists and equals $\frac{\boldsymbol{x}}{\sqrt{1-\left\|\boldsymbol{x}\right\|^{2}/R^{2}}}$., $\mathcal{J}_{R,d}$ denotes the Jacobian of $\boldsymbol{\rho}_{R,d}$, and $R,\eta>0$ are two constants being determined later.
Before moving on, we state some useful properties of $\boldsymbol{\rho}_{R,d}$ here.
Lemma 5 (Lemma 13 of [3] and Proposition 29 of [7]).
Let $\left\|\cdot\right\|_{\mathsf{op}}$ denotes the operator norm, then for $\boldsymbol{\rho}_{R,d}:\mathbb{R}^{d}\to\mathbb{R}^{d},\boldsymbol{x}\mapsto\frac{\boldsymbol{x}}{\sqrt{1+\left\|\boldsymbol{x}\right\|^{2}/R^{2}}}$, there are
  1. 1.
    $\nabla(\boldsymbol{x}^{\top}\boldsymbol{\rho}_{R,d}(\boldsymbol{x}))=\left(2-\left\|\boldsymbol{\rho}_{R,d}(\boldsymbol{x})\right\|^{2}/R^{2}\right)\boldsymbol{\rho}_{R,d}(\boldsymbol{x})$,
  2. 2.
    $\left\|\boldsymbol{\rho}_{R,d}(\boldsymbol{x})-\boldsymbol{\rho}_{R,d}(\boldsymbol{y})\right\|\leq\left\|\boldsymbol{x}-\boldsymbol{y}\right\|$,
  3. 3.
    $\mathcal{J}_{R,d}(\boldsymbol{x})=\frac{I_{d}-\boldsymbol{\rho}_{R,d}(\boldsymbol{x})\boldsymbol{\rho}_{R,d}(\boldsymbol{x})^{\top}/R^{2}}{\sqrt{1+\left\|\boldsymbol{x}\right\|^{2}/R^{2}}}$,
  4. 4.
    $\left\|\mathcal{J}_{R,d}(\boldsymbol{x})\right\|_{\mathsf{op}}=\frac{1}{\sqrt{1+\left\|\boldsymbol{x}\right\|^{2}/R^{2}}}\leq 1$,
  5. 5.
    $\left\|\mathcal{J}_{R,d}(\boldsymbol{x})-\mathcal{J}_{R,d}(\boldsymbol{y})\right\|_{\mathsf{op}}\leq\frac{3}{R}\left\|\boldsymbol{x}-\boldsymbol{y}\right\|$.
We then can combine (58), (59), (60) and Lemma 5 to have
$\widehat{F}_{T,U}(\boldsymbol{x})$ $=\bar{F}_{T,U}(\boldsymbol{\rho}_{R,d}(\boldsymbol{x}))+\eta\boldsymbol{x}^{\top}\boldsymbol{\rho}_{R,d}(\boldsymbol{x})=F_{T}(U^{\top}\boldsymbol{\rho}_{R,d}(\boldsymbol{x}))+\eta\boldsymbol{x}^{\top}\boldsymbol{\rho}_{R,d}(\boldsymbol{x}),$ (61)
$\nabla\widehat{F}_{T,U}(\boldsymbol{x})$ $=\mathcal{J}_{R,d}(\boldsymbol{x})^{\top}\nabla\bar{F}_{T,U}(\boldsymbol{\rho}_{R,d}(\boldsymbol{x}))+\eta\left(2-\left\|\boldsymbol{\rho}_{R,d}(\boldsymbol{x})\right\|^{2}/R^{2}\right)\boldsymbol{\rho}_{R,d}(\boldsymbol{x})$ (62)
$=\mathcal{J}_{R,d}(\boldsymbol{x})^{\top}U\nabla F_{T}(U^{\top}\boldsymbol{\rho}_{R,d}(\boldsymbol{x}))+\eta\left(2-\left\|\boldsymbol{\rho}_{R,d}(\boldsymbol{x})\right\|^{2}/R^{2}\right)\boldsymbol{\rho}_{R,d}(\boldsymbol{x}),$ (63)
$\widehat{\boldsymbol{g}}_{T,U}(\boldsymbol{x},r)$ $=\mathcal{J}_{R,d}(\boldsymbol{x})^{\top}\bar{\boldsymbol{g}}_{T,U}(\boldsymbol{\rho}_{R,d}(\boldsymbol{x}),r)+\eta\left(2-\left\|\boldsymbol{\rho}_{R,d}(\boldsymbol{x})\right\|^{2}/R^{2}\right)\boldsymbol{\rho}_{R,d}(\boldsymbol{x})$ (64)
$=\mathcal{J}_{R,d}(\boldsymbol{x})^{\top}U\boldsymbol{g}_{T}(U^{\top}\boldsymbol{\rho}_{R,d}(\boldsymbol{x}),r)+\eta\left(2-\left\|\boldsymbol{\rho}_{R,d}(\boldsymbol{x})\right\|^{2}/R^{2}\right)\boldsymbol{\rho}_{R,d}(\boldsymbol{x}).$ (65)
Now we are ready to give the following Lemma 6, which is almost identical to Lemma 25 in [7] by differing up to numerical constants. Since Lemma 6 is particularly important, we therefore provide its full proof here.
Lemma 6 (Lemma 25 of [7]).
For $T\in\mathbb{\mathbb{N}}$, suppose that $F_{T}$ satisfies the following two requirements:
  1. 1.
    For all $\boldsymbol{x}\in\mathbb{R}^{T}$, $\left\|\nabla F_{T}(\boldsymbol{x})\right\|\leq\gamma\sqrt{T}$ where $\gamma>0$ is a constant satisfying $\frac{1}{\sqrt{5}}-\frac{39}{140}\geq\frac{10}{63}+\frac{5\sqrt{5}}{126\gamma}\Leftrightarrow\gamma\geq\frac{50\sqrt{5}}{252\sqrt{5}-551}\approx 8.95$.
  2. 2.
    For all $\boldsymbol{x}\in\mathbb{R}^{T}$, if $\mathsf{prog}_{1}(\boldsymbol{x})<T$, then $\left\|\nabla F_{T}(\boldsymbol{x})\right\|\geq\left|\nabla F_{T}(\boldsymbol{x})\left[\mathsf{prog}_{1}(\boldsymbol{x})+1\right]\right|>1$.
Given $q,\iota\in\left(0,1\right)$, $\eta=\frac{1}{\sqrt{5}}-\frac{39}{140}$, $R=7\gamma\sqrt{T}$, $d\in\mathbb{\mathbb{N}}$ satisfying $d\geq T+32R^{2}\log\frac{2T^{2}}{q\iota}$ and a randomized algorithm $\mathsf{A}\in\mathsf{A}_{\mathsf{rand}}$, let $\boldsymbol{x}_{t},\forall t\in\mathbb{\mathbb{N}}$ be the trajectory produced by applying $\mathsf{A}$ to optimize $\widehat{F}_{T,U}$ interacting with the stochastic first-order oracle $\widehat{\boldsymbol{g}}_{T,U}$, where $U$ is drawn from $\mathsf{Ortho}(d,T)$ uniformly, then there is
$\Pr\left[\left\|\nabla\widehat{F}_{T,U}(\boldsymbol{x}_{t})\right\|\geq\frac{1}{2},\forall t\leq\frac{T-\log(2/\iota)}{2q}\right]\geq 1-\iota.$
Here $\Pr$ takes into account all randomness over $\mathsf{A}$, $\boldsymbol{g}_{T}$ and $U$.
Proof.
By definition of randomized algorithms (see Definition 4), we have
$\boldsymbol{x}_{t}=\mathsf{A}_{t}\left(s,\widehat{\boldsymbol{g}}_{T,U}(\boldsymbol{x}_{t-1},r_{t-1}),\dots,\widehat{\boldsymbol{g}}_{T,U}(\boldsymbol{x}_{1},r_{1})\right),\forall t\in\mathbb{\mathbb{N}}.$ (66)
Now let us consider another sequence
$\boldsymbol{y}_{t}\triangleq\boldsymbol{\rho}_{R,d}(\boldsymbol{x}_{t})=\frac{\boldsymbol{x}_{t}}{\sqrt{1+\left\|\boldsymbol{x}_{t}\right\|/R^{2}}}\in\mathsf{int}\mathcal{B}^{d}(R),\forall t\in\mathbb{\mathbb{N}}.$ (67)
Note that
[…]
$=\mathcal{G}(\boldsymbol{y}_{t},\bar{\boldsymbol{g}}_{T,U}(\boldsymbol{y}_{t},r_{t})),$ (68)
where $\mathcal{G}(\boldsymbol{u},\boldsymbol{v}):\mathsf{int}\mathcal{B}^{d}(R)\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ is a measurable mapping defined as
$\mathcal{G}(\boldsymbol{u},\boldsymbol{v})\triangleq\mathcal{J}_{R,d}(\boldsymbol{\rho}_{R,d}^{-1}(\boldsymbol{u}))^{\top}\boldsymbol{v}+\eta\left(2-\left\|\boldsymbol{u}\right\|^{2}/R^{2}\right)\boldsymbol{u}.$
Thus, we can write
[…]
By a simple induction, there exists a sequence of measurable mappings $\mathsf{A}_{t}^{\boldsymbol{y}},\forall t\in\mathbb{\mathbb{N}}$ such that
$\boldsymbol{y}_{t}=\mathsf{A}_{t}^{\boldsymbol{y}}(s,\bar{\boldsymbol{g}}_{T,U}(\boldsymbol{y}_{t-1},r_{t-1}),\dots,\bar{\boldsymbol{g}}_{T,U}(\boldsymbol{y}_{1},r_{1})),\forall t\in\mathbb{\mathbb{N}}.$
[…]
$\Pr\left[\mathsf{prog}_{1/4}\left(U^{\top}\boldsymbol{y}_{t}\right)<T,\forall t\leq\frac{T-\log(2/\iota)}{2q}\right]\geq 1-\iota.$ (69)
Now we recall
[…]
Moreover, by the first requirement on $F_{T}$, there is
[…]
We fix $t\leq\frac{T-\log(2/\iota)}{2q}$ and consider the following two cases:
  1. Case 1.
    $\left\|\boldsymbol{x}_{t}\right\|\geq\frac{R}{2}$. In this case, we first have
[…]
    $\eta\left(2-\left\|\boldsymbol{y}_{t}\right\|^{2}/R^{2}\right)\left\|\boldsymbol{y}_{t}\right\|\geq\eta R\min_{\frac{1}{\sqrt{5}}\leq a\leq 1}(2-a^{2})a=\frac{9\eta R}{5\sqrt{5}}.$
    In addition, there is
[…]
    As such, by our choices of $\eta$ and $R$,
    $\left\|\nabla\widehat{F}_{T,U}(\boldsymbol{x}_{t})\right\|\geq\frac{9\eta R}{5\sqrt{5}}-\frac{2\gamma\sqrt{T}}{\sqrt{5}}\geq\frac{1}{2}.$
  2. Case 2.
    $\left\|\boldsymbol{x}_{t}\right\|<\frac{R}{2}$ . In this case, we introduce
[…]
    Let $\boldsymbol{u}_{j_{t}}\in\mathbb{R}^{d}$ denotes the $j_{t}$-th column of $U$. By the definition of $j_{t}$, there is
    $\left|\left\langle\boldsymbol{u}_{j_{t}},\boldsymbol{y}_{t}\right\rangle\right|=\left|\left(U^{\top}\boldsymbol{y}_{t}\right)\left[j_{t}\right]\right|<1.$ (72)
    In addition, we have
[…]
    where the last step is by the second requirement on $F_{T}$. Now we compute
    $\left\|\nabla\widehat{F}_{T,U}(\boldsymbol{x}_{t})\right\|\overset{\left\|\boldsymbol{u}_{j_{t}}\right\|=1}{\geq}\left|\left\langle\boldsymbol{u}_{j_{t}},\nabla\widehat{F}_{T,U}(\boldsymbol{x}_{t})\right\rangle\right|$
[…]
    $\geq$ $\frac{\left|\left\langle\boldsymbol{u}_{j_{t}},\nabla\bar{F}_{T,U}(\boldsymbol{y}_{t})\right\rangle\right|}{\sqrt{1+\left\|\boldsymbol{x}_{t}\right\|^{2}/R^{2}}}-\left|\left\langle\boldsymbol{u}_{j_{t}},\boldsymbol{y}_{t}\right\rangle\right|\left[\frac{\left|\left\langle\boldsymbol{y}_{t},\nabla\bar{F}_{T,U}(\boldsymbol{y}_{t})\right\rangle\right|}{R^{2}\sqrt{1+\left\|\boldsymbol{x}_{t}\right\|^{2}/R^{2}}}+\eta\left(2-\left\|\boldsymbol{y}_{t}\right\|^{2}/R^{2}\right)\right]$
    $\overset{(a)}{\geq}$ $\frac{1}{\sqrt{1+\left\|\boldsymbol{x}_{t}\right\|^{2}/R^{2}}}-\frac{\left\|\boldsymbol{y}_{t}\right\|\left\|\nabla\bar{F}_{T,U}(\boldsymbol{y}_{t})\right\|}{R^{2}\sqrt{1+\left\|\boldsymbol{x}_{t}\right\|^{2}/R^{2}}}-2\eta\overset{(b)}{\geq}\frac{2}{\sqrt{5}}-\frac{2\gamma\sqrt{T}}{5R}-2\eta\overset{(c)}{=}\frac{1}{2},$
[…]
We combine two cases to conclude
$\Pr\left[\left\|\nabla\widehat{F}_{T,U}(\boldsymbol{x}_{t})\right\|\geq\frac{1}{2},\forall t\leq\frac{T-\log(2/\iota)}{2q}\right]\geq 1-\iota.$
Lastly, we recall the following hard instance and its stochastic first-order oracle studied in [3].
Lemma 7 (Lemma 2 of [3]).
For any $T\in\mathbb{\mathbb{N}}$, there exists a differentiable function $F_{T}:\mathbb{R}^{T}\to\mathbb{R}$ satisfying the following properties:
  1. 1.
    $F_{T}(\boldsymbol{0})=0$ and $\inf_{\boldsymbol{x}\in\mathbb{R}^{T}}F_{T}(\boldsymbol{x})\geq-fT$, where $f=12$.
  2. 2.
    $F_{T}$ is $\ell$-smooth, where $\ell=152$.
  3. 3.
    For all $\boldsymbol{x}\in\mathbb{R}^{T}$, $\left\|\nabla F_{T}(\boldsymbol{x})\right\|_{\infty}\leq\gamma$, where $\gamma=23$.
  4. 4.
    For all $\boldsymbol{x}\in\mathbb{R}^{T}$, $\mathsf{prog}_{0}(\nabla F_{T}(\boldsymbol{x}))\leq\mathsf{prog}_{1/2}(\boldsymbol{x})+1$.
  5. 5.
    For all $\boldsymbol{x}\in\mathbb{R}^{T}$ and $i\triangleq\mathsf{prog}_{1/2}(\boldsymbol{x})$, $\nabla F_{T}(\boldsymbol{x})=\nabla F_{T}(\boldsymbol{x}_{\leq i+1})$ and $\left[\nabla F_{T}(\boldsymbol{x})\right]_{\leq i}=\left[\nabla F_{T}(\boldsymbol{x}_{\leq i})\right]_{\leq i}$.
  6. 6.
    For all $\boldsymbol{x}\in\mathbb{R}^{T}$, if $\mathsf{prog}_{1}(\boldsymbol{x})<T$, then $\left\|\nabla F_{T}(\boldsymbol{x})\right\|\geq\left|\nabla F_{T}(\boldsymbol{x})\left[\mathsf{prog}_{1}(\boldsymbol{x})+1\right]\right|>1$.
Lemma 8 (Lemma 3 of [3]).
For any $T\in\mathbb{\mathbb{N}}$ and $F_{T}$ in Lemma 7, the following $\boldsymbol{g}_{T}:\mathbb{R}^{T}\times\mathcal{R}\to\mathbb{R}^{T}$, where $\mathcal{R}\triangleq\left\{0,1\right\}$, is a stochastic first-order oracle of $F_{T}$:
$\boldsymbol{g}_{T}(\boldsymbol{x},r)\left[i\right]\triangleq\begin{cases}\nabla F_{T}(\boldsymbol{x})\left[i\right]&i\neq\mathsf{prog}_{1/4}(\boldsymbol{x})+1\\ \frac{r}{q}\nabla F_{T}(\boldsymbol{x})\left[i\right]&i=\mathsf{prog}_{1/4}(\boldsymbol{x})+1\end{cases},\forall i\in\left[T\right],$
where $r=\mathsf{Bernoulli}(q)$ for some $q\in\left(0,1\right)$. Moreover, $\boldsymbol{g}_{T}$ is a probability-$q$ zero-chain.
E.6.3 Analysis under Heavy-Tailed Noise
From now on, we need to diverge from [3] and [7], since both of which are under the finite variance case (i.e., $\mathsf{p}=2$) instead of heavy-tailed noise (i.e., $\mathsf{p}\in\left(1,2\right]$).
Lemma 9 (variation of Lemma 7 in [3] and Lemma 26 in [7]).
The instances $\widehat{F}_{T,U}$ and $\widehat{\boldsymbol{g}}_{T,U}$ (see (59) and (60)) constructed based on $F_{T}$ in Lemma 7 and $\boldsymbol{g}_{T}$ in Lemma 8 under $\eta=\frac{1}{\sqrt{5}}-\frac{39}{140}$ and $R=7\gamma\sqrt{T}$ for $\gamma=23$ satisfy the following properties:
  1. 1.
    $\widehat{F}_{T,U}(\boldsymbol{0})-\inf_{\boldsymbol{x}\in\mathbb{R}^{d}}\widehat{F}_{T,U}(\boldsymbol{x})\leq\widehat{f}T$, where $\widehat{f}=12$.
  2. 2.
    $\widehat{F}_{T,U}$ is $^\widehat{\ell}$-smooth, where $\widehat{\ell}=154$.
  3. 3.
    For all $\boldsymbol{x}\in\mathbb{R}^{d}$, $\left\|\nabla\widehat{F}_{T,U}(\boldsymbol{x})\right\|\leq\widehat{\gamma}\sqrt{T}$, where $\widehat{\gamma}=53$.
  4. 4.
    For all $\boldsymbol{x}\in\mathbb{R}^{d}$ and $\mathsf{p}\in\left(1,2\right]$, $\mathbb{\mathbb{E}}_{r}\left[\left\|\nabla\widehat{F}_{T,U}(\boldsymbol{x})-\widehat{\boldsymbol{g}}_{T,U}(\boldsymbol{x},r)\right\|^{\mathsf{p}}\right]\leq\frac{\left(2\gamma\right)^{\mathsf{p}}\left(1-q^{\mathsf{p}-1}\right)}{(\mathsf{p}-1)q^{\mathsf{p}-1}}$.
Proof.
First, we know
[…]
Next, for any $\boldsymbol{x}\in\mathbb{R}^{d}$,
[…]
By Lemma 14 in [3] and Lemma 7, $\mathcal{J}_{R,d}(\boldsymbol{x})^{\top}U\nabla F_{T}(U^{\top}\boldsymbol{\rho}_{R,d}(\boldsymbol{x}))$ is $\ell+\frac{3\gamma\sqrt{T}}{R}\overset{R=7\gamma\sqrt{T}}{=}\ell+\frac{3}{7}$-Lipschitz for $\ell=152$. Moreover, we have
$\left\|\left(2-\left\|\boldsymbol{\rho}_{R,d}(\boldsymbol{x})\right\|^{2}/R^{2}\right)\boldsymbol{\rho}_{R,d}(\boldsymbol{x})-\left(2-\left\|\boldsymbol{\rho}_{R,d}(\boldsymbol{y})\right\|^{2}/R^{2}\right)\boldsymbol{\rho}_{R,d}(\boldsymbol{y})\right\|$
$\leq$ $2\left\|\boldsymbol{\rho}_{R,d}(\boldsymbol{x})-\boldsymbol{\rho}_{R,d}(\boldsymbol{y})\right\|+\frac{\left\|\boldsymbol{\rho}_{R,d}(\boldsymbol{y})\right\|^{2}}{R^{2}}\left\|\boldsymbol{\rho}_{R,d}(\boldsymbol{y})-\boldsymbol{\rho}_{R,d}(\boldsymbol{x})\right\|$
$+\frac{\left|\left\|\boldsymbol{\rho}_{R,d}(\boldsymbol{y})\right\|^{2}-\left\|\boldsymbol{\rho}_{R,d}(\boldsymbol{x})\right\|^{2}\right|}{R^{2}}\left\|\boldsymbol{\rho}_{R,d}(\boldsymbol{x})\right\|$
$\overset{(a)}{\leq}$ $2\left\|\boldsymbol{x}-\boldsymbol{y}\right\|+\left\|\boldsymbol{y}-\boldsymbol{x}\right\|+2\left\|\boldsymbol{y}-\boldsymbol{x}\right\|=5\left\|\boldsymbol{x}-\boldsymbol{y}\right\|,$
[…]
Moreover, we observe that
[…]
Note that for $\gamma=23$,
[…]
and
$\eta\left(2-\left\|\boldsymbol{\rho}_{R,d}(\boldsymbol{x})\right\|^{2}/R^{2}\right)\left\|\boldsymbol{\rho}_{R,d}(\boldsymbol{x})\right\|$ $\overset{\left\|\boldsymbol{\rho}_{R,d}(\cdot)\right\|\leq R}{\leq}\eta R\max_{0\leq a\leq 1}(2-a^{2})a=\eta R\cdot\frac{4}{3}\sqrt{\frac{2}{3}}$
$=\left(\frac{1}{\sqrt{5}}-\frac{39}{140}\right)\cdot 7\cdot\frac{4}{3}\sqrt{\frac{2}{3}}\cdot\gamma\sqrt{T}\leq 1.3\gamma\sqrt{T}.$
We hence have $\left\|\nabla\widehat{F}_{T,U}(\boldsymbol{x})\right\|\leq 2.3\gamma\sqrt{T}\leq 53\sqrt{T}=\widehat{\gamma}\sqrt{T}$.
Lastly, still for $\gamma=23$, we compute
[…]
$\leq\left\|\mathcal{J}_{R,d}(\boldsymbol{x})\right\|_{\mathsf{op}}\left\|U\left(\nabla F_{T}(U^{\top}\boldsymbol{\rho}_{R,d}(\boldsymbol{x}))-\boldsymbol{g}_{T}(U^{\top}\boldsymbol{\rho}_{R,d}(\boldsymbol{x}),r)\right)\right\|$
[…]
$\overset{U^{\top}U=I_{T}}{=}\left\|\nabla F_{T}(U^{\top}\boldsymbol{\rho}_{R,d}(\boldsymbol{x}))-\boldsymbol{g}_{T}(U^{\top}\boldsymbol{\rho}_{R,d}(\boldsymbol{x}),r)\right\|$
[…]
which implies
$\mathbb{\mathbb{E}}_{r}\left[\left\|\nabla\widehat{F}_{T,U}(\boldsymbol{x})-\widehat{\boldsymbol{g}}_{T,U}(\boldsymbol{x},r)\right\|^{\mathsf{p}}\right]$ $\leq\gamma^{\mathsf{p}}\left(1-q+\frac{(1-q)^{\mathsf{p}}}{q^{\mathsf{p}-1}}\right)=\gamma^{\mathsf{p}}(1-q)\left(\frac{q^{\mathsf{p}-1}+(1-q)^{\mathsf{p}-1}}{q^{\mathsf{p}-1}}\right)$
[…]
Next, inspired by [7], we first prove a lower bound for heavy-tailed smooth nonconvex optimization, as presented in the following Theorem 11.
Theorem 11.
For any $\Delta>0$, $H>0$, $\mathsf{p}\in\left(1,2\right]$, $\sigma\geq 0$, $0<\epsilon\leq\sqrt{\frac{\Delta H}{96\cdot 12\cdot 154}}$, let $d$ be in the order of $\frac{\Delta H}{\epsilon^{2}}\log\left(\frac{\Delta H}{\epsilon^{2}}\left(1+\left(\frac{\sigma}{\epsilon}\right)^{\frac{\mathsf{p}}{2(\mathsf{p}-1)}}\right)\right)$ (see proof for the precise definition), there exists a distribution over functions $F:\mathbb{R}^{d}\to\mathbb{R}$ and stochastic first-order oracles $\boldsymbol{g}$ such that with probability $1$, $(\boldsymbol{0})-F_{\star}\leq\Delta, FF$ is $H$-smooth and $\frac{3}{2}\sqrt{H\Delta}$-Lipschitz, and $\boldsymbol{g}$ has a finite $\mathsf{p}$-th centered moment $\sigma^{\mathsf{p}}$. Moreover, for any randomized $\mathsf{A}\in\mathsf{A}_{\mathsf{rand}}$ employed to optimize a randomly selected $F$ interacting with $\boldsymbol{g}$, to output a point $\boldsymbol{x}$ such that $\mathbb{\mathbb{E}}\left[\left\|\nabla F(\boldsymbol{x})\right\|\right]\leq\epsilon$, the number of queries of $\boldsymbol{g}$ by $\mathsf{A}$ satisfies $\gtrsim\Delta H\epsilon^{-2}+\Delta H\sigma^{\frac{\mathsf{p}}{\mathsf{p}-1}}\epsilon^{-\frac{3\mathsf{p}-2}{\mathsf{p}-1}}$.
Remark 11.
The reader familiar with the literature may find that a similar lower bound (in fact, exactly the same order) has been established by [26, 51] before, and hence may wonder about the difference. Here, we note that the lower bounds in [26, 51] are shown for a special algorithmic class known as zero-respecting algorithms [^4] A first-order algorithm is called zero-respecting if it satisfies $\boldsymbol{x}_{t}\in\cup_{s<t}\mathrm{support}(\boldsymbol{g}_{s}),\forall t\in\mathbb{\mathbb{N}}$. For more details, see Definition 1 of [3].. However, our Theorem 11 is proved for a broader family, i.e., randomized algorithms. This fact is important because Algorithm 4 is a randomized algorithm but not a zero-respecting algorithm.
Proof.
For $T\in\mathbb{\mathbb{N}}$ and $q\in\left(0,1\right)$ being determined later, let $\iota=1/2$, $\eta=\frac{1}{\sqrt{5}}-\frac{39}{140}$, $R=7\gamma\sqrt{T}$ where $\gamma=23$ and $d=\left\lceil T+32R^{2}\log\frac{2T^{2}}{q\iota}\right\rceil=\left\lceil T+32R^{2}\log\frac{4T^{2}}{q}\right\rceil=\Theta(T\log\frac{T^{2}}{q})$, we construct $\widehat{F}_{T,U}:\mathbb{R}^{d}\to\mathbb{R}$ and $\widehat{\boldsymbol{g}}_{T,U}(\boldsymbol{x},r):\mathbb{R}^{d}\times\mathcal{R}\to\mathbb{R}^{d}$ based on $F_{T}$ in Lemma 7 and $\boldsymbol{g}_{T}$ in Lemma 8.
By Lemma 6 (note that $F_{T}$ satisfies the requirements of Lemma 6 due to Lemma 7), for any $\mathsf{A}\in\mathsf{A}_{\mathsf{rand}}$ employed to optimize $\widehat{F}_{T,U}$ interacting with the stochastic first-order oracle $\widehat{\boldsymbol{g}}_{T,U}$, where $U$ is drawn from $\mathsf{Ortho}(d,T)$ uniformly, we have
$\Pr\left[\left\|\nabla\widehat{F}_{T,U}(\boldsymbol{x}_{t})\right\|\geq\frac{1}{2},\forall t\leq\frac{T-\log 4}{2q}\right]\geq\frac{1}{2}.$ (74)
By Lemma 9, $\widehat{F}_{T,U}(\boldsymbol{0})-\inf_{\boldsymbol{x}\in\mathbb{R}^{d}}\widehat{F}_{T,U}(\boldsymbol{x})\leq\widehat{f}T$ for $\widehat{f}=12$, $\widehat{F}_{T,U}$ is $^\widehat{\ell}$-smooth for $\widehat{\ell}=154$ and $\widehat{\gamma}\sqrt{T}$-Lipschitz for $\widehat{\gamma}=53$, $\mathbb{\mathbb{E}}_{r}\left[\left\|\nabla\widehat{F}_{T,U}(\boldsymbol{x})-\widehat{\boldsymbol{g}}_{T,U}(\boldsymbol{x},r)\right\|^{\mathsf{p}}\right]\leq\frac{\left(2\gamma\right)^{\mathsf{p}}\left(1-q^{\mathsf{p}-1}\right)}{(\mathsf{p}-1)q^{\mathsf{p}-1}},\forall\boldsymbol{x}\in\mathbb{R}^{d}$.
Now we set
$\lambda\triangleq\frac{4\widehat{\ell}\epsilon}{H},$ $T\triangleq\left\lfloor\frac{\widehat{\ell}\Delta}{\widehat{f}H\lambda^{2}}\right\rfloor,$ $q\triangleq\frac{1}{\left[1+(\mathsf{p}-1)\left(\frac{\widehat{\ell}\sigma}{2\gamma H\lambda}\right)^{\mathsf{p}}\right]^{\frac{1}{\mathsf{p}-1}}}.$ (75)
For any $U\in\mathsf{Ortho}(d,T)$, we introduce
$F_{U}(\boldsymbol{x})\triangleq\frac{H\lambda^{2}}{\widehat{\ell}}\widehat{F}_{T,U}(\boldsymbol{x}/\lambda)$ and $\boldsymbol{g}_{U}(\boldsymbol{x})=\frac{H\lambda}{\widehat{\ell}}\widehat{\boldsymbol{g}}_{T,U}(\boldsymbol{x}/\lambda,r).$
We observe that
[…]
$\mathbb{\mathbb{E}}_{r}\left[\left\|\nabla F_{U}(\boldsymbol{x})-\boldsymbol{g}_{U}(\boldsymbol{x})\right\|^{\mathsf{p}}\right]$ $=\left(\frac{H\lambda}{\widehat{\ell}}\right)^{\mathsf{p}}\mathbb{\mathbb{E}}_{r}\left[\left\|\nabla\widehat{F}_{T,U}(\boldsymbol{x}/\lambda)-\widehat{\boldsymbol{g}}_{T,U}(\boldsymbol{x}/\lambda,r)\right\|^{\mathsf{p}}\right]$
[…]
For any $\mathsf{A}\in\mathsf{A}_{\mathsf{rand}}$ used to optimize $F_{U}$ with $\mathsf{StocGrad}_{U}$ when $U$ is drawn from $\mathsf{Ortho}(d,T)$ uniformly, we can view as $\boldsymbol{x}_{t}/\lambda$ as the output of another $\mathsf{A}^{\lambda}\in\mathsf{A}_{\mathsf{rand}}$ interacting with $\widehat{F}_{T,U}$ and $\widehat{\boldsymbol{g}}_{T,U}$ (a similar argument is used in the proof of Lemma 6). As such, by (74),
$\Pr\left[\left\|\nabla\widehat{F}_{T,U}(\boldsymbol{x}_{t}/\lambda)\right\|\geq\frac{1}{2},\forall t\leq\frac{T-\log 4}{2q}\right]\geq\frac{1}{2},$
which implies with probability at least $1/2$
$\left\|\nabla F_{U}(\boldsymbol{x}_{t})\right\|$ $=\frac{H\lambda}{\widehat{\ell}}\left\|\nabla\widehat{F}_{T,U}(\boldsymbol{x}_{t}/\lambda)\right\|\geq\frac{H\lambda}{2\widehat{\ell}},\forall t\leq\frac{T-\log 4}{2q}.$
Therefore, we have
[…]
Lastly, we compute
[…]
$\overset{(a)}{\geq}\frac{1}{2}\left(\frac{\Delta H}{16\widehat{f}\widehat{\ell}\epsilon^{2}}-3\right)\left[1+(\mathsf{p}-1)\left(\frac{\widehat{\ell}\sigma}{2\gamma H\lambda}\right)^{\mathsf{p}}\right]^{\frac{1}{\mathsf{p}-1}}$
$\overset{(b)}{\geq}\frac{\Delta H}{64\widehat{f}\widehat{\ell}\epsilon^{2}}\left(1+(\mathsf{p}-1)^{\frac{1}{\mathsf{p}-1}}\left(\frac{\widehat{\ell}\sigma}{2\gamma H\lambda}\right)^{\frac{\mathsf{p}}{\mathsf{p}-1}}\right)$
[…]
where the $(a)$ is by $\left\lfloor\cdot\right\rfloor\geq\cdot-1$ and $\log 4\leq 2$, and $(b)$ holds due to the condition $\epsilon^{2}\leq\frac{\Delta H}{96\cdot 12\cdot 154}=\frac{\Delta H}{96\widehat{f}\widehat{\ell}}$, implying $\frac{\Delta H}{16\widehat{f}\widehat{\ell}\epsilon^{2}}-3\geq\frac{\Delta H}{32\widehat{f}\widehat{\ell}\epsilon^{2}}$, and $(1+x)^{\frac{1}{\mathsf{p}-1}}\geq 1+x^{\frac{1}{\mathsf{p}-1}}$ for $x\geq 0$ when $\mathsf{p}\in\left(1,2\right]$. So to achieve $\mathbb{\mathbb{E}}\left[\left\|\nabla F_{U}(\boldsymbol{x}_{t})\right\|\right]<\epsilon$, it requires at least $\Delta H\epsilon^{-2}+\Delta H\sigma^{\frac{\mathsf{p}}{\mathsf{p}-1}}\epsilon^{-\frac{3\mathsf{p}-2}{\mathsf{p}-1}}$ many iterations. In particular, we note that $d=\Theta(T\log\frac{T^{2}}{q})=\Theta\left(\frac{\Delta H}{\epsilon^{2}}\log\left(\frac{\Delta H}{\epsilon^{2}}\left(1+\left(\frac{\sigma}{\epsilon}\right)^{\frac{\mathsf{p}}{2(\mathsf{p}-1)}}\right)\right)\right)$. ∎
Equipped with Theorem 11, we can show the lower bound on the distributional complexity [29] for nonsmooth nonconvex optimization with heavy tails in the following Theorem 12.
Theorem 12.
For any given $\Delta>0$, $G>0$, $\mathsf{p}\in\left(1,2\right]$, $\sigma\geq 0$, $\delta>0$ and $0<\epsilon\leq\frac{\Delta}{96\cdot 12\cdot 154\delta}\land\frac{4G^{2}\delta}{9\Delta}$, let $d$ be in the order of $\frac{\Delta}{\delta\epsilon}\log\left(\frac{\Delta}{\delta\epsilon}\left(1+\left(\frac{\sigma}{\epsilon}\right)^{\frac{\mathsf{p}}{2(\mathsf{p}-1)}}\right)\right)$, there exists a distribution over functions $F:\mathbb{R}^{d}\to\mathbb{R}$, and stochastic first-order oracles $\boldsymbol{g}$ such that with probability $1$, $(\boldsymbol{0})-F_{\star}\leq\Delta, FF$ is $G$-Lipschitz and $\boldsymbol{g}$ has a finite $\mathsf{p}$-th centered moment $\sigma^{\mathsf{p}}$. Moreover, for any randomized $\mathsf{A}\in\mathsf{A}_{\mathsf{rand}}$ employed to optimize a randomly selected $F$ interacting with $\boldsymbol{g}$, to output a point $\boldsymbol{x}$ such that $\mathbb{\mathbb{E}}\left[\left\|\nabla F(\boldsymbol{x})\right\|_{\delta}\right]\leq\epsilon$, the number of queries of $\boldsymbol{g}$ by $\mathsf{A}$ satisfies $\gtrsim\Delta\delta^{-1}\epsilon^{-1}+\Delta\sigma^{\frac{\mathsf{p}}{\mathsf{p}-1}}\delta^{-1}\epsilon^{-\frac{2\mathsf{p}-1}{\mathsf{p}-1}}$.
Proof.
Let $H\triangleq\epsilon/\delta$ in the following. Note that $\epsilon\leq\sqrt{\frac{\Delta H}{96\cdot 12\cdot 154}}$ due to our requirement on $\epsilon$. So we can consider the same distribution on $F$ and $\boldsymbol{g}$ as in Theorem 11. Importantly, $F$ is $H$-smooth and $\frac{3}{2}\sqrt{H\Delta}=\frac{3}{2}\sqrt{\frac{\epsilon\Delta}{\delta}}\leq G$-Lipschitz (again due to the condition on $\epsilon$) with probability $1$.
If $\mathsf{A}\in\mathsf{A}_{\mathsf{rand}}$ finds a point $\boldsymbol{x}$ such that $\mathbb{\mathbb{E}}\left[\left\|\nabla F(\boldsymbol{x})\right\|_{\delta}\right]\leq\epsilon$ (i.e., a $(\delta,\epsilon)$-stationary point of $F$), then by Proposition 14 of [7] (or $\nu=1$ in Lemma 12 given later), it also satisfies $\mathbb{\mathbb{E}}\left[\left\|\nabla F(\boldsymbol{x})\right\|\right]\leq\epsilon+H\delta=2\epsilon$. Therefore, Theorem 11 implies the number of queries of $\boldsymbol{g}$ by $\mathsf{A}$ is at least $\gtrsim\Delta H\epsilon^{-2}+\Delta H\sigma^{\frac{\mathsf{p}}{\mathsf{p}-1}}\epsilon^{-\frac{3\mathsf{p}-2}{\mathsf{p}-1}}=\Delta\delta^{-1}\epsilon^{-1}+\Delta\sigma^{\frac{\mathsf{p}}{\mathsf{p}-1}}\delta^{-1}\epsilon^{-\frac{2\mathsf{p}-1}{\mathsf{p}-1}}$. ∎
Finally, we are able to prove Theorem 6.
Proof of Theorem 6.
We recall the fact that lower bounds on the distributional complexity imply lower bounds on the minimax complexity [29]. Thus, Theorem 6 can be directly concluded from Theorem 12. ∎
Appendix F More Details about Further Extensions
This section contains more details of Section 5.
F.1 Full Results for Smooth $\ell_{t}$
We present the full version of Theorem 7 and Corollary 4.
Theorem 13 (full version of Theorem 7).
Under Assumption 1 (with replacing the third point by Condition 1) and let $S_{T}(\boldsymbol{x})\triangleq\left(H\sum_{t=1}^{T}\ell_{t}(\boldsymbol{x})-\ell_{t}^{\inf}\right)\land\left(H\sum_{t=1}^{T}\ell_{t}(\boldsymbol{x})-\ell_{t}^{\star}+\sum_{t=1}^{T}\left\|\nabla\ell_{t}(\boldsymbol{x}_{t}^{\star})\right\|^{2}\right)$ where $\ell_{t}^{\inf}\triangleq\inf_{\boldsymbol{x}\in\mathbb{R}^{d}}\ell_{t}(\boldsymbol{x})$, $\boldsymbol{x}_{t}^{\star}\triangleq\mathbb{\mathrm{argmin}}_{\boldsymbol{x}\in\mathcal{X}}\ell_{t}(\boldsymbol{x})$, and $\ell_{t}^{\star}\triangleq\ell_{t}(\boldsymbol{x}_{t}^{\star})$:
  • taking $\eta_{t}=\frac{1}{4H}\land\gamma D\land\frac{D}{\sigma t^{1/\mathsf{p}}}$ for any $\gamma>0$ in $\mathsf{A}\in\left\{\mathsf{OGD},\mathsf{DA}\right\}$, we have
    $\mathbb{\mathbb{E}}\left[\mathsf{R}_{T}^{\mathsf{A}}(\boldsymbol{x})\right]\lesssim HD^{2}+D\left(\frac{1}{\gamma}+\gamma S_{T}(\boldsymbol{x})\right)+\sigma DT^{1/\mathsf{p}},\forall\boldsymbol{x}\in\mathcal{X}.$
  • taking $\eta=D/\sqrt{2}$ in $\mathsf{A}=\mathsf{AdaGrad}$, we have
    $\mathbb{\mathbb{E}}\left[\mathsf{R}_{T}^{\mathsf{A}}(\boldsymbol{x})\right]\lesssim HD^{2}+D\sqrt{S_{T}(\boldsymbol{x})}+\sigma DT^{1/\mathsf{p}},\forall\boldsymbol{x}\in\mathcal{X}.$
Remark 12.
First, $\boldsymbol{x}_{t}^{\star}$ must exist since $\mathcal{X}$ is a nonempty compact convex set and $\ell_{t}$ is closed convex. Next, when $\ell_{t}\geq 0$ on $\mathbb{R}^{d}$, we have $S_{T}(\boldsymbol{x})\leq H\sum_{t=1}^{T}\ell_{t}(\boldsymbol{x})-\ell_{t}^{\inf}\leq H\sum_{t=1}^{T}\ell_{t}(\boldsymbol{x})$. This fact, together with the replacement of $\gamma$ by $\gamma/\sqrt{H}$ for $\mathsf{OGD}$ and $\mathsf{DA}$, recovers Theorem 7.
Proof.
Under Condition 1, the following famous inequality holds (regardless of the convexity of $\ell_{t}$)
$\left\|\nabla\ell_{t}(\boldsymbol{x})\right\|^{2}\leq 2H(\ell_{t}(\boldsymbol{x})-\ell_{t}^{\inf}),\forall\boldsymbol{x}\in\mathbb{R}^{d}.$ (76)
For convex $\ell_{t}$, another similar inequality is also true
$\left\|\nabla\ell_{t}(\boldsymbol{x})-\nabla\ell_{t}(\boldsymbol{x}_{t}^{\star})\right\|^{2}\leq 2H(\ell_{t}(\boldsymbol{x})-\ell_{t}^{\star}),\forall\boldsymbol{x}\in\mathcal{X},$
which implies
$\left\|\nabla\ell_{t}(\boldsymbol{x})\right\|^{2}\leq\left(\sqrt{2H(\ell_{t}(\boldsymbol{x})-\ell_{t}^{\star})}+\left\|\nabla\ell_{t}(\boldsymbol{x}_{t}^{\star})\right\|\right)^{2}\leq 3H(\ell_{t}(\boldsymbol{x})-\ell_{t}^{\star})+3\left\|\nabla\ell_{t}(\boldsymbol{x}_{t}^{\star})\right\|^{2},\forall\boldsymbol{x}\in\mathcal{X}.$ (77)
By (76) and (77), there is
$\sum_{t=1}^{T}\left\|\nabla\ell_{t}(\boldsymbol{x}_{t})\right\|^{2}$ $\leq 3\sum_{t=1}^{T}\left[H(\ell_{t}(\boldsymbol{x}_{t})-\ell_{t}^{\inf})\land\left(H(\ell_{t}(\boldsymbol{x}_{t})-\ell_{t}^{\star})+\left\|\nabla\ell_{t}(\boldsymbol{x}_{t}^{\star})\right\|^{2}\right)\right]$
$=3H\mathsf{R}_{T}^{\mathsf{A}}(\boldsymbol{x})+3\sum_{t=1}^{T}\left[H(\ell_{t}(\boldsymbol{x})-\ell_{t}^{\inf})\land\left(H(\ell_{t}(\boldsymbol{x})-\ell_{t}^{\star})+\left\|\nabla\ell_{t}(\boldsymbol{x}_{t}^{\star})\right\|^{2}\right)\right]$
$\leq 3H\mathsf{R}_{T}^{\mathsf{A}}(\boldsymbol{x})+3S_{T}(\boldsymbol{x}).$ (78)
With (78) on hand, we can extend the regret bound of $\mathsf{A}\in\left\{\mathsf{OGD},\mathsf{DA},\mathsf{AdaGrad}\right\}$ to smooth $\ell_{t}$.
For $\mathsf{OGD}$, note that (13) still holds up to the change of $G$ by $\left\|\nabla\ell_{t}(\boldsymbol{x}_{t})\right\|$, meaning that for any $\boldsymbol{x}\in\mathcal{X}$,
$\sum_{t=1}^{T}\left\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t}-\boldsymbol{x}\right\rangle\leq\frac{D^{2}}{\eta_{T}}+\sum_{t=1}^{T}\eta_{t}\left\|\nabla\ell_{t}(\boldsymbol{x}_{t})\right\|^{2}+\mathsf{C}(\mathsf{p})\eta_{t}^{\mathsf{p}-1}\left\|\boldsymbol{\epsilon}_{t}\right\|^{\mathsf{p}}D^{2-\mathsf{p}}.$ (79)
For $\mathsf{DA}$, note that (19) still holds up to the change of $G$ by $\left\|\nabla\ell_{t}(\boldsymbol{x}_{t})\right\|$, meaning that for any $\boldsymbol{x}\in\mathcal{X}$,
$\sum_{t=1}^{T}\left\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t}-\boldsymbol{x}\right\rangle\leq\frac{D^{2}}{\eta_{T}}+\sum_{t=1}^{T}\eta_{t-1}\left\|\nabla\ell_{t}(\boldsymbol{x}_{t})\right\|^{2}+\mathsf{C}(\mathsf{p})\eta_{t-1}^{\mathsf{p}-1}\left\|\boldsymbol{\epsilon}_{t}\right\|^{\mathsf{p}}D^{2-\mathsf{p}},$ (80)
where $\eta_{0}=\eta_{1}$. Now we write $\eta_{t}=\eta\land\frac{D}{\sigma t^{1/\mathsf{p}}}$ for $\eta=\frac{1}{4H}\land\gamma D$. For $\mathsf{A}\in\left\{\mathsf{OGD},\mathsf{DA}\right\}$, taking expectations on both sides of (79) or (79) implies
$\mathbb{\mathbb{E}}\left[\mathsf{R}_{T}^{\mathsf{A}}(\boldsymbol{x})\right]$ $\lesssim\frac{D^{2}}{\eta}+\sigma DT^{1/\mathsf{p}}+\eta\mathbb{\mathbb{E}}\left[\sum_{t=1}^{T}\left\|\nabla\ell_{t}(\boldsymbol{x}_{t})\right\|^{2}\right]$
[…]
Since $\eta=\frac{1}{4H}\land\gamma D\Rightarrow 1-3\eta H\geq\frac{1}{4}$, $\eta\leq\gamma D$ and $\frac{1}{\eta}\leq 4H+\frac{1}{\gamma D}$, we hence conclude
$\mathbb{\mathbb{E}}\left[\mathsf{R}_{T}^{\mathsf{A}}(\boldsymbol{x})\right]\lesssim\frac{D^{2}}{\eta(1-3\eta H)}+\frac{3\eta S_{T}(\boldsymbol{x})}{1-3\eta H}+\frac{\sigma DT^{1/\mathsf{p}}}{1-3\eta H}\lesssim HD^{2}+D\left(\frac{1}{\gamma}+\gamma S_{T}(\boldsymbol{x})\right)+\sigma DT^{1/\mathsf{p}}.$
For $\mathsf{A}=\mathsf{AdaGrad}$, note that (25) still holds, i.e., for any $\boldsymbol{x}\in\mathcal{X}$,
$\sum_{t=1}^{T}\left\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t}-\boldsymbol{x}\right\rangle\leq\sqrt{2}\left(\frac{D^{2}}{2\eta}+\eta\right)\left[\sqrt{\sum_{t=1}^{T}\left\|\nabla\ell_{t}(\boldsymbol{x}_{t})\right\|^{2}}+\left(\sum_{t=1}^{T}\left\|\boldsymbol{\epsilon}_{t}\right\|^{\mathsf{p}}\right)^{\frac{1}{\mathsf{p}}}\right].$ (81)
We take expectations on both sides of (81), then apply Hölder’s inequality to have $\mathbb{\mathbb{E}}\left[\left(\sum_{t=1}^{T}\left\|\boldsymbol{\epsilon}_{t}\right\|^{\mathsf{p}}\right)^{\frac{1}{\mathsf{p}}}\right]\leq\left(\sum_{t=1}^{T}\mathbb{\mathbb{E}}\left[\left\|\boldsymbol{\epsilon}_{t}\right\|^{\mathsf{p}}\right]\right)^{\frac{1}{\mathsf{p}}}\leq\sigma T^{\frac{1}{\mathsf{p}}}$, and plug in $\eta=D/\sqrt{2}$ to obtain
$\mathbb{\mathbb{E}}\left[\mathsf{R}_{T}^{\mathsf{A}}(\boldsymbol{x})\right]\lesssim D\mathbb{\mathbb{E}}\left[\sqrt{\sum_{t=1}^{T}\left\|\nabla\ell_{t}(\boldsymbol{x}_{t})\right\|^{2}}\right]+\sigma DT^{1/\mathsf{p}}\leq D\sqrt{\mathbb{\mathbb{E}}\left[\sum_{t=1}^{T}\left\|\nabla\ell_{t}(\boldsymbol{x}_{t})\right\|^{2}\right]}+\sigma DT^{1/\mathsf{p}},$
where the last step is again by Hölder’s inequality. Combine the above inequality with (78) to conclude
$\mathbb{\mathbb{E}}\left[\mathsf{R}_{T}^{\mathsf{A}}(\boldsymbol{x})\right]$ $\lesssim D\sqrt{H\mathbb{\mathbb{E}}\left[\mathsf{R}_{T}^{\mathsf{A}}(\boldsymbol{x})\right]+S_{T}(\boldsymbol{x})}+\sigma DT^{1/\mathsf{p}}$
$\Rightarrow\mathbb{\mathbb{E}}\left[\mathsf{R}_{T}^{\mathsf{A}}(\boldsymbol{x})\right]$ $\lesssim HD^{2}+D\sqrt{S_{T}(\boldsymbol{x})}+\sigma DT^{1/\mathsf{p}}.$
Corollary 8 (full version of Corollary 4).
Under Assumption 1 (with replacing the third point by Condition 1) for $\ell_{t}(\boldsymbol{x})=F(\boldsymbol{x})$ and let $\boldsymbol{x}^{\star}\triangleq\mathbb{\mathrm{argmin}}_{\boldsymbol{x}\in\mathcal{X}}F(\boldsymbol{x})$ and $\bar{\boldsymbol{x}}_{T}\triangleq\frac{1}{T}\sum_{t=1}^{T}\boldsymbol{x}_{t}$, taking $\eta_{t}=\frac{1}{4H}\land\frac{D}{\left\|\nabla F(\boldsymbol{x}^{\star})\right\|\sqrt{T}}\land\frac{D}{\sigma t^{1/\mathsf{p}}}$ in $\mathsf{A}\in\left\{\mathsf{OGD},\mathsf{DA}\right\}$ or $\eta=D/\sqrt{2}$ in $\mathsf{A}=\mathsf{AdaGrad}$, we have
$\mathbb{\mathbb{E}}\left[F(\bar{\boldsymbol{x}}_{T})-F(\boldsymbol{x}^{\star})\right]\leq\frac{\mathbb{\mathbb{E}}\left[\mathsf{R}_{T}^{\mathsf{A}}(\boldsymbol{x}^{\star})\right]}{T}\lesssim\frac{HD^{2}}{T}+\frac{\left\|\nabla F(\boldsymbol{x}^{\star})\right\|D}{\sqrt{T}}+\frac{\sigma D}{T^{1-\frac{1}{\mathsf{p}}}}.$
Remark 13.
When $\boldsymbol{x}^{\star}\in\mathcal{X}$ is also a global minimizer (i.e., $\boldsymbol{x}^{\star}\in\mathrm{arginf}_{\boldsymbol{x}\in\mathbb{R}^{d}}F(\boldsymbol{x})$), we have $\left\|\nabla F(\boldsymbol{x}^{\star})\right\|=0$ since $F$ is differentiable. This fact recovers Corollary 4.
Proof.
By convexity, $F(\bar{\boldsymbol{x}}_{T})-F(\boldsymbol{x}^{\star})\leq\frac{\sum_{t=1}^{T}F(\boldsymbol{x}_{t})-F(\boldsymbol{x}^{\star})}{T}=\frac{\mathsf{R}_{T}^{\mathsf{A}}(\boldsymbol{x}^{\star})}{T}$ is valid for any OCO algorithm $\mathsf{A}$. We conclude from invoking Theorem 13 with $\gamma=\frac{1}{\left\|\nabla F(\boldsymbol{x}^{\star})\right\|\sqrt{T}}$ for $\mathsf{A}\in\left\{\mathsf{OGD},\mathsf{DA}\right\}$ and $\eta=D/\sqrt{2}$ for $\mathsf{A}=\mathsf{AdaGrad}$ and combining the fact $S_{T}(\boldsymbol{x}^{\star})\leq H\sum_{t=1}^{T}F(\boldsymbol{x}^{\star})-F(\boldsymbol{x}^{\star})+\sum_{t=1}^{T}\left\|\nabla F(\boldsymbol{x}^{\star})\right\|^{2}=\left\|\nabla F(\boldsymbol{x}^{\star})\right\|^{2}T$ in this case. ∎
F.2 An Optimistic Algorithm under Heavy Tails
As discussed in Section 5, our goal is to handle broader cases with optimistic algorithms. To do so, we first introduce the following new Assumption 3:
Assumption 3.
We consider the following series of assumptions:
  • $\mathcal{X}\subset\mathbb{R}^{d}$ is a nonempty closed convex set bounded by $D$, i.e., $\sup_{\boldsymbol{x},\boldsymbol{y}\in\mathcal{X}}\left\|\boldsymbol{x}-\boldsymbol{y}\right\|\leq D$.
  • $\ell_{t}:\mathcal{X}\to\mathbb{R}$ is closed convex for all $t\in\left[T\right]$.
  • $\ell_{t}$ is $(G_{t},H_{t},\nu)$-general nonsmooth on $\mathcal{X}$, i.e., there exists $G_{t}\geq 0$, $H_{t}\geq 0$ and $\nu\in\left(0,1\right]$ such that $G_{t}+H_{t}>0$ and $\left\|\nabla\ell_{t}(\boldsymbol{x})-\nabla\ell_{t}(\boldsymbol{y})\right\|\leq 2G_{t}+H_{t}\left\|\boldsymbol{x}-\boldsymbol{y}\right\|^{\nu},\forall\boldsymbol{z}\in\mathcal{X},\nabla\ell_{t}(\boldsymbol{z})\in\partial\ell_{t}(\boldsymbol{z}),\boldsymbol{z}\in\left\{\boldsymbol{x},\boldsymbol{y}\right\}$, for all $t\in\left[T\right]$.
  • Given a point $\boldsymbol{x}_{t}\in\mathcal{X}$ at the $t$-th iteration, one can query $\boldsymbol{g}_{t}\in\mathbb{R}^{d}$ satisfying $\nabla\ell_{t}(\boldsymbol{x}_{t})\triangleq\mathbb{\mathbb{E}}\left[\boldsymbol{g}_{t}\mid\mathcal{F}_{t-1}\right]\in\partial\ell_{t}(\boldsymbol{x}_{t})$ and $\mathbb{\mathbb{E}}\left[\left\|\boldsymbol{\epsilon}_{t}\right\|^{\mathsf{p}}\right]\leq\sigma_{t}^{\mathsf{p}}$ for some $\mathsf{p}\in\left(1,2\right]$ and $\sigma_{t}\geq 0$, where $\mathcal{F}_{t}\triangleq\sigma(\boldsymbol{g}_{1},\dots,\boldsymbol{g}_{t})$ denotes the natural filtration and $\boldsymbol{\epsilon}_{t}\triangleq\boldsymbol{g}_{t}-\nabla\ell_{t}(\boldsymbol{x}_{t})$ is the stochastic noise.
Assumption 3 generalizes Assumption 1 since the latter is a special case of the former, i.e., when $(G_{t},H_{t},\nu)=(G,0,\nu)$ (for any $\nu\in\left(0,1\right]$) and $\sigma_{t}=\sigma$. When $G_{t}=0$, this new assumption means that each $\ell_{t}$ is locally Hölder smooth with time-varying parameters $(H_{t},\nu)$ on $\mathcal{X}$. If we further let $\nu=1$, then each $\ell_{t}$ is standard smooth with a parameter $H_{t}$.
F.2.1 Optimistic $\mathsf{AdaGrad}$
Algorithm 5 Optimistic $\mathsf{AdaGrad}$ ($\mathsf{OAdaGrad}$)
Input: initial point $\boldsymbol{x}_{1}\in\mathcal{X}$, initial hint $\boldsymbol{h}_{1}\in\mathbb{R}^{d}$, stepsize $\eta>0$ and $\gamma_{t}>0$
for $t=1$ to $T$ do
 Query a hint $\boldsymbol{h}_{t+1}$
$\eta_{t}=\eta V_{t}^{-1/2}\land\gamma_{t}$ where $V_{t}=\sum_{s=1}^{t}\left\|\boldsymbol{g}_{s}-\boldsymbol{h}_{s}\right\|^{2}$
$\boldsymbol{x}_{t+1}=\mathbb{\mathrm{argmin}}_{\boldsymbol{x}\in\mathcal{X}}\left\langle\boldsymbol{g}_{t}+\boldsymbol{h}_{t+1}-\boldsymbol{h}_{t},\boldsymbol{x}\right\rangle+\frac{\left\|\boldsymbol{x}-\boldsymbol{x}_{t}\right\|^{2}}{2\eta_{t}}$
end for
Following the famous idea of optimism, we consider an optimistic version of $\mathsf{AdaGrad}$ named $\mathsf{OAdaGrad}$ in Algorithm 5. To start with, we give the following Lemma 10. As a sanity check, one can take $\boldsymbol{h}_{t}=\boldsymbol{0}$ and $\gamma_{t}=+\infty$, then $\mathsf{OAdaGrad}$ reduces to the standard $\mathsf{AdaGrad}$ method, and Lemma 10 recovers the path-wise regret in (7). Therefore, the following result can be viewed as a further extension of $\mathsf{AdaGrad}$, meaning that we can apply our idea described before (see the paragraph under (7) or the proof around (25)) to overcome heavy-tailed noise.
Lemma 10.
Under Assumption 3, taking any hint sequence $\boldsymbol{h}_{t}\in\mathbb{R}^{d}$, $\eta=D/\sqrt{2}$, and any nonincreasing stepsize $\gamma_{t}$ in $\mathsf{OAdaGrad}$ (Algorithm 5), we have
$\sum_{t=1}^{T}\left\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t}-\boldsymbol{x}\right\rangle\lesssim\left\langle\boldsymbol{h}_{1},\boldsymbol{x}_{1}-\boldsymbol{x}\right\rangle+\frac{D^{2}}{\gamma_{T}}+D\sqrt{\sum_{t=1}^{T}\left\|\boldsymbol{g}_{t}-\boldsymbol{h}_{t}\right\|^{2}}-\sum_{t=1}^{T-1}\frac{\left\|\boldsymbol{x}_{t+1}-\boldsymbol{x}_{t}\right\|^{2}}{4\gamma_{t}},\forall\boldsymbol{x}\in\mathcal{X}.$
Proof.
Given $t\in\left[T\right]$, by the optimality condition of the update rule in Algorithm 5, for any $\boldsymbol{x}\in\mathcal{X}$,
$\left\langle\boldsymbol{g}_{t}+\boldsymbol{h}_{t+1}-\boldsymbol{h}_{t}+\frac{\boldsymbol{x}_{t+1}-\boldsymbol{x}_{t}}{\eta_{t}},\boldsymbol{x}_{t+1}-\boldsymbol{x}\right\rangle\leq 0,$
which implies
$\left\langle\boldsymbol{g}_{t}+\boldsymbol{h}_{t+1}-\boldsymbol{h}_{t},\boldsymbol{x}_{t+1}-\boldsymbol{x}\right\rangle$ $\leq\frac{\left\langle\boldsymbol{x}_{t}-\boldsymbol{x}_{t+1},\boldsymbol{x}_{t+1}-\boldsymbol{x}\right\rangle}{\eta_{t}}$
$=\frac{\left\|\boldsymbol{x}-\boldsymbol{x}_{t}\right\|^{2}-\left\|\boldsymbol{x}-\boldsymbol{x}_{t+1}\right\|^{2}-\left\|\boldsymbol{x}_{t+1}-\boldsymbol{x}_{t}\right\|^{2}}{2\eta_{t}},$
Therefore, we know
$\left\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t}-\boldsymbol{x}\right\rangle=$ $\left\langle\boldsymbol{g}_{t}+\boldsymbol{h}_{t+1}-\boldsymbol{h}_{t},\boldsymbol{x}_{t+1}-\boldsymbol{x}\right\rangle+\left\langle\boldsymbol{g}_{t}-\boldsymbol{h}_{t},\boldsymbol{x}_{t}-\boldsymbol{x}_{t+1}\right\rangle+\left\langle\boldsymbol{h}_{t},\boldsymbol{x}_{t}-\boldsymbol{x}\right\rangle-\left\langle\boldsymbol{h}_{t+1},\boldsymbol{x}_{t+1}-\boldsymbol{x}\right\rangle$
$\leq$ $\frac{\left\|\boldsymbol{x}-\boldsymbol{x}_{t}\right\|^{2}-\left\|\boldsymbol{x}-\boldsymbol{x}_{t+1}\right\|^{2}}{2\eta_{t}}+\left\langle\boldsymbol{g}_{t}-\boldsymbol{h}_{t},\boldsymbol{x}_{t}-\boldsymbol{x}_{t+1}\right\rangle-\frac{\left\|\boldsymbol{x}_{t+1}-\boldsymbol{x}_{t}\right\|^{2}}{2\eta_{t}}$
$+\left\langle\boldsymbol{h}_{t},\boldsymbol{x}_{t}-\boldsymbol{x}\right\rangle-\left\langle\boldsymbol{h}_{t+1},\boldsymbol{x}_{t+1}-\boldsymbol{x}\right\rangle$
$\leq$ $\frac{\left\|\boldsymbol{x}-\boldsymbol{x}_{t}\right\|^{2}-\left\|\boldsymbol{x}-\boldsymbol{x}_{t+1}\right\|^{2}}{2\eta_{t}}+\eta_{t}\left\|\boldsymbol{g}_{t}-\boldsymbol{h}_{t}\right\|^{2}-\frac{\left\|\boldsymbol{x}_{t+1}-\boldsymbol{x}_{t}\right\|^{2}}{4\eta_{t}}$
$+\left\langle\boldsymbol{h}_{t},\boldsymbol{x}_{t}-\boldsymbol{x}\right\rangle-\left\langle\boldsymbol{h}_{t+1},\boldsymbol{x}_{t+1}-\boldsymbol{x}\right\rangle,$
sum up which from $t=1$ to $T$ and drop the term $\frac{\left\|\boldsymbol{x}_{T+1}-\boldsymbol{x}\right\|^{2}}{2\eta_{T}}$ to have
$\sum_{t=1}^{T}\left\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t}-\boldsymbol{x}\right\rangle\leq$ $\frac{\left\|\boldsymbol{x}-\boldsymbol{x}_{1}\right\|^{2}}{2\eta_{1}}+\sum_{t=1}^{T-1}\left(\frac{1}{\eta_{t+1}}-\frac{1}{\eta_{t}}\right)\frac{\left\|\boldsymbol{x}_{t+1}-\boldsymbol{x}\right\|^{2}}{2}+\sum_{t=1}^{T}\eta_{t}\left\|\boldsymbol{g}_{t}-\boldsymbol{h}_{t}\right\|^{2}-\frac{\left\|\boldsymbol{x}_{t+1}-\boldsymbol{x}_{t}\right\|^{2}}{4\eta_{t}}$
$+\left\langle\boldsymbol{h}_{1},\boldsymbol{x}_{1}-\boldsymbol{x}\right\rangle-\left\langle\boldsymbol{h}_{T+1},\boldsymbol{x}_{T+1}-\boldsymbol{x}\right\rangle$
$\overset{(a)}{\leq}$ $\frac{D^{2}}{2\eta_{T}}+\sum_{t=1}^{T}\eta_{t}\left\|\boldsymbol{g}_{t}-\boldsymbol{h}_{t}\right\|^{2}-\frac{\left\|\boldsymbol{x}_{t+1}-\boldsymbol{x}_{t}\right\|^{2}}{4\eta_{t}}+\left\langle\boldsymbol{h}_{1},\boldsymbol{x}_{1}-\boldsymbol{x}\right\rangle-\left\langle\boldsymbol{h}_{T+1},\boldsymbol{x}_{T+1}-\boldsymbol{x}\right\rangle$
$\overset{(b)}{\leq}$ $\frac{D^{2}}{2\gamma_{T}}+\frac{D^{2}\sqrt{V_{T}}}{2\eta}+\sum_{t=1}^{T}\frac{\eta\left\|\boldsymbol{g}_{t}-\boldsymbol{h}_{t}\right\|^{2}}{\sqrt{V_{t}}}-\frac{\left\|\boldsymbol{x}_{t+1}-\boldsymbol{x}_{t}\right\|^{2}}{4\gamma_{t}}+\left\langle\boldsymbol{h}_{1},\boldsymbol{x}_{1}-\boldsymbol{x}\right\rangle-\left\langle\boldsymbol{h}_{T+1},\boldsymbol{x}_{T+1}-\boldsymbol{x}\right\rangle$
$\overset{(c)}{\leq}$ $\frac{D^{2}}{2\gamma_{T}}+\left(\frac{D^{2}}{2\eta}+2\eta\right)\sqrt{\sum_{t=1}^{T}\left\|\boldsymbol{g}_{t}-\boldsymbol{h}_{t}\right\|^{2}}-\sum_{t=1}^{T}\frac{\left\|\boldsymbol{x}_{t+1}-\boldsymbol{x}_{t}\right\|^{2}}{4\gamma_{t}}+\left\langle\boldsymbol{h}_{1},\boldsymbol{x}_{1}-\boldsymbol{x}\right\rangle-\left\langle\boldsymbol{h}_{T+1},\boldsymbol{x}_{T+1}-\boldsymbol{x}\right\rangle,$
where $(a)$ is by $\left\|\boldsymbol{x}_{t}-\boldsymbol{x}\right\|\leq D,\forall t\in\left[T\right]$ and $\eta_{t+1}\leq\eta_{t},\forall t\in\left[T-1\right]$ (since $\gamma_{t}$ is assumed to be nonincreasing), $(b)$ is due to $\eta_{t}=\frac{\eta}{\sqrt{V_{t}}}\land\gamma_{t}\Rightarrow\frac{1}{\eta_{T}}=\frac{\sqrt{V_{T}}}{\eta}\lor\frac{1}{\gamma_{T}}\leq\frac{\sqrt{V_{T}}}{\eta}+\frac{1}{\gamma_{T}}$, and $(c)$ follows a similar step as proving (24). Finally, we drop the term $\frac{\left\|\boldsymbol{x}_{T+1}-\boldsymbol{x}_{T}\right\|^{2}}{4\gamma_{T}}$, use $\eta=D/\sqrt{2}$, and set $\boldsymbol{h}_{T+1}=\boldsymbol{0}$ to obtain the desired bound (this step is without loss of generality, otherwise, one can simply change the $\boldsymbol{x}_{T+1}$ used above to be $\mathbb{\mathrm{argmin}}_{\boldsymbol{x}\in\mathcal{X}}\left\langle\boldsymbol{g}_{T}-\boldsymbol{h}_{T},\boldsymbol{x}\right\rangle+\frac{\left\|\boldsymbol{x}-\boldsymbol{x}_{T}\right\|^{2}}{2\eta_{T}}$). ∎
F.2.2 New Regret for $\mathsf{OAdaGrad}$
Equipped with Lemma 10, we first prove the following Theorem 14, which establishes the regret of $\mathsf{OAdaGrad}$ under heavy tails.
Theorem 14.
Under Assumption 3, taking $\boldsymbol{h}_{t}=\boldsymbol{g}_{t-1}$ where $\boldsymbol{g}_{0}\triangleq\boldsymbol{0}$, $\eta=D/\sqrt{2}$, and $\gamma_{t}=\frac{D^{1-\nu}}{\max_{s\in\left[t\right]}H_{s}}$ in $\mathsf{OAdaGrad}$ (Algorithm 5), we have
$\mathbb{\mathbb{E}}\left[\mathsf{R}_{T}^{\mathsf{OAdaGrad}}(\boldsymbol{x})\right]\lesssim D^{1+\nu}\left(\sum_{t=1}^{T}H_{t}^{\frac{2-\nu}{1-\nu}}\right)^{\frac{1-\nu}{2-\nu}}+D\sqrt{A_{T}+\sum_{t=1}^{T}G_{t}^{2}}+D\left(\sum_{t=1}^{T}\sigma_{t}^{\mathsf{p}}\right)^{\frac{1}{\mathsf{p}}},$
where $A_{T}\triangleq\left\|\nabla\ell_{1}(\boldsymbol{x}_{1})\right\|+\sum_{t=2}^{T}\sup_{\boldsymbol{x}\in\mathcal{X}}\left\|\nabla\ell_{t}(\boldsymbol{x})-\nabla\ell_{t-1}(\boldsymbol{x})\right\|^{2}$ and $\ell_{0}\triangleq 0$.
We briefly discuss Theorem 14 before proving it. First, the quantity $A_{T}$ is standard and well-known in the literature as gradient variation [6]. Next, let us consider the case $H_{t}=0$, then the bound degenerates to $D\sqrt{A_{T}+\sum_{t=1}^{T}G_{t}^{2}}+D\left(\sum_{t=1}^{T}\sigma_{t}^{\mathsf{p}}\right)^{1/\mathsf{p}}$, which further reduces to the optimal regret $GD\sqrt{T}+\sigma DT^{1/\mathsf{p}}$ under Assumption 1. In the case $G_{t}=0$, i.e., locally Hölder smooth $\ell_{t}$, it gives a regret $D^{1+\nu}\left(\sum_{t=1}^{T}H_{t}^{\frac{2-\nu}{1-\nu}}\right)^{\frac{1-\nu}{2-\nu}}+D\sqrt{A_{T}}+D\left(\sum_{t=1}^{T}\sigma_{t}^{\mathsf{p}}\right)^{1/\mathsf{p}}$. In the special situation, deterministic OCO under standard smoothness (i.e., $\sigma_{t}=0$, $H_{t}=H$, and $\nu=1$), this matches the classical result $HD^{2}+\sqrt{A_{T}}D$ [6]. But as one can see, our Theorem 14 is more general and, as far as we know, is the first bound containing gradient variation for heavy-tailed OCO.
Proof.
By Lemma 10 with $\boldsymbol{h}_{t}=\boldsymbol{g}_{t-1}$, we have
$\sum_{t=1}^{T}\left\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t}-\boldsymbol{x}\right\rangle\lesssim\frac{D^{2}}{\gamma_{T}}+D\sqrt{\sum_{t=1}^{T}\left\|\boldsymbol{g}_{t}-\boldsymbol{g}_{t-1}\right\|^{2}}-\sum_{t=1}^{T-1}\frac{\left\|\boldsymbol{x}_{t+1}-\boldsymbol{x}_{t}\right\|^{2}}{4\gamma_{t}}.$ (82)
Let $\boldsymbol{\epsilon}_{0}\triangleq\boldsymbol{0}$, we can find
$\sum_{t=1}^{T}\left\|\boldsymbol{g}_{t}-\boldsymbol{g}_{t-1}\right\|^{2}$ $=\sum_{t=1}^{T}\left\|\nabla\ell_{t}(\boldsymbol{x}_{t})-\nabla\ell_{t-1}(\boldsymbol{x}_{t})+\nabla\ell_{t-1}(\boldsymbol{x}_{t})-\nabla\ell_{t-1}(\boldsymbol{x}_{t-1})+\boldsymbol{\epsilon}_{t}-\boldsymbol{\epsilon}_{t-1}\right\|^{2}$
$\lesssim\sum_{t=1}^{T}\left\|\nabla\ell_{t}(\boldsymbol{x}_{t})-\nabla\ell_{t-1}(\boldsymbol{x}_{t})\right\|^{2}+\sum_{t=1}^{T}\left\|\nabla\ell_{t-1}(\boldsymbol{x}_{t})-\nabla\ell_{t-1}(\boldsymbol{x}_{t-1})\right\|^{2}+\sum_{t=1}^{T}\left\|\boldsymbol{\epsilon}_{t}\right\|^{2}$
$\leq A_{T}+\sum_{t=2}^{T}\left\|\nabla\ell_{t-1}(\boldsymbol{x}_{t})-\nabla\ell_{t-1}(\boldsymbol{x}_{t-1})\right\|^{2}+\sum_{t=1}^{T}\left\|\boldsymbol{\epsilon}_{t}\right\|^{2},$ (83)
where the last step is by the definition of $A_{T}$ and $\ell_{0}=0$. Now by Assumption 3, $\left\|\nabla\ell_{t-1}(\boldsymbol{x}_{t})-\nabla\ell_{t-1}(\boldsymbol{x}_{t-1})\right\|^{2}\lesssim G_{t-1}^{2}+H_{t-1}^{2}\left\|\boldsymbol{x}_{t}-\boldsymbol{x}_{t-1}\right\|^{2\nu}$, which implies
$\sum_{t=2}^{T}\left\|\nabla\ell_{t-1}(\boldsymbol{x}_{t})-\nabla\ell_{t-1}(\boldsymbol{x}_{t-1})\right\|^{2}\lesssim\sum_{t=2}^{T}G_{t-1}^{2}+H_{t-1}^{2}\left\|\boldsymbol{x}_{t}-\boldsymbol{x}_{t-1}\right\|^{2\nu}=\sum_{t=1}^{T-1}G_{t}^{2}+H_{t}^{2}\left\|\boldsymbol{x}_{t+1}-\boldsymbol{x}_{t}\right\|^{2\nu}.$ (84)
Combine (83) and (84) and use $\sqrt{a+b}\leq\sqrt{a}+\sqrt{b},\forall a,b\geq 0$ to have
$\sqrt{\sum_{t=1}^{T}\left\|\boldsymbol{g}_{t}-\boldsymbol{g}_{t-1}\right\|^{2}}$ $\lesssim\sqrt{A_{T}+\sum_{t=1}^{T-1}G_{t}^{2}}+\sqrt{\sum_{t=1}^{T-1}H_{t}^{2}\left\|\boldsymbol{x}_{t+1}-\boldsymbol{x}_{t}\right\|^{2\nu}}+\sqrt{\sum_{t=1}^{T}\left\|\boldsymbol{\epsilon}_{t}\right\|^{2}}$
$\leq\sqrt{A_{T}+\sum_{t=1}^{T-1}G_{t}^{2}}+\sqrt{\sum_{t=1}^{T-1}H_{t}^{2}\left\|\boldsymbol{x}_{t+1}-\boldsymbol{x}_{t}\right\|^{2\nu}}+\left(\sum_{t=1}^{T}\left\|\boldsymbol{\epsilon}_{t}\right\|^{\mathsf{p}}\right)^{\frac{1}{\mathsf{p}}},$ (85)
where the last step is due to $\left\|\cdot\right\|_{2}\leq\left\|\cdot\right\|_{\mathsf{p}}$ for any $\mathsf{p}\in\left[1,2\right]$.
Therefore, by (82) and (85), the following inequality holds
$\sum_{t=1}^{T}\left\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t}-\boldsymbol{x}\right\rangle\lesssim$ $\frac{D^{2}}{\gamma_{T}}+D\sqrt{\sum_{t=1}^{T-1}H_{t}^{2}\left\|\boldsymbol{x}_{t+1}-\boldsymbol{x}_{t}\right\|^{2\nu}}-\sum_{t=1}^{T-1}\frac{\left\|\boldsymbol{x}_{t+1}-\boldsymbol{x}_{t}\right\|^{2}}{4\gamma_{t}}$
$+D\sqrt{A_{T}+\sum_{t=1}^{T-1}G_{t}^{2}}+D\left(\sum_{t=1}^{T}\left\|\boldsymbol{\epsilon}_{t}\right\|^{\mathsf{p}}\right)^{\frac{1}{\mathsf{p}}}.$ (86)
We use Hölder’s inequality to bound
$\sum_{t=1}^{T-1}H_{t}^{2}\left\|\boldsymbol{x}_{t+1}-\boldsymbol{x}_{t}\right\|^{2\nu}=\sum_{t=1}^{T-1}H_{t}^{2}\left(2\nu\gamma_{t}\right)^{\nu}\frac{\left\|\boldsymbol{x}_{t+1}-\boldsymbol{x}_{t}\right\|}{\left(2\nu\gamma_{t}\right)^{\nu}}^{2\nu}\leq\left(\sum_{t=1}^{T-1}H_{t}^{\frac{2}{1-\nu}}\left(2\nu\gamma_{t}\right)^{\frac{\nu}{1-\nu}}\right)^{1-\nu}\left(\sum_{t=1}^{T-1}\frac{\left\|\boldsymbol{x}_{t+1}-\boldsymbol{x}_{t}\right\|}{2\nu\gamma_{t}}^{2}\right)^{\nu},$
which implies
$D\sqrt{\sum_{t=1}^{T-1}H_{t}^{2}\left\|\boldsymbol{x}_{t+1}-\boldsymbol{x}_{t}\right\|^{2\nu}}$ $\leq D\left(\sum_{t=1}^{T-1}H_{t}^{\frac{2}{1-\nu}}\left(2\nu\gamma_{t}\right)^{\frac{\nu}{1-\nu}}\right)^{\frac{1-\nu}{2}}\left(\sum_{t=1}^{T-1}\frac{\left\|\boldsymbol{x}_{t+1}-\boldsymbol{x}_{t}\right\|}{2\nu\gamma_{t}}^{2}\right)^{\frac{\nu}{2}}$
$\leq\frac{D^{\frac{2}{2-\nu}}\left(\sum_{t=1}^{T-1}H_{t}^{\frac{2}{1-\nu}}\left(2\nu\gamma_{t}\right)^{\frac{\nu}{1-\nu}}\right)^{\frac{1-\nu}{2-\nu}}}{2/(2-\nu)}+\frac{\sum_{t=1}^{T-1}\frac{\left\|\boldsymbol{x}_{t+1}-\boldsymbol{x}_{t}\right\|}{2\nu\gamma_{t}}^{2}}{2/\nu}$
$\lesssim D^{\frac{2}{2-\nu}}\left(\sum_{t=1}^{T-1}H_{t}^{\frac{2}{1-\nu}}\gamma_{t}^{\frac{\nu}{1-\nu}}\right)^{\frac{1-\nu}{2-\nu}}+\sum_{t=1}^{T-1}\frac{\left\|\boldsymbol{x}_{t+1}-\boldsymbol{x}_{t}\right\|^{2}}{4\gamma_{t}},$ (87)
where the second step is due to Young’s inequality. Plug (87) back into (86) to know
$\sum_{t=1}^{T}\left\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t}-\boldsymbol{x}\right\rangle\lesssim\frac{D^{2}}{\gamma_{T}}+D^{\frac{2}{2-\nu}}\left(\sum_{t=1}^{T-1}H_{t}^{\frac{2}{1-\nu}}\gamma_{t}^{\frac{\nu}{1-\nu}}\right)^{\frac{1-\nu}{2-\nu}}+D\sqrt{A_{T}+\sum_{t=1}^{T-1}G_{t}^{2}}+D\left(\sum_{t=1}^{T}\left\|\boldsymbol{\epsilon}_{t}\right\|^{\mathsf{p}}\right)^{\frac{1}{\mathsf{p}}}.$
Recall that $\gamma_{t}=\frac{D^{1-\nu}}{\max_{s\in\left[t\right]}H_{s}}$, we hence have
$\sum_{t=1}^{T}\left\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t}-\boldsymbol{x}\right\rangle\lesssim D^{1+\nu}\left(\sum_{t=1}^{T-1}H_{t}^{\frac{2-\nu}{1-\nu}}\right)^{\frac{1-\nu}{2-\nu}}+D\sqrt{A_{T}+\sum_{t=1}^{T-1}G_{t}^{2}}+D\left(\sum_{t=1}^{T}\left\|\boldsymbol{\epsilon}_{t}\right\|^{\mathsf{p}}\right)^{\frac{1}{\mathsf{p}}}.$
Finally, we conclude by taking expectations on both sides of the above inequality and using Hölder’s inequality again to obtain
$\mathbb{\mathbb{E}}\left[\left(\sum_{t=1}^{T}\left\|\boldsymbol{\epsilon}_{t}\right\|^{\mathsf{p}}\right)^{\frac{1}{\mathsf{p}}}\right]\leq\left(\mathbb{\mathbb{E}}\left[\sum_{t=1}^{T}\left\|\boldsymbol{\epsilon}_{t}\right\|^{\mathsf{p}}\right]\right)^{\frac{1}{\mathsf{p}}}\leq D\left(\sum_{t=1}^{T}\sigma_{t}^{\mathsf{p}}\right)^{\frac{1}{\mathsf{p}}}.$
An undesired point of the above Theorem 14 is requiring the knowledge of both $\nu$ and $H_{t}$ to set the stepsize $\gamma_{t}$. In the following, we show that under a slightly variant nonsmooth notion, one can relax this requirement. To do so, we consider Condition 2 to substitute the third point in Assumption 3.
Condition 2.
$\ell_{t}$ is $(G_{t},H_{t},\nu,\star)$-general nonsmooth on $\mathcal{X}$, i.e., there exists $G_{t}\geq 0$, $H_{t}\geq 0$ and $\nu\in\left(0,1\right]$ such that $G_{t}+H_{t}>0$ and $\left\|\nabla\ell_{t}(\boldsymbol{x})-\nabla\ell_{t}(\boldsymbol{x}_{t}^{\star})\right\|\lesssim G_{t}+H_{t}^{\frac{1}{1+\nu}}(\ell_{t}(\boldsymbol{x})-\ell_{t}^{\star})^{\frac{\nu}{1+\nu}},\forall\boldsymbol{x}\in\mathcal{X},\nabla\ell_{t}(\boldsymbol{x})\in\partial\ell_{t}(\boldsymbol{x})$ where $\boldsymbol{x}_{t}^{\star}\triangleq\mathbb{\mathrm{argmin}}_{\boldsymbol{x}\in\mathcal{X}}\ell_{t}(\boldsymbol{x})$ and $\ell_{t}^{\star}\triangleq\ell_{t}(\boldsymbol{x}_{t}^{\star})$, for all $t\in\left[T\right]$.
To gain some intuition of the new Condition 2, we first let $H_{t}=0$ to have $\left\|\nabla\ell_{t}(\boldsymbol{x})-\nabla\ell_{t}(\boldsymbol{x}_{t}^{\star})\right\|\lesssim G_{t}$, which is similar to $\left\|\nabla\ell_{t}(\boldsymbol{x})-\nabla\ell_{t}(\boldsymbol{y})\right\|\lesssim G_{t},\forall\boldsymbol{x},\boldsymbol{y}\in\mathcal{X}$ in Assumption 3. Next, we consider the case $G_{t}=0$, where Condition 2 states an inequality $\left\|\nabla\ell_{t}(\boldsymbol{x})-\nabla\ell_{t}(\boldsymbol{x}_{t}^{\star})\right\|\lesssim H_{t}^{\frac{1}{1+\nu}}(\ell_{t}(\boldsymbol{x})-\ell_{t}^{\star})^{\frac{\nu}{1+\nu}}$, which is known to hold if $\ell_{t}$ is $(H_{t},\nu)$-Hölder smooth on $\mathbb{R}^{d}$ and convex, which can be viewed as a generalization of Assumption 3 (though strictly speaking, we only require $\ell_{t}$ to be Hölder smooth on $\mathcal{X}$ in Assumption 3). For general $G_{t}$, $H_{t}$, and $\nu$, whether Condition 2 is strictly general than the third point in Assumption 3 is unclear. Thus, we consider it as a separate condition here.
Now, we prove a new regret for $\mathsf{OAdaGrad}$ under Condition 2, which no longer needs to know $H_{t}$ and $\nu$.
Theorem 15.
Under Assumption 3 (with replacing the third point by Condition 2), taking $\boldsymbol{h}_{t}=\boldsymbol{g}_{t-1}$ where $\boldsymbol{g}_{0}\triangleq\boldsymbol{0}$, $\eta=D/\sqrt{2}$, and $\gamma_{t}=+\infty$ in $\mathsf{OAdaGrad}$ (Algorithm 5), we have
$\mathbb{\mathbb{E}}\left[\mathsf{R}_{T}^{\mathsf{OAdaGrad}}(\boldsymbol{x})\right]\lesssim D^{1+\nu}\left(\sum_{t=1}^{T}H_{t}^{\frac{2}{1-\nu}}\right)^{\frac{1-\nu}{2}}+D\left(\sum_{t=1}^{T}H_{t}^{\frac{2}{1-\nu}}\right)^{\frac{1-\nu}{2(1+\nu)}}\left[C_{T}(\boldsymbol{x})\right]^{\frac{\nu}{1+\nu}}+D\sqrt{B_{T}+\sum_{t=1}^{T}G_{t}^{2}}+D\left(\sum_{t=1}^{T}\sigma_{t}^{\mathsf{p}}\right)^{\frac{1}{\mathsf{p}}},$
where $B_{T}\triangleq\sum_{t=1}^{T}\left\|\nabla\ell_{t}(\boldsymbol{x}_{t}^{\star})-\nabla\ell_{t-1}(\boldsymbol{x}_{t-1}^{\star})\right\|^{2}$, $C_{T}(\boldsymbol{x})\triangleq\sum_{t=1}^{T}\ell_{t}(\boldsymbol{x})-\ell_{t}^{\star}$, and $\ell_{0}\triangleq 0$.
Note that Theorem 15 and the previous Theorem 14 are not directly comparable due to different assumptions. But we can consider some special cases to better understand the difference. For example, let $H_{t}=0$, Theorem 15 degenerates to $D\sqrt{B_{T}+\sum_{t=1}^{T}G_{t}^{2}}+D\left(\sum_{t=1}^{T}\sigma_{t}^{\mathsf{p}}\right)^{\frac{1}{\mathsf{p}}}$, similar to the regret $D\sqrt{A_{T}+\sum_{t=1}^{T}G_{t}^{2}}+D\left(\sum_{t=1}^{T}\sigma_{t}^{\mathsf{p}}\right)^{1/\mathsf{p}}$ by Theorem 14 in this case (note that both $B_{T}$ and $A_{T}$ are at most in the same order of $T$). Next, if $G_{t}=0$, then these two bounds are hard to compare due to the extra term $D\left(\sum_{t=1}^{T}H_{t}^{\frac{2}{1-\nu}}\right)^{\frac{1-\nu}{2(1+\nu)}}\left[C_{T}(\boldsymbol{x})\right]^{\frac{\nu}{1+\nu}}$ in Theorem 15. But as one will see later, in convex optimization, Theorem 15 possibly leads to a better rate than Theorem 14.
Proof.
By Lemma 10 with $\boldsymbol{h}_{t}=\boldsymbol{g}_{t-1}$ and $\gamma_{t}=+\infty$, we have
$\sum_{t=1}^{T}\left\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t}-\boldsymbol{x}\right\rangle\lesssim D\sqrt{\sum_{t=1}^{T}\left\|\boldsymbol{g}_{t}-\boldsymbol{g}_{t-1}\right\|^{2}}.$ (88)
Let $\boldsymbol{\epsilon}_{0}\triangleq\boldsymbol{0}$, we can find
$\sum_{t=1}^{T}\left\|\boldsymbol{g}_{t}-\boldsymbol{g}_{t-1}\right\|^{2}$
$=$ $\sum_{t=1}^{T}\left\|\nabla\ell_{t}(\boldsymbol{x}_{t})-\nabla\ell_{t}(\boldsymbol{x}_{t}^{\star})+\nabla\ell_{t}(\boldsymbol{x}_{t}^{\star})-\nabla\ell_{t-1}(\boldsymbol{x}_{t-1}^{\star})+\nabla\ell_{t-1}(\boldsymbol{x}_{t-1}^{\star})-\nabla\ell_{t-1}(\boldsymbol{x}_{t-1})+\boldsymbol{\epsilon}_{t}-\boldsymbol{\epsilon}_{t-1}\right\|^{2}$
$\lesssim$ $\sum_{t=1}^{T}\left\|\nabla\ell_{t}(\boldsymbol{x}_{t}^{\star})-\nabla\ell_{t-1}(\boldsymbol{x}_{t-1}^{\star})\right\|^{2}+\sum_{t=1}^{T}\left\|\nabla\ell_{t}(\boldsymbol{x}_{t})-\nabla\ell_{t}(\boldsymbol{x}_{t}^{\star})\right\|^{2}+\sum_{t=1}^{T}\left\|\boldsymbol{\epsilon}_{t}\right\|^{2}$
$=$ $B_{T}+\sum_{t=1}^{T}\left\|\nabla\ell_{t}(\boldsymbol{x}_{t})-\nabla\ell_{t}(\boldsymbol{x}_{t}^{\star})\right\|^{2}+\sum_{t=1}^{T}\left\|\boldsymbol{\epsilon}_{t}\right\|^{2}.$ (89)
By Condition 2, we know $\left\|\nabla\ell_{t}(\boldsymbol{x}_{t})-\nabla\ell_{t}(\boldsymbol{x}_{t}^{\star})\right\|^{2}\lesssim G_{t}^{2}+H_{t}^{\frac{2}{1+\nu}}(\ell_{t}(\boldsymbol{x}_{t})-\ell_{t}^{\star})^{\frac{2\nu}{1+\nu}}$, which implies
$\sum_{t=1}^{T}\left\|\nabla\ell_{t}(\boldsymbol{x}_{t})-\nabla\ell_{t}(\boldsymbol{x}_{t}^{\star})\right\|^{2}\lesssim\sum_{t=1}^{T}G_{t}^{2}+H_{t}^{\frac{2}{1+\nu}}(\ell_{t}(\boldsymbol{x}_{t})-\ell_{t}^{\star})^{\frac{2\nu}{1+\nu}}.$ (90)
Combine (89) and (90) and use $\sqrt{a+b}\leq\sqrt{a}+\sqrt{b},\forall a,b\geq 0$ to have
$\sqrt{\sum_{t=1}^{T}\left\|\boldsymbol{g}_{t}-\boldsymbol{g}_{t-1}\right\|^{2}}$ $\lesssim\sqrt{B_{T}+\sum_{t=1}^{T}G_{t}^{2}}+\sqrt{\sum_{t=1}^{T}H_{t}^{\frac{2}{1+\nu}}(\ell_{t}(\boldsymbol{x}_{t})-\ell_{t}^{\star})^{\frac{2\nu}{1+\nu}}}+\sqrt{\sum_{t=1}^{T}\left\|\boldsymbol{\epsilon}_{t}\right\|^{2}}$
$\leq\sqrt{B_{T}+\sum_{t=1}^{T}G_{t}^{2}}+\sqrt{\sum_{t=1}^{T}H_{t}^{\frac{2}{1+\nu}}(\ell_{t}(\boldsymbol{x}_{t})-\ell_{t}^{\star})^{\frac{2\nu}{1+\nu}}}+\left(\sum_{t=1}^{T}\left\|\boldsymbol{\epsilon}_{t}\right\|^{\mathsf{p}}\right)^{\frac{1}{\mathsf{p}}},$ (91)
where the last step is due to $\left\|\cdot\right\|_{2}\leq\left\|\cdot\right\|_{\mathsf{p}}$ for any $\mathsf{p}\in\left[1,2\right]$.
Therefore, by (88) and (91), the following inequality holds
$\sum_{t=1}^{T}\left\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t}-\boldsymbol{x}\right\rangle\lesssim D\sqrt{\sum_{t=1}^{T}H_{t}^{\frac{2}{1+\nu}}(\ell_{t}(\boldsymbol{x}_{t})-\ell_{t}^{\star})^{\frac{2\nu}{1+\nu}}}+D\sqrt{B_{T}+\sum_{t=1}^{T}G_{t}^{2}}+D\left(\sum_{t=1}^{T}\left\|\boldsymbol{\epsilon}_{t}\right\|^{\mathsf{p}}\right)^{\frac{1}{\mathsf{p}}}.$
We use Hölder’s inequality to bound
$\sum_{t=1}^{T}H_{t}^{\frac{2}{1+\nu}}(\ell_{t}(\boldsymbol{x}_{t})-\ell_{t}^{\star})^{\frac{2\nu}{1+\nu}}\leq\left(\sum_{t=1}^{T}H_{t}^{\frac{2}{1-\nu}}\right)^{\frac{1-\nu}{1+\nu}}\left(\sum_{t=1}^{T}\ell_{t}(\boldsymbol{x}_{t})-\ell_{t}^{\star}\right)^{\frac{2\nu}{1+\nu}},$
which implies
$\sum_{t=1}^{T}\left\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t}-\boldsymbol{x}\right\rangle\lesssim D\left(\sum_{t=1}^{T}H_{t}^{\frac{2}{1-\nu}}\right)^{\frac{1-\nu}{2(1+\nu)}}\left(\sum_{t=1}^{T}\ell_{t}(\boldsymbol{x}_{t})-\ell_{t}^{\star}\right)^{\frac{\nu}{1+\nu}}+D\sqrt{B_{T}+\sum_{t=1}^{T}G_{t}^{2}}+D\left(\sum_{t=1}^{T}\left\|\boldsymbol{\epsilon}_{t}\right\|^{\mathsf{p}}\right)^{\frac{1}{\mathsf{p}}}.$
Now take expectations on both sides of the above inequality and use the fact $\mathbb{\mathbb{E}}\left[\left(\sum_{t=1}^{T}\left\|\boldsymbol{\epsilon}_{t}\right\|^{\mathsf{p}}\right)^{\frac{1}{\mathsf{p}}}\right]\leq\left(\mathbb{\mathbb{E}}\left[\sum_{t=1}^{T}\left\|\boldsymbol{\epsilon}_{t}\right\|^{\mathsf{p}}\right]\right)^{\frac{1}{\mathsf{p}}}\leq D\left(\sum_{t=1}^{T}\sigma_{t}^{\mathsf{p}}\right)^{\frac{1}{\mathsf{p}}}$ (due to Hölder’s inequality) to find for $\mathsf{A}=\mathsf{OAdaGrad}$
$\mathbb{\mathbb{E}}\left[\mathsf{R}_{T}^{\mathsf{A}}(\boldsymbol{x})\right]$ $\lesssim D\left(\sum_{t=1}^{T}H_{t}^{\frac{2}{1-\nu}}\right)^{\frac{1-\nu}{2(1+\nu)}}\mathbb{\mathbb{E}}\left[\left(\sum_{t=1}^{T}\ell_{t}(\boldsymbol{x}_{t})-\ell_{t}^{\star}\right)^{\frac{\nu}{1+\nu}}\right]+D\sqrt{B_{T}+\sum_{t=1}^{T}G_{t}^{2}}+D\left(\sum_{t=1}^{T}\sigma_{t}^{\mathsf{p}}\right)^{\frac{1}{\mathsf{p}}}$
$\leq D\left(\sum_{t=1}^{T}H_{t}^{\frac{2}{1-\nu}}\right)^{\frac{1-\nu}{2(1+\nu)}}\left(\mathbb{\mathbb{E}}\left[\sum_{t=1}^{T}\ell_{t}(\boldsymbol{x}_{t})-\ell_{t}^{\star}\right]\right)^{\frac{\nu}{1+\nu}}+D\sqrt{B_{T}+\sum_{t=1}^{T}G_{t}^{2}}+D\left(\sum_{t=1}^{T}\sigma_{t}^{\mathsf{p}}\right)^{\frac{1}{\mathsf{p}}}$
$=D\left(\sum_{t=1}^{T}H_{t}^{\frac{2}{1-\nu}}\right)^{\frac{1-\nu}{2(1+\nu)}}\left(\mathbb{\mathbb{E}}\left[\mathsf{R}_{T}^{\mathsf{A}}(\boldsymbol{x})\right]+C_{T}(\boldsymbol{x})\right)^{\frac{\nu}{1+\nu}}+D\sqrt{B_{T}+\sum_{t=1}^{T}G_{t}^{2}}+D\left(\sum_{t=1}^{T}\sigma_{t}^{\mathsf{p}}\right)^{\frac{1}{\mathsf{p}}},$ (92)
where the second step is due to Hölder’s inequality.
If $\mathbb{\mathbb{E}}\left[\mathsf{R}_{T}^{\mathsf{A}}(\boldsymbol{x})\right]\leq 0$, we are done. Hence, we only need to consider the case $\mathbb{\mathbb{E}}\left[\mathsf{R}_{T}^{\mathsf{A}}(\boldsymbol{x})\right]\geq 0$. If $\mathbb{\mathbb{E}}\left[\mathsf{R}_{T}^{\mathsf{A}}(\boldsymbol{x})\right]\leq C_{T}(\boldsymbol{x})$, then (92) implies
$\mathbb{\mathbb{E}}\left[\mathsf{R}_{T}^{\mathsf{A}}(\boldsymbol{x})\right]\lesssim D\left(\sum_{t=1}^{T}H_{t}^{\frac{2}{1-\nu}}\right)^{\frac{1-\nu}{2(1+\nu)}}\left[C_{T}(\boldsymbol{x})\right]^{\frac{\nu}{1+\nu}}+D\sqrt{B_{T}+\sum_{t=1}^{T}G_{t}^{2}}+D\left(\sum_{t=1}^{T}\sigma_{t}^{\mathsf{p}}\right)^{\frac{1}{\mathsf{p}}}.$
Otherwise, we have
$\mathbb{\mathbb{E}}\left[\mathsf{R}_{T}^{\mathsf{A}}(\boldsymbol{x})\right]\lesssim D\left(\sum_{t=1}^{T}H_{t}^{\frac{2}{1-\nu}}\right)^{\frac{1-\nu}{2(1+\nu)}}\left(\mathbb{\mathbb{E}}\left[\mathsf{R}_{T}^{\mathsf{A}}(\boldsymbol{x})\right]\right)^{\frac{\nu}{1+\nu}}+D\sqrt{B_{T}+\sum_{t=1}^{T}G_{t}^{2}}+D\left(\sum_{t=1}^{T}\sigma_{t}^{\mathsf{p}}\right)^{\frac{1}{\mathsf{p}}},$
which implies
$\mathbb{\mathbb{E}}\left[\mathsf{R}_{T}^{\mathsf{A}}(\boldsymbol{x})\right]\lesssim D^{1+\nu}\left(\sum_{t=1}^{T}H_{t}^{\frac{2}{1-\nu}}\right)^{\frac{1-\nu}{2}}+D\left(\sum_{t=1}^{T}H_{t}^{\frac{2}{1-\nu}}\right)^{\frac{1-\nu}{2(1+\nu)}}\left[C_{T}(\boldsymbol{x})\right]^{\frac{\nu}{1+\nu}}+D\sqrt{B_{T}+\sum_{t=1}^{T}G_{t}^{2}}+D\left(\sum_{t=1}^{T}\sigma_{t}^{\mathsf{p}}\right)^{\frac{1}{\mathsf{p}}}.$
So in all three cases, we have
$\mathbb{\mathbb{E}}\left[\mathsf{R}_{T}^{\mathsf{A}}(\boldsymbol{x})\right]\lesssim D^{1+\nu}\left(\sum_{t=1}^{T}H_{t}^{\frac{2}{1-\nu}}\right)^{\frac{1-\nu}{2}}+D\left(\sum_{t=1}^{T}H_{t}^{\frac{2}{1-\nu}}\right)^{\frac{1-\nu}{2(1+\nu)}}\left[C_{T}(\boldsymbol{x})\right]^{\frac{\nu}{1+\nu}}+D\sqrt{B_{T}+\sum_{t=1}^{T}G_{t}^{2}}+D\left(\sum_{t=1}^{T}\sigma_{t}^{\mathsf{p}}\right)^{\frac{1}{\mathsf{p}}}.$
Appendix G More Applications
Based on the results for $\mathsf{OAdaGrad}$ given in Appendix F, we provide more applications in this section.
G.1 General Nonsmooth Convex Optimization
With Theorems 14 and 15, we immediately obtain the following convergence rates for convex optimization. The proof of which is omitted to save space.
Corollary 9.
Under Assumption 3 for $\ell_{t}(\boldsymbol{x})=F(\boldsymbol{x})$ (meaning that $G_{t}=G$, $H_{t}=H$, and $\sigma_{t}=\sigma$) and let $\boldsymbol{x}^{\star}\triangleq\mathbb{\mathrm{argmin}}_{\boldsymbol{x}\in\mathcal{X}}F(\boldsymbol{x})$ and $\bar{\boldsymbol{x}}_{T}\triangleq\frac{1}{T}\sum_{t=1}^{T}\boldsymbol{x}_{t}$:
  • considering the same setting for $\mathsf{A}=\mathsf{OAdaGrad}$ as in Theorem 14, we have
    $\mathbb{\mathbb{E}}\left[F(\bar{\boldsymbol{x}}_{T})-F(\boldsymbol{x}^{\star})\right]\leq\frac{\mathbb{\mathbb{E}}\left[\mathsf{R}_{T}^{\mathsf{A}}(\boldsymbol{x}^{\star})\right]}{T}\lesssim\frac{\left\|\nabla F(\boldsymbol{x}^{\star})\right\|D}{T}+\frac{HD^{1+\nu}}{T^{\frac{1}{2-\nu}}}+\frac{GD}{\sqrt{T}}+\frac{\sigma D}{T^{1-\frac{1}{\mathsf{p}}}}.$
  • replacing the third point in Assumption 3 by Condition 2 and considering the same setting for $\mathsf{A}=\mathsf{OAdaGrad}$ as in Theorem 15, we have
    $\mathbb{\mathbb{E}}\left[F(\bar{\boldsymbol{x}}_{T})-F(\boldsymbol{x}^{\star})\right]\leq\frac{\mathbb{\mathbb{E}}\left[\mathsf{R}_{T}^{\mathsf{A}}(\boldsymbol{x}^{\star})\right]}{T}\lesssim\frac{\left\|\nabla F(\boldsymbol{x}^{\star})\right\|D}{T}+\frac{HD^{1+\nu}}{T^{\frac{1+\nu}{2}}}+\frac{GD}{\sqrt{T}}+\frac{\sigma D}{T^{1-\frac{1}{\mathsf{p}}}}.$
If we consider Assumption 1, (i.e., $H=0$ and $\left\|\nabla F(\boldsymbol{x})\right\|\leq G$), Corollary 9 is as fast as Corollary 1 and hence optimal. Under the same assumption used in Corollary 8 (i.e., $G=0$ and $\nu=1$), both rates degenerates to $\frac{\left\|\nabla F(\boldsymbol{x}^{\star})\right\|D}{T}+\frac{HD^{2}}{T}+\frac{\sigma D}{T^{1-\frac{1}{\mathsf{p}}}}$ and are faster than the bound $\frac{HD^{2}}{T}+\frac{\left\|\nabla F(\boldsymbol{x}^{\star})\right\|D}{\sqrt{T}}+\frac{\sigma D}{T^{1-\frac{1}{\mathsf{p}}}}$ given in Corollary 8. If we specialize in globally Hölder smooth functions (i.e., $\left\|\nabla F(\boldsymbol{x})-\nabla F(\boldsymbol{y})\right\|\leq H\left\|\boldsymbol{x}-\boldsymbol{y}\right\|^{\nu},\forall,\boldsymbol{x},\boldsymbol{y}\in\mathbb{R}^{d}$), then both bounds are new under heavy tails and the second one is faster.
G.2 General Nonsmooth Nonconvex Optimization
In this section, we move back to the nonsmooth nonconvex optimization studied before in Section 4. But instead of assuming $F$ is Lipschitz as in Assumption 2, we will use the following general Assumption 4. Note that a function satisfying Assumption 2 with a Lipschitz parameter $G$ also fits Assumption 4 for $H=0$.
Assumption 4.
We consider the following series of assumptions:
  • The objective $F$ is lower bounded by $F_{\star}\triangleq\inf_{\boldsymbol{x}\in\mathbb{R}^{d}}F(\boldsymbol{x})\in\mathbb{R}$.
  • $F$ is differentiable and well-behaved, i.e., $F(\boldsymbol{x})-F(\boldsymbol{y})=\int_{0}^{1}\left\langle\nabla F(\boldsymbol{y}+t(\boldsymbol{x}-\boldsymbol{y})),\boldsymbol{x}-\boldsymbol{y}\right\rangle\mathrm{d}t$.
  • $F$ is $(G,H,\nu)$-general nonsmooth on $\mathbb{R}^{d}$, i.e., there exists $G\geq 0$, $H\geq 0$ and $\nu\in\left(0,1\right]$ such that $G+H>0$ and $\left\|\nabla F(\boldsymbol{x})-\nabla F(\boldsymbol{y})\right\|\leq 2G+H\left\|\boldsymbol{x}-\boldsymbol{y}\right\|^{\nu},\forall\boldsymbol{x},\boldsymbol{y}\in\mathbb{R}^{d}$.
  • Given $\boldsymbol{z}_{t}\in\mathbb{R}^{d}$ at the $t$-th iteration, one can query $\boldsymbol{g}_{t}\in\mathbb{R}^{d}$ satisfying $\mathbb{\mathbb{E}}\left[\boldsymbol{g}_{t}\mid\mathcal{F}_{t-1}\right]=\nabla F(\boldsymbol{z}_{t})$ and $\mathbb{\mathbb{E}}\left[\left\|\boldsymbol{\epsilon}_{t}\right\|^{\mathsf{p}}\right]\leq\sigma^{\mathsf{p}}$ for some $\mathsf{p}\in\left(1,2\right]$ and $\sigma\geq 0$, where $\mathcal{F}_{t}$ denotes the natural filtration and $\boldsymbol{\epsilon}_{t}\triangleq\boldsymbol{g}_{t}-\nabla F(\boldsymbol{z}_{t})$ is the stochastic noise.
Our goal is still to find a $(\delta,\epsilon)$-stationary point, but under Assumption 4 this time. Fortunately, we can still start from the $\mathsf{O2NC}$ framework (Algorithm 4), since the proof of Theorem 4 only relies on the first two and the last one conditions in Assumption 2, which are the same as Assumption 4. In other words, Theorem 4 holds under this more general Assumption 4.
G.2.1 $\mathsf{OAdaGrad}$ with Reset
Algorithm 6 $\mathsf{OAdaGrad}$ with Reset ($\mathsf{OAdaGradR}$)
Input: initial point $\boldsymbol{x}_{1}\in\mathcal{B}^{d}(D)$, parameter $D>0$.
Set $\boldsymbol{g}_{0}=\boldsymbol{0}$ and $V_{0}=0$
for $n=1$ to $KT$ do
$V_{n}=V_{n-1}\mathds{1}\left[n\mod T\neq 1\right]+\left\|\boldsymbol{g}_{n}-\boldsymbol{g}_{n-1}\right\|^{2}$
$\eta_{n}=\sqrt{2}D/\sqrt{V_{n}}$
$\boldsymbol{x}_{n+1}=\mathbb{\mathrm{argmin}}_{\boldsymbol{x}\in\mathcal{B}^{d}(D)}\left\langle 2\boldsymbol{g}_{n}-\boldsymbol{g}_{n-1},\boldsymbol{x}\right\rangle+\frac{\left\|\boldsymbol{x}-\boldsymbol{x}_{n}\right\|^{2}}{2\eta_{n}}$
if $n+1\mod T=1$ do
  $\boldsymbol{x}_{n+1}=-D\frac{\boldsymbol{g}_{n}}{\left\|\boldsymbol{g}_{n}\right\|}$
end if
end for
With Theorem 4 on hand, our next task is naturally to find an online learning algorithm $\mathsf{A}$ that has a proper $K$-shifting regret under Assumption 4 (especially the third point in which). Thanks to the framework presented in Appendix F, we already have some clues, i.e., employing an optimistic algorithm. In fact, such an idea has also been studied in [7] but only for the deterministic smooth case, i.e., $\sigma=G=0$ and $\nu=1$ in Assumption 4.
To handle the more general case, particularly including heavy-tailed noise, we first present a new method called $\mathsf{OAdaGrad}$ with Reset ($\mathsf{OAdaGradR}$) in Algorithm 6, which can be viewed as running $\mathsf{OAdaGrad}$ for $T$ iterations with the hint $\boldsymbol{h}_{t}=\boldsymbol{g}_{t-1}$ and then resetting in total of $K$ times.
We clarify that the idea of reset is not new, which was originally suggested by [7], as also used previously in Theorem 5. However, the step of how to reset in $\mathsf{OAdaGradR}$, i.e., $\boldsymbol{x}_{n+1}=-D\frac{\boldsymbol{g}_{n}}{\left\|\boldsymbol{g}_{n}\right\|}$ if $n+1\mod T=1$, is critical and novel as far as we know. Why is it important? This is because the hint $\boldsymbol{h}_{t}$ is not reset to $\boldsymbol{0}$. Hence, if we reset $\boldsymbol{x}_{n+1}$ to an arbitrary point in $\mathcal{B}^{d}(D)$, then one can imagine that we will face a redundant term $\sum_{k=1}^{K}\left\langle\boldsymbol{h}_{(k-1)T},\boldsymbol{x}_{(k-1)T+1}-\boldsymbol{v}_{k}\right\rangle$ in total (see Lemma 10), which is however undesired. Instead, our specially designed way of resetting $\boldsymbol{x}_{n+1}$ can resolve this potential issue, as reflected in the following Lemma 11.
Lemma 11.
For any initial point $\boldsymbol{x}_{1}\in\mathcal{B}^{d}(D)$ and any sequence $\boldsymbol{v}_{1},\dots,\boldsymbol{v}_{K}$ satisfying $\boldsymbol{v}_{k}\in\mathcal{B}^{d}(D),\forall k\in\left[K\right]$, the online learning algorithm $\mathsf{A}=\mathsf{OAdaGradR}$ (Algorithm 6) guarantees
$\mathsf{R}_{T}^{\mathsf{A}}\left(\boldsymbol{v}_{1},\dots,\boldsymbol{v}_{K}\right)\lesssim D\sum_{k=1}^{K}\sqrt{\sum_{n=(k-1)T+1}^{kT}\left\|\boldsymbol{g}_{n}-\boldsymbol{g}_{n-1}\right\|^{2}},$
where $\mathsf{R}_{T}^{\mathsf{A}}(\boldsymbol{v}_{1},\dots,\boldsymbol{v}_{K})$ is the $K$-shifting regret defined in (9).
Proof.
Given $k\in\left[K\right]$, for $n\in\left\{(k-1)T+1,\dots,kT\right\}$, $\mathsf{OAdaGradR}$ (Algorithm 6) is the same as running $\mathsf{OAdaGrad}$ (Algorithm 5) for $T$ iterations with the feasible set $\mathcal{X}=\mathcal{B}^{d}(D)$ (which implies $\sup_{\boldsymbol{x},\boldsymbol{y}\in\mathcal{X}}\left\|\boldsymbol{x}-\boldsymbol{y}\right\|\leq 2D$), initial point $\boldsymbol{x}_{(k-1)T+1}$, hint sequence $\boldsymbol{h}_{t}=\boldsymbol{g}_{(k-1)T+t-1}$, stepsize $\eta=2D/\sqrt{2}$ and $\gamma_{t}=+\infty$, and the stochastic gradient sequence $\boldsymbol{g}_{(k-1)T+t}$. Therefore, Lemma 10 implies that for any $\boldsymbol{v}_{k}\in\mathcal{B}^{d}(D)$,
$\sum_{n=(k-1)T+1}^{kT}\left\langle\boldsymbol{g}_{n},\boldsymbol{x}_{n}-\boldsymbol{v}_{k}\right\rangle$ $\lesssim\left\langle\boldsymbol{g}_{(k-1)T},\boldsymbol{x}_{(k-1)T+1}-\boldsymbol{v}_{k}\right\rangle+D\sqrt{\sum_{t=1}^{T}\left\|\boldsymbol{g}_{(k-1)T+t}-\boldsymbol{g}_{(k-1)T+t-1}\right\|^{2}}.$
$=\left\langle\boldsymbol{g}_{(k-1)T},\boldsymbol{x}_{(k-1)T+1}-\boldsymbol{v}_{k}\right\rangle+D\sqrt{\sum_{n=(k-1)T+1}^{kT}\left\|\boldsymbol{g}_{n}-\boldsymbol{g}_{n-1}\right\|^{2}}$ (93)
Observe that if $k=1$, we have
$\left\langle\boldsymbol{g}_{(k-1)T},\boldsymbol{x}_{(k-1)T+1}-\boldsymbol{v}_{k}\right\rangle=\left\langle\boldsymbol{g}_{0},\boldsymbol{x}_{1}-\boldsymbol{v}_{1}\right\rangle=\left\langle\boldsymbol{0},\boldsymbol{x}_{1}-\boldsymbol{v}_{1}\right\rangle=0.$
If $k\neq 1$, we use $\boldsymbol{x}_{(k-1)T+1}=-D\frac{\boldsymbol{x}_{(k-1)T}}{\left\|\boldsymbol{x}_{(k-1)T}\right\|}$ and $\left\|\boldsymbol{v}_{k}\right\|\leq D$ to have
$\left\langle\boldsymbol{g}_{(k-1)T},\boldsymbol{x}_{(k-1)T+1}-\boldsymbol{v}_{k}\right\rangle=-D\left\|\boldsymbol{g}_{(k-1)T}\right\|+\left\langle\boldsymbol{g}_{(k-1)T},-\boldsymbol{v}_{k}\right\rangle\leq 0.$
Thus there is always
$\left\langle\boldsymbol{g}_{(k-1)T},\boldsymbol{x}_{(k-1)T+1}-\boldsymbol{v}_{k}\right\rangle\leq 0.$ (94)
Finally, we combine (93) and (94) to know for any $k\in\left[K\right]$,
$\sum_{n=(k-1)T+1}^{kT}\left\langle\boldsymbol{g}_{n},\boldsymbol{x}_{n}-\boldsymbol{v}_{k}\right\rangle\lesssim D\sqrt{\sum_{n=(k-1)T+1}^{kT}\left\|\boldsymbol{g}_{n}-\boldsymbol{g}_{n-1}\right\|^{2}},\forall\boldsymbol{v}_{k}\in\mathcal{B}^{d}(D),$
sum up which from $k=1$ to $K$ to obtain the following desired result
$\mathsf{R}_{T}^{\mathsf{A}}\left(\boldsymbol{v}_{1},\dots,\boldsymbol{v}_{K}\right)=\sum_{k=1}^{K}\sum_{n=(k-1)T+1}^{kT}\left\langle\boldsymbol{g}_{n},\boldsymbol{x}_{n}-\boldsymbol{v}_{k}\right\rangle\lesssim D\sum_{k=1}^{K}\sqrt{\sum_{n=(k-1)T+1}^{kT}\left\|\boldsymbol{g}_{n}-\boldsymbol{g}_{n-1}\right\|^{2}}.$
G.2.2 Convergence Rates
Equipped with Lemma 11 for $\mathsf{OAdaGradR}$, we are ready to show convergence rates under the new Assumption 4. First, we prove the following Theorem 16, which can be viewed as a generalization of Theorem 5.
Theorem 16.
Under Assumption 4 and let $\nabla\triangleq\left\|\nabla F(\boldsymbol{y}_{0})\right\|$, $\Delta\triangleq F(\boldsymbol{y}_{0})-F_{\star}$, and $\bar{\boldsymbol{z}}_{k}\triangleq\frac{1}{T}\sum_{n=(k-1)T+1}^{kT}\boldsymbol{z}_{n},\forall k\in\left[K\right]$, setting $\mathsf{A}=\mathsf{OAdaGradR}$ (Algorithm 6) in $\mathsf{O2NC}$ (Algorithm 4) with $D=\delta/T$, we have
$\mathbb{\mathbb{E}}\left[\frac{1}{K}\sum_{k=1}^{K}\left\|\nabla F(\bar{\boldsymbol{z}}_{k})\right\|_{\delta}\right]\lesssim\frac{\nabla}{KT}+\frac{\Delta}{\delta K}+\frac{H\delta^{\nu}}{T^{\frac{1}{2}+\nu}}+\frac{G}{\sqrt{T}}+\frac{\sigma}{T^{1-\frac{1}{\mathsf{p}}}}.$
Proof.
By Theorem 4 and Lemma 11, there is
$\mathbb{\mathbb{E}}\left[\sum_{k=1}^{K}\frac{1}{K}\left\|\frac{1}{T}\sum_{n=(k-1)T+1}^{kT}\nabla F(\boldsymbol{z}_{n})\right\|\right]\lesssim\frac{F(\boldsymbol{y}_{0})-F_{\star}}{DKT}+\frac{\sum_{k=1}^{K}\mathbb{\mathbb{E}}\left[\sqrt{\sum_{n=(k-1)T+1}^{kT}\left\|\boldsymbol{g}_{n}-\boldsymbol{g}_{n-1}\right\|^{2}}\right]}{KT}+\frac{\sigma}{T^{1-\frac{1}{\mathsf{p}}}}.$ (95)
We first lower bound the L.H.S. of (95). Same as (55), we still have
$\boldsymbol{z}_{n}\in\mathcal{B}^{d}(\bar{\boldsymbol{z}}_{k},\delta),\forall n\in\left\{(k-1)T+1,\dots,kT\right\}.$
By the definition of $\left\|\nabla F(\bar{\boldsymbol{z}}_{k})\right\|_{\delta}$ (see Definition 1), there is
$\left\|\nabla F(\bar{\boldsymbol{z}}_{k})\right\|_{\delta}\leq\left\|\frac{1}{T}\sum_{n=(k-1)T+1}^{kT}\nabla F(\boldsymbol{z}_{n})\right\|.$ (96)
Next, we upper bound the R.H.S. of (95). Before moving on, we list two facts that will be frequently used later, i.e., for any $n\in\left[KT\right]$,
$\left\|\boldsymbol{x}_{n}\right\|\leq D$ and $s_{n}\in\left[0,1\right].$ (97)
Now, let us upper bound $\left\|\boldsymbol{g}_{n}-\boldsymbol{g}_{n-1}\right\|^{2}$ on the R.H.S. of (95) as follows.
  • $n=1$. In this case, we recall $\boldsymbol{g}_{0}=\boldsymbol{0}$ and hence have
    $\left\|\boldsymbol{g}_{n}-\boldsymbol{g}_{n-1}\right\|^{2}$ $=\left\|\boldsymbol{g}_{1}\right\|^{2}=\left\|\boldsymbol{\epsilon}_{1}+\nabla F(\boldsymbol{z}_{1})-\nabla F(\boldsymbol{y}_{0})+\nabla F(\boldsymbol{y}_{0})\right\|^{2}$
    $\lesssim\left\|\boldsymbol{\epsilon}_{1}\right\|^{2}+\left\|\nabla F(\boldsymbol{z}_{1})-\nabla F(\boldsymbol{y}_{0})\right\|^{2}+\left\|\nabla F(\boldsymbol{y}_{0})\right\|^{2}$
[…]
    $\overset{(a)}{\lesssim}\left\|\boldsymbol{\epsilon}_{1}\right\|^{2}+G^{2}+H^{2}D^{2\nu}+\left\|\nabla F(\boldsymbol{y}_{0})\right\|^{2},$ (98)
[…]
  • $n\neq 1$. In this case, we have
    $\left\|\boldsymbol{g}_{n}-\boldsymbol{g}_{n-1}\right\|^{2}$ $=\left\|\boldsymbol{\epsilon}_{n}-\boldsymbol{\epsilon}_{n-1}+\nabla F(\boldsymbol{z}_{n})-\nabla F(\boldsymbol{z}_{n-1})\right\|^{2}$
    $\lesssim\left\|\boldsymbol{\epsilon}_{n}\right\|^{2}+\left\|\boldsymbol{\epsilon}_{n-1}\right\|^{2}+\left\|\nabla F(\boldsymbol{z}_{n})-\nabla F(\boldsymbol{z}_{n-1})\right\|^{2}$
[…]
    $\overset{(b)}{\lesssim}\left\|\boldsymbol{\epsilon}_{n}\right\|^{2}+\left\|\boldsymbol{\epsilon}_{n-1}\right\|^{2}+G^{2}+H^{2}D^{2\nu},$ (99)
[…]
Thus, we can find for any $k\in\left[K\right]$,
[…]
where $\boldsymbol{\epsilon}{}_{0}\triangleq\boldsymbol{0}$ for simplicity. As such, we obtain
$\sqrt{\sum_{n=(k-1)T+1}^{kT}\left\|\boldsymbol{g}_{n}-\boldsymbol{g}_{n-1}\right\|^{2}}$ $\lesssim\left\|\nabla F(\boldsymbol{y}_{0})\right\|\mathds{1}\left[k=1\right]+\left(HD^{\nu}+G\right)\sqrt{T}+\sqrt{\sum_{n=(k-1)T}^{kT}\left\|\boldsymbol{\epsilon}_{n}\right\|^{2}}$
$\leq\left\|\nabla F(\boldsymbol{y}_{0})\right\|\mathds{1}\left[k=1\right]+\left(HD^{\nu}+G\right)\sqrt{T}+\left(\sum_{n=(k-1)T}^{kT}\left\|\boldsymbol{\epsilon}_{n}\right\|^{\mathsf{p}}\right)^{\frac{1}{\mathsf{p}}},$
where the last step is by $\left\|\cdot\right\|_{2}\leq\left\|\cdot\right\|_{\mathsf{p}}$ for $\mathsf{p}\in\left[1,2\right]$. By Hölder’s inequality, there is
[…]
Thus, we know
$\mathbb{\mathbb{E}}\left[\sqrt{\sum_{n=(k-1)T+1}^{kT}\left\|\boldsymbol{g}_{n}-\boldsymbol{g}_{n-1}\right\|^{2}}\right]\lesssim\left\|\nabla F(\boldsymbol{y}_{0})\right\|\mathds{1}\left[k=1\right]+\left(HD^{\nu}+G\right)\sqrt{T}+\sigma T^{\frac{1}{\mathsf{p}}},\forall k\in\left[K\right].$ (100)
Finally, we plug (96) and (100) back into (95), then use $D=\delta/T$, $\nabla=\left\|\nabla F(\boldsymbol{y}_{0})\right\|$, and $\Delta=F(\boldsymbol{y}_{0})-F_{\star}$ to have
$\mathbb{\mathbb{E}}\left[\frac{1}{K}\sum_{k=1}^{K}\left\|\nabla F(\bar{\boldsymbol{z}}_{k})\right\|_{\delta}\right]\lesssim\frac{\nabla}{KT}+\frac{\Delta}{\delta K}+\frac{H\delta^{\nu}}{T^{\frac{1}{2}+\nu}}+\frac{G}{\sqrt{T}}+\frac{\sigma}{T^{1-\frac{1}{\mathsf{p}}}}.$
Similar to before, we can consider two situations where problem-dependent parameters are known or unknown, corresponding to the following Corollaries 10 and 11, respectively. To save space, the proofs are omitted since they can be easily checked.
Corollary 10.
Under the same setting in Theorem 16, suppose we have $N\geq 2$ stochastic gradient budgets, taking $K=\left\lfloor N/T\right\rfloor$ and $T=\left\lceil N/2\right\rceil\land\left(\left\lceil\left(\frac{\delta^{1+\nu}HN}{\Delta}\right)^{\frac{2}{3+2\nu}}\right\rceil\lor\left\lceil\left(\frac{\delta GN}{\Delta}\right)^{\frac{2}{3}}\right\rceil\lor\left\lceil\left(\frac{\delta\sigma N}{\Delta}\right)^{\frac{\mathsf{p}}{2\mathsf{p}-1}}\right\rceil\right)$, we have
$\mathbb{\mathbb{E}}\left[\frac{1}{K}\sum_{k=1}^{K}\left\|\nabla F(\bar{\boldsymbol{z}}_{k})\right\|_{\delta}\right]\lesssim$ $\frac{\nabla}{N}+\frac{H\delta^{\nu}}{N^{\frac{1}{2}+\nu}}+\frac{G}{\sqrt{N}}+\frac{\sigma}{N^{1-\frac{1}{\mathsf{p}}}}+\frac{\Delta}{\delta N}$
$+\frac{H^{\frac{2}{3+2\nu}}\Delta^{\frac{1+2\nu}{3+2\nu}}}{\delta^{\frac{1}{3+2\nu}}N^{\frac{1+2\nu}{3+2\nu}}}+\frac{G^{\frac{2}{3}}\Delta^{\frac{1}{3}}}{(\delta N)^{\frac{1}{3}}}+\frac{\sigma^{\frac{\mathsf{p}}{2\mathsf{p}-1}}\Delta^{\frac{\mathsf{p}-1}{2\mathsf{p}-1}}}{(\delta N)^{\frac{\mathsf{p}-1}{2\mathsf{p}-1}}}.$
Corollary 11.
Under the same setting in Theorem 16, suppose we have $N\geq 2$ stochastic gradient budgets, taking $K=\left\lfloor N/T\right\rfloor$ and $T=\left\lceil N/2\right\rceil\land\left\lceil(\delta N)^{\frac{2}{3}}\right\rceil$, we have
$\mathbb{\mathbb{E}}\left[\frac{1}{K}\sum_{k=1}^{K}\left\|\nabla F(\bar{\boldsymbol{z}}_{k})\right\|_{\delta}\right]\lesssim$ $\frac{\nabla}{N}+\frac{H\delta^{\nu}}{N^{\frac{1}{2}+\nu}}+\frac{H}{\delta^{\frac{2-\nu}{5}}N^{\frac{1+2\nu}{5}}}$
$+\frac{\Delta}{(\delta N)\land(\delta N)^{\frac{1}{3}}}+\frac{G}{\sqrt{N}\land(\delta N)^{\frac{1}{3}}}+\frac{\sigma}{N^{1-\frac{1}{\mathsf{p}}}\land(\delta N)^{\frac{2(\mathsf{p}-1)}{3\mathsf{p}}}}.$
G.2.3 A Special Case: Hölder Smooth Nonconvex Functions
We now consider a special case of $G=0$ in Assumption 4, meaning that $F$ is Hölder smooth. Now, due to the smoothness, finding an $\epsilon$-stationary point instead of a $(\delta,\epsilon)$-stationary point is more reasonable. In the following Lemma 12, we connect these two notions. Especially, when $\nu=1$, Lemma 12 recovers Proposition 14 of [7].
Lemma 12.
If $F$ is $(H,\nu)$-Hölder smooth, i.e., there exists $H>0$ and $\nu\in\left(0,1\right]$ such that $\left\|\nabla F(\boldsymbol{x})-\nabla F(\boldsymbol{y})\right\|\leq H\left\|\boldsymbol{x}-\boldsymbol{y}\right\|^{\nu},\forall\boldsymbol{x},\boldsymbol{y}\in\mathbb{R}^{d}$, then $\left\|\nabla F(\boldsymbol{x})\right\|\leq\left\|\nabla F(\boldsymbol{x})\right\|_{\delta}+H\delta^{\nu},\forall\boldsymbol{x}\in\mathbb{R}^{d},\delta>0$.
Proof.
For any fixed $\boldsymbol{x}\in\mathbb{R}^{d}$ and $\delta,\epsilon>0$, by the definition of $\left\|\nabla F(\boldsymbol{x})\right\|_{\delta}$, there exists a finite set $S\subset\mathcal{B}^{d}(\boldsymbol{x},\delta)$ such that $\frac{1}{\left|S\right|}\sum_{\boldsymbol{y}\in S}\boldsymbol{y}=\boldsymbol{x}$ and $\left\|\frac{1}{\left|S\right|}\sum_{\boldsymbol{y}\in S}\nabla F(\boldsymbol{y})\right\|\leq\left\|\nabla F(\boldsymbol{x})\right\|_{\delta}+\epsilon$. Hence, we know
$\left\|\nabla F(\boldsymbol{x})\right\|$ $\leq\left\|\frac{1}{\left|S\right|}\sum_{\boldsymbol{y}\in S}\nabla F(\boldsymbol{y})\right\|+\left\|\nabla F(\boldsymbol{x})-\frac{1}{\left|S\right|}\sum_{\boldsymbol{y}\in S}\nabla F(\boldsymbol{y})\right\|$
$\leq\left\|\nabla F(\boldsymbol{x})\right\|_{\delta}+\epsilon+\left\|\nabla F(\boldsymbol{x})-\frac{1}{\left|S\right|}\sum_{\boldsymbol{y}\in S}\nabla F(\boldsymbol{y})\right\|$
$\leq\left\|\nabla F(\boldsymbol{x})\right\|_{\delta}+\epsilon+\frac{1}{\left|S\right|}\sum_{\boldsymbol{y}\in S}\left\|\nabla F(\boldsymbol{x})-\nabla F(\boldsymbol{y})\right\|$
$\leq\left\|\nabla F(\boldsymbol{x})\right\|_{\delta}+\epsilon+\frac{H}{\left|S\right|}\sum_{\boldsymbol{y}\in S}\left\|\boldsymbol{x}-\boldsymbol{y}\right\|^{\nu}\leq\left\|\nabla F(\boldsymbol{x})\right\|_{\delta}+\epsilon+H\delta^{\nu}.$
Take $\epsilon\to 0$ to conclude. ∎
Armed with Lemma 12, we can prove the following Theorem 17.
Theorem 17.
Under Assumption 4 (with $G=0$) and let $\nabla\triangleq\left\|\nabla F(\boldsymbol{y}_{0})\right\|$, $\Delta\triangleq F(\boldsymbol{y}_{0})-F_{\star}$, and $\bar{\boldsymbol{z}}_{k}\triangleq\frac{1}{T}\sum_{n=(k-1)T+1}^{kT}\boldsymbol{z}_{n},\forall k\in\left[K\right]$, setting $\mathsf{A}=\mathsf{OAdaGradR}$ (Algorithm 6) in $\mathsf{O2NC}$ (Algorithm 4) with $D=\delta/T$, we have
$\mathbb{\mathbb{E}}\left[\frac{1}{K}\sum_{k=1}^{K}\left\|\nabla F(\bar{\boldsymbol{z}}_{k})\right\|_{\delta}\right]\lesssim\frac{\nabla}{KT}+\frac{\Delta}{\delta K}+\frac{\sigma}{T^{1-\frac{1}{\mathsf{p}}}}+H\delta^{\nu}.$
Proof.
We first invoke Lemma 12 and Theorem 16 with $G=0$, then use $T\geq 1$ to conclude. ∎
Note that now $\delta$ should also be viewed as a parameter decided by the user. Therefore, we will also choose the value of $\delta$ in the following two corollaries, corresponding to the cases where problem-dependent parameters are known and unknown, respectively. Again, the proofs are omitted to save space, and they are easy to check.
Corollary 12.
Under the same setting in Theorem 17, let $r\triangleq\mathsf{p}\nu+(\mathsf{p}-1)(1+\nu)$, suppose we have $N\geq 2$ stochastic gradient budgets, taking $K=\left\lfloor N/T\right\rfloor$, $T=\left\lceil N/2\right\rceil\land\left\lceil\frac{\sigma^{\frac{\mathsf{p}(1+\nu)}{r}}N^{\frac{\mathsf{p}\nu}{r}}}{H^{\frac{\mathsf{p}}{r}}\Delta^{\frac{\mathsf{p}\nu}{r}}}+1\right\rceil$, and $\delta=\left(\frac{\Delta T}{HN}\right)^{\frac{1}{1+\nu}}$, we have
$\mathbb{\mathbb{E}}\left[\frac{1}{K}\sum_{k=1}^{K}\left\|\nabla F(\bar{\boldsymbol{z}}_{k})\right\|_{\delta}\right]\lesssim\frac{\nabla}{N}+\frac{H^{\frac{1}{1+\nu}}\Delta^{\frac{\nu}{1+\nu}}}{N^{\frac{\nu}{1+\nu}}}+\frac{\sigma}{N^{1-\frac{1}{\mathsf{p}}}}+\frac{\sigma^{\frac{\mathsf{p}\nu}{r}}H^{\frac{\mathsf{p}-1}{r}}\Delta^{\frac{(\mathsf{p}-1)\nu}{r}}}{N^{\frac{(\mathsf{p}-1)\nu}{r}}}.$
Corollary 13.
Under the same setting in Theorem 17, suppose we have $N\geq 2$ stochastic gradient budgets, taking $K=\left\lfloor N/T\right\rfloor$, $T=\left\lceil\sqrt{N}\right\rceil$, and $\delta=1/N^{1/4}$, we have
$\mathbb{\mathbb{E}}\left[\frac{1}{K}\sum_{k=1}^{K}\left\|\nabla F(\bar{\boldsymbol{z}}_{k})\right\|_{\delta}\right]\lesssim\frac{\nabla}{N}+\frac{\Delta}{N^{\frac{1}{4}}}+\frac{H}{N^{\frac{\nu}{4}}}+\frac{\sigma}{N^{\frac{\mathsf{p}-1}{2\mathsf{p}}}}.$
As a sanity check, when $\nu=1$ (i.e., the standard smooth case), using the fact $\nabla=\left\|\nabla F(\boldsymbol{y}_{0})\right\|\lesssim\sqrt{H(F(\boldsymbol{y}_{0})-F_{\star})}=\sqrt{H\Delta}$, Corollary 12 reduces to a rate $\sqrt{\frac{H\Delta}{N}}+\frac{\sigma}{N^{1-\frac{1}{\mathsf{p}}}}+\left(\frac{\sigma^{\frac{\mathsf{p}}{\mathsf{p}-1}}H\Delta}{N}\right)^{\frac{\mathsf{p}-1}{3\mathsf{p}-2}}$ and Corollary 13 gives a rate $\frac{\Delta+H}{N^{\frac{1}{4}}}+\frac{\sigma}{N^{\frac{\mathsf{p}-1}{2\mathsf{p}}}}$, both of which match the best possible results in the respective situations [15, 26].
For general $\nu\in\left(0,1\right)$, as far as we know, no previous works consider heavy-tailed noise. Hence, both corollaries are the first and new.
Appendix H Algebraic Facts
We give three useful algebraic facts in this section.
Fact 1.
For any $T\in\mathbb{\mathbb{N}}$ and $a\in\left(0,1\right)$, there is
$\sum_{t=1}^{T-1}\frac{\sum_{s=t+1}^{T}s^{a}}{t(T-t)^{2}}\lesssim\frac{1+\log T}{T^{1-a}}.$
Proof.
Note that $\sum_{s=t+1}^{T}s^{a}\leq(T-t)T^{a}$, which implies
$\sum_{t=1}^{T-1}\frac{\sum_{s=t+1}^{T}s^{a}}{t(T-t)^{2}}\leq\sum_{t=1}^{T-1}\frac{T^{a}}{t(T-t)}=\frac{1}{T^{1-a}}\sum_{t=1}^{T-1}\frac{1}{t}+\frac{1}{T-t}=\frac{2}{T^{1-a}}\sum_{t=1}^{T-1}\frac{1}{t}\lesssim\frac{1+\log T}{T^{1-a}}.$
Fact 2.
Given $2\leq N\in\mathbb{\mathbb{N}}$, $K=\left\lfloor N/T\right\rfloor$ and $T\in\mathbb{\mathbb{N}}$ satisfying $T\leq\left\lceil N/2\right\rceil$, there is $KT\geq N/4$.
Proof.
Note that $KT=\left\lfloor N/T\right\rfloor T\geq N-T\geq(N-1)/2\geq N/4$. ∎
Fact 3.
Given $\mathsf{p}\in\left(1,2\right]$ and $q\in\left(0,1\right)$, there are $q^{\mathsf{p}-1}+(1-q)^{\mathsf{p}-1}\leq 2^{2-\mathsf{p}}$ and $1-q\leq\frac{1-q^{\mathsf{p}-1}}{\mathsf{p}-1}$.
Proof.
Note that $x^{\mathsf{p}-1}$ is concave when $\mathsf{p}\in\left(1,2\right]$, we hence have $\frac{q^{\mathsf{p}-1}+(1-q)^{\mathsf{p}-1}}{2}\leq\left(\frac{q+1-q}{2}\right)^{\mathsf{p}-1}\Rightarrow q^{\mathsf{p}-1}+(1-q)^{\mathsf{p}-1}\leq 2^{2-\mathsf{p}}$. Next, let $x=1-q\in\left(0,1\right)$, we have
$1-q\leq\frac{1-q^{\mathsf{p}-1}}{\mathsf{p}-1}\Leftrightarrow(1-x)^{\mathsf{p}-1}\leq 1-(\mathsf{p}-1)x,$
which is true by Bernoulli’s inequality. ∎